### v1

# Memento-II: Learning by Stateful Reflective Memory

Abstract
We present a new theoretical framework for continual and experiential learning in Large Language Model (LLM) agents that integrates episodic memory with reinforcement learning. We identify reflection as the central mechanism enabling agents to adapt through interaction without back-propagation or model fine-tuning, thereby breaking the conventional separation between training and deployment. Although our prior work on Memento [zhou2025memento] and case-based reasoning LLMs [guo2025optimizing, guo2024ds] demonstrated strong empirical performance in open-ended, long-horizon tasks, the underlying learning and control principles remain theoretically unexplained. This paper further addresses this gap by introducing the Stateful Reflective Decision Process (SRDP), a formal abstraction of reflective memory dynamics, where an agent maintains episodic memory and performs two key operations: writing, which stores interaction outcomes and corresponds to policy evaluation, and reading, which retrieves relevant cases to make informed reflective decisions and corresponds to policy improvement. This equivalence allows us to analyse reflection using classical tools from dynamic programming and reinforcement learning, thereby transforming reflective memory from an empirical design pattern into a well-defined control-theoretic object. We further instantiate this stateful reflective learning by integrating retrieval with entropy-regularised policy iteration. The resulting algorithm performs soft policy iteration with guaranteed convergence. We also show that as memory expands and covers the state space more densely, the composite policy converges to the optimal solution. This theoretical framework unifies heuristic approaches such as case-based reasoning, retrieval-augmented generation with principled reinforcement learning, and provides a rigorous mathematical basis for building reflective, memory-augmented agents capable of continual adaptation and general-purpose learning.
Figure 1: Learning by Stateful Reflection with Memory realises continual learning without fine-tuning LLMs by iteratively reading from (policy improvement) and writing to (policy evaluation) an evolving episodic memory that drives the agent’s online adaptation. Empirically, this Read-Write Learning paradigm has shown strong effectiveness across diverse domains, including software testing [guo2025optimizing], automated data science [guo2024ds], and deep research agents [zhou2025memento], demonstrating that memory-driven reflection can endow LLM agents with genuine self-improving capabilities.
1 Introduction
Modern machine learning is about learning from experience [turing1996intelligent, silver2025welcome]; both supervised learning and reinforcement learning can be viewed as processes of optimising a parameterised model through experience, either from human labelled examples or data-driven feedback. The objective is to learn a mapping from inputs to outputs, or in the case of reinforcement learning, from states to actions, such that the resulting function generalises effectively to new, unseen situations [shalev2014understanding, valiant1984theory]. Formally, the learner assumes that the target function is constrained by a family of parameterised models, for instance, neural networks, where the parameters determine the precise form of the mapping.
In supervised learning, these parameters are typically adjusted by minimising a loss function defined over labelled examples, whereas in reinforcement learning (RL), the supervision signal is indirect: instead of labelled outputs, the agent interacts with an environment, receives rewards, and adjusts its parameters to maximise the expected cumulative return [SuttonBarto2018]. In both cases, the parameters are updated by backpropagation, a gradient-based optimisation algorithm that propagates errors through the network to refine the model [rumelhart1986learning].
Despite their substantial empirical successes, current machine learning paradigms suffer from fundamental inefficiencies in sample complexity. Learning via backpropagation and gradient-based optimisation typically requires an enormous number of training examples, whether labelled data in supervised learning or interaction trajectories in reinforcement learning, before competent behaviour emerges [goodfellow2016deep, guo2025deepseek]. From a theoretical perspective, this reflects the difficulty of identifying generalisable structure in high-dimensional, non-linear hypothesis classes when learning is driven almost entirely by statistical correlations [shalev2014understanding]. In the absence of explicit access to semantic abstractions, compositional structure, or task-level meaning, learning algorithms must recover regularities indirectly through averaging over large numbers of samples, leading to intrinsically high sample complexity.
The training of large language models continuously exemplify this frustration. Although massive pre-training endows them with broad linguistic and reasoning abilities, adapting them to specific tasks or domains still has to resort to additional fine-tuning, such as supervised fine-tuning (SFT) or reinforcement learning with human feedback (RLHF) [ouyang2022training], each demanding further large quantities of curated data or human annotations [lightman2023let].
In contrast, humans exploit semantic representations, episodic memory, and reflective reuse of experience to dramatically reduce the number of examples required for effective adaptation [GershmanDaw2017EpisodicRL]. This gap suggests that learning mechanisms operating beyond purely statistical optimisation, and instead leveraging stateful memory and semantic reflection, may offer a principled route to lower-complexity learning without reliance on large-scale gradient updates.
While current LLMs remain data-hungry and inefficient in training, they nonetheless exhibit emerging cognitive abilities that enable iterative generalisation to unseen tasks with only a few examples. These abilities can be broadly categorised into three forms of what this paper called generalised reflection (Figure 2):
  1. 1.
    In-context learning. As early as GPT-3, researchers observed that LLMs could learn from a few human-provided examples directly within a prompt, a process known as in-context learning, where a few labelled examples guide behaviour on new tasks [brown2020language] (Figure 2 (a)). For instance, given a few examples of English-to-French translations in the prompt, GPT-3 can correctly translate a new English sentence into French without any parameter updates.
  2. 2.
    Feedback-driven reflection. Subsequent studies showed that given a specific unknown task, LLMs could refine their outputs through interaction with the environment, adjusting responses based on external feedback [Yao2022ReAct] (Figure 2 (b)). For instance, in code generation, the LLM proposes a function, runs tests to observe errors or runtime feedback, and iteratively refines the code to reach a correct solution.
  3. 3.
    Internal reasoning. More recently, LLM models have displayed the ability to perform internal reasoning via chain-of-thought processes before generating final outputs [Wei2022CoT] (Figure 2 (c)). For instance, when solving a math word problem, the model explicitly writes intermediate reasoning steps (e.g., decomposing a problem into smaller parts) before computing the final numerical answer.
Although these mechanisms vary in complexity from human-guided examples (Figure 2 (a)) to environment-driven feedback (Figure 2 (b)) to purely internal reasoning (Figure 2 (c)), they share a common principle: each relies on intermediate reflection to enable adaptation and generalisation beyond the training data. Such generalised reflection has been observed not only in pre-trained large language models (LLMs) [brown2020language, guo2025deepseek], but also in multimodal LLMs [driess2023palm] and non-language transformers [hollmann2025accurate].
(a) In-context learning (b) Feedback-driven reflection (c) Chain-of-thought
Figure 2: Three modes of generalised reflection in LLM agents: (a) in-context learning where few-shot exemplars in the prompt steer behaviour on new tasks; (b) post-output reflection via environmental feedback; (c) internal chain-of-thought and planning before emitting an answer.
In our earlier empirical work [guo2024ds, guo2025optimizing, zhou2025memento], we investigated whether reflection through memory could enable learning without parameter updates. We first introduced this paradigm with case-based reasoning LLM agents [guo2024ds], where adaptation arises from updating memory rather than fine-tuning model weights. Building on this foundation, Memento [zhou2025memento] formalised the idea within a Memory-Augmented Markov Decision Process (M-MDP) framework. Across domains such as software testing [guo2025optimizing], automated data science [guo2024ds], and deep research [zhou2025memento], this memory-driven reflection paradigm has consistently shown that LLMs can achieve continual self-improvement through accumulated experience rather than model retraining.
In this paper, we further provide a theoretical analysis and formation and leverage the capability of generalised reflection to develop a new paradigm for sample-efficient learning from experience without relying on data-intensive model training. Instead of updating model parameters, the proposed learning process operates through intermediate reflection, driven by relevant past experiences retrieved from an episodic memory [GershmanDaw2017EpisodicRL], as shown in Fig 1. This framework can unify diverse sources of experience, whether originating from internal reasoning (e.g., chain-of-thought), interactions with the environment (e.g., ReAct [Yao2022ReAct]), or human-provided labelled data, under a single, memory-based learning mechanism.
We establish a theory of redefining how learning and decision-making are modelled through a proposed Stateful Reflective Decision Process (SRDP). Unlike the traditional Markov Decision Process (MDP), our framework introduces several conceptual shifts. First, the agent’s state now combines the current situation with its episodic memory, allowing it to incorporate past experiences directly while still maintaining a coherent Markov structure in this enlarged space (namely reflected MDP). Second, each action unfolds in two stages: the agent first recalls a relevant past case from memory, and then the language model produces a concrete response or behaviour based on that case. Third, the environment’s transitions and rewards are defined through the probabilistic reasoning of the language model itself, meaning that the model’s internal uncertainty is embedded in how the world evolves. Finally, memory evolves over time as the agent writes new experiences into it. This continual update makes the system dynamic and self-adaptive, without violating the overall Markovian formulation.
Within this formulation, learning naturally emerges from the process of retrieval: policy improvement is achieved through more effective retrieval, while policy evaluation is realised through writing experiences back to memory. This mechanism closely mirrors the principles of episodic control and reinforcement learning in humans [GershmanDaw2017EpisodicRL, fountas2024human]. Consequently, such learning can be efficiently implemented using retrieval-augmented generation architectures equipped with large-scale memory systems [Borgeaud2022RETRO, Lewis2020RAG].
A key desire property is its self-learnability. With a frozen LLM [^1] While LLM parameters may evolve over time through offline updates given adequate data, such updates typically occur at a much slower timescale. Our approach does not rule out parameter learning, but instead addresses the intermediate regime in which adaptation must occur without immediate model updates. as its reasoning and reflection core, the agent improves through trial and error when data is limited and later reuses and combines past knowledge as experience grows. It balances exploitation (using known cases), exploration (targeted search in uncertain areas), and discovery (seeking new ones) through adaptive memory operations, achieving continual learning without parameter updates.
The paper is organised as follows. We first introduce the stateful reflective decision process framework. We then present theoretical results on convergence and analyse the model’s optimality and its properties as a continuously-learnable agent. Related work is reviewed next, and the paper concludes with directions for future research.
2 Stateful Reflection as Generalised Learning Methods
Recent advances in reflective prompting, such as Tree of Thoughts [yao2023tree] and Graph of Thoughts [besta2024graph], have demonstrated the potential of structured reflection to enhance LLM reasoning. In contrast, we explore a new learning paradigm that enables models to learn from experience through reflection grounded in episodic memory, rather than through parameter fine-tuning [guo2024ds, guo2025optimizing, zhou2025memento]. To support this paradigm, we introduce the Stateful Reflective Decision Process (SRDP), which formalises continual learning as a sequence of reflective interactions with memory. We begin by introducing the core concepts, leaving the detailed mathematical formulation to the sections that follow.
Covering the three modes of generalised reflection introduced earlier (Fig. 2), we define reflection in this work as follows.
Definition 1 (Reflection).
Reflection (generalised) is an iterative mechanism by which a large language model (LLM) agent progressively improves its effective policy through coordinated interaction with episodic memory, internal reasoning processes, and environmental feedback. It is not a single operation but a two-step procedure consisting of a read step followed by a write step, which together form a closed learning loop.
Figure 1 illustrates the overall workflow of an SRDP. Episodic memory refers to the structured storage of past experiences, each recording the relationship among states, actions, and outcomes, in line with the classical notion of episodic memory in cognitive psychology [tulving1983elements]. Given a task or query represented as a state (1), the agent first enters an internal reflection stage (2), during which the LLM retrieves relevant experiences from episodic memory to guide reasoning and decision-making. Based on this reflective conditioning, the LLM produces an answer or action (3), which interacts with the external environment (4) and yields evaluative feedback, such as rewards, correctness signals, or task-level assessments. This feedback is then used to update episodic memory (5), storing new experiences that capture the consequences of the agent’s decisions. The process alternates between reading from memory for policy improvement and writing to memory for policy evaluation, enabling continual learning without modifying the underlying LLM parameters.
Formally, let $\pi_{t}$ denote the agent’s effective policy at iteration $k$, implicitly represented by its prompt, episodic memory contents, and internal reasoning configuration. Reflection defines an operator $\mathcal{T}$ such that
$\pi_{t+1}=\mathcal{T}(\pi_{t})=\text{Read}\bigl(\text{Write}(\pi_{t})\bigr).$
The read step corresponds to policy improvement: the agent conditions on retrieved episodic memories, contextual exemplars, or intermediate reasoning traces to construct an improved action distribution for the current state. In LLM agents, such conditioning reshapes the effective policy directly, without parameter updates,
$\pi_{t+1}\leftarrow\mathcal{R}(\pi_{t}\mid s_{t},M_{t}),$
where $s_{t}$ denotes the current environment state and $M_{t}$ the episodic memory state. The write step corresponds to policy evaluation: the agent generates an action, plan, or answer under $\pi_{t+1}$, producing trajectories
$\tau_{t}\sim\pi_{t+1},$
whose outcomes yield evaluative feedback,
$\hat{V}(\pi_{t+1})\leftarrow\text{Eval}(\tau_{t}),$
which is subsequently stored in episodic memory for future reads.
As we shall show in the following sections, this alternating read–write mechanism is formally analogous to policy iteration in reinforcement learning. Policy improvement is realised through memory-conditioned inference, while policy evaluation is realised through interaction-driven feedback and experience accumulation. Under mild assumptions that reading yields non-decreasing policy quality with respect to accumulated feedback and that writing provides informative evaluation signals, the iterative reflection process converges to a fixed point corresponding to an optimal or self-consistent policy.
The SRDP generalises and unifies several existing paradigms. Compared with case-based reasoning (CBR), which retrieves and adapts discrete past cases for problem solving [guo2024ds, AamodtPlaza1994], SRDP embeds such retrieval in a sequential decision process, explicitly modelling temporal dependencies and memory dynamics under the Markov assumption. In contrast to retrieval-augmented generation (RAG), which conditions generation on static external documents [Borgeaud2022RETRO], SRDP treats memory as a continually evolving component, updated through interaction and grounded in feedback. From the perspective of memory-based reinforcement learning (RL), SRDP can be viewed as replacing parameter updates with LLM reflective memory updates, allowing the policy to improve through structured recall rather than gradient-based optimisation [ramani2019short].
This formulation is also strongly inspired by neuroscientific findings on the role of the hippocampus in episodic memory and reflective cognition [hami2025, McClelland1995CLS]. In humans and animals, episodic recall supports flexible reasoning and planning by replaying past experiences, an ability mirrored by the reflective cycles of SRDP. Thus, the Stateful Reflective Decision Process provides both a computational and cognitive bridge: it formalises how memory and reflection together enable continual, self-improving intelligence without explicit retraining.
Experimentally, it has been demonstrated that this approach enables learning without fine-tuning across various agentic tasks, including data-science agents [guo2025optimizing], software-testing agents [guo2024ds], and deep-research agents [zhou2025memento].
3 Learning by Stateful Reflective Experience
Mathematically, the core idea is to extend the Markov Decision Process [Puterman1994] with an intermediate reflection stage, where the agent retrieves relevant experiences from external memory before generating the final output, enabling continual and memory-based learning.
3.1 Preliminary: Markov Decision Process
We begin by introducing the Markov Decision Process (MDP) [Puterman1994], the standard framework for modelling decision-making in agents. An LLM agent, viewed as a large language model capable of reasoning and acting through interaction with its environment [christianos2023pangu, yang2025agentic, wang2025tutorial, wan2024alphazero], can be naturally described within this framework. Formally, an MDP is defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$. Here, $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $\mathcal{P}(s^{\prime}\mid s,a)$ denotes the transition kernel, specifying the probability of reaching state $s^{\prime}$ after taking action $a$ in state $s$; $\mathcal{R}(s,a)\in\mathbb{R}$ represents the expected immediate reward; and $\gamma\in[0,1)$ is the discount factor. Here, the state $s$ corresponds to the current prompt or task context. Specifically, in the context of LLM reasoning, states correspond to intermediate reasoning steps [wang2025tutorial, wang2024openr], and when interacting with the external environment, they may take the form of textual descriptions of the environment state [christianos2023pangu]. The action $a$ is the model’s generated output or decision, and the transition describes how the environment or subsequent prompt evolves after the model’s response [wang2025tutorial]. A (stochastic) policy $\pi(a\mid s)$ defines a probability distribution over actions conditioned on the current state and, in turn, induces the value function that measures the expected return when following the policy from a given state.
$V^{\pi}(s)=\mathbb{E}_{\pi,\mathcal{P}}\!\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})\,\bigg|\,s_{0}=s\right],$
with the state–action value function
$Q^{\pi}(s,a)=\mathcal{R}(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot\mid s,a)}\!\left[V^{\pi}(s^{\prime})\right].$
The optimal value functions are
$V^{\star}(s)=\max_{\pi}V^{\pi}(s),\qquad Q^{\star}(s,a)=\mathcal{R}(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot\mid s,a)}\!\left[\max_{a^{\prime}}Q^{\star}(s^{\prime},a^{\prime})\right].$
For large language model (LLM) agents [christianos2023pangu, wan2024alphazero, zeng2024token], the policy is typically given by the LLM itself, which maps input states or prompts directly to action distributions through its generative reasoning process:
$\pi_{\mathrm{LLM}}(a\mid s)\equiv p_{\mathrm{LLM}}(a\mid s;\,\theta),$
where $\theta$ denotes the (frozen in our case) parameters of the pre-trained model. Next, we extend the LLM’s stochastic policy by incorporating contributions from episodic memory, allowing past experiences to influence current decisions.
3.2 Stateful Reflective Decision Process
Building on the standard Markov Decision Process (MDP) framework introduced in the previous section, we now present the proposed Stateful Reflective Decision Process (SRDP). The classical MDP $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$ captures agent–environment interaction in its simplest form but lacks mechanisms for reflection and memory. To enable reflective, memory-driven behaviour, we extend this formulation by augmenting the agent with an episodic memory space $\mathcal{M}$, a memory retrieval policy $\mu$, and a stochastic LLM kernel $p_{\mathrm{LLM}}$ that generates context-dependent actions conditioned on both the current environment state and retrieved memory, yielding the Stateful Reflective Decision Process (SRDP). Finally, by absorbing the retrieval and LLM-generation mechanisms into effective transition and reward functions defined over the augmented state space $\mathcal{X}=\mathcal{S}\times\mathcal{M}$, we obtain a simplified induced process, referred to as the Reflected MDP. This induced process captures the full reflective dynamics of the LLM agent in a Markovian form, where a composite policy $\pi_{M}$ governs decision-making through the joint interaction of memory access and language-based reasoning. The following illustrates the relationships:
$\underbrace{\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle}_{\text{Underlying MDP}}$ $\xrightarrow{\;ext.\;}\underbrace{\Big\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{M},\mu,p_{\mathrm{LLM}}\Big\rangle}_{\text{SRDP}}\xrightarrow{\;equiv.\;}\underbrace{\Big\langle\mathcal{X}{:=}\mathcal{S}\times\mathcal{M},\ \mathcal{C},\ \mathcal{P}^{\mathrm{LLM}},\ \mathcal{R}^{\mathrm{LLM}},\ \gamma,\ \pi_{M}\Big\rangle}_{\text{Reflected MDP}}$ (1)
3.2.1 Definition
Definition 2 (Stateful Reflective Decision Process (SRDP)).
SRDP is a tuple
$\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{M},p_{\mathrm{LLM}}\rangle$
that extends the standard MDP formalism by incorporating episodic memory and a language-model-based decision kernel. The components are defined as follows:
  • $\mathcal{S}$: state space;
  • $\mathcal{A}$: environment action space;
  • $\mathcal{P}(\cdot\mid s,a)$: transition kernel;
  • $\mathcal{R}(s,a)\in\mathbb{R}$: reward function;
  • $\gamma\in[0,1)$: discount factor;
  • $\mathcal{M}$: space of finite episodic memories, where each memory $M\in\mathcal{M}$ is a multiset of cases c=(s,a,r)c=(s,a,r) [^2] In multi-state settings, $r$ can be replaced by an estimated value $Q$, where Bellman or TD updates are propagated across memory.;
  • $p_{\mathrm{LLM}}(a\mid s,c)$: the LLM kernel that generates an environment action conditioned on the current state $s$ and a retrieved case $c$.
At time $t$, the agent maintains episodic memory $M_{t}=\{c_{i}\}_{i=1}^{N_{t}}$, where $N_{t}$ is the number of cases in the current memeory, and observes the current environment state $s_{t}$. The decision process unfolds in two stages:
  1. 1.
    Retrieval Stage: A case $c_{t}$ is sampled from memory according to the retrieval policy
    $c_{t}\sim\mu(\cdot\mid s_{t},M_{t}),$
    where $\mu$ is the retrieval policy (e.g., similarity-based or learned).
  2. 2.
    Action Stage: The LLM generates an action conditioned on the retrieved case:
    $a_{t}\sim p_{\mathrm{LLM}}(\cdot\mid s_{t},c_{t}).$
This defines a composite policy:
$\pi(a\mid s,M)=\sum_{c\in M}\mu(c\mid s,M)\,p_{\mathrm{LLM}}(a\mid s,c).$ (2)
After executing the action $a_{t}$, the environment changes to $s_{t+1}\sim\mathcal{P}(\cdot\mid s_{t},a_{t})$, yields a reward $r_{t}=\mathcal{R}(s_{t},a_{t})$, and the memory updates to $M_{t+1}=\mathsf{Write}(M_{t},s_{t},a_{t},r_{t})$.
$s_{0}$$M_{0}$$c_{0}$$a_{0}$$r_{0}$$s_{1}$$M_{1}$$c_{1}$$a_{1}$$r_{1}$$s_{t}$$M_{t}$$c_{t}$$a_{t}$$r_{t}$$x_{0}$$x_{1}$$x_{t}$Environment DynamicsRetrieval Policy $\mu(c|x)$LLM Conditioning $p_{\mathrm{LLM}}(a|s,c)$Memory Write $\mathsf{Write}(M,s,a,r)$Reward Generation
Figure 3: A graphical model of the Stateful Reflective Decision Process (SRDP) showing the augmented state space $x_{t}=(s_{t},M_{t})$, retrieval actions $c_{t}$, LLM conditioning, and memory evolution. Dashed boxes indicate augmented states combining environment state and memory.
Figure 3 depicts the graphical model of the Stateful Reflective Decision Process (SRDP), characterised by a dual-action structure. The agent first performs a retrieval action $c_{t}\in M_{t}$ (green nodes), selecting a relevant past case from memory to guide its current decision. This retrieval is controlled by the policy $\mu(c\mid s,M)$, represented by the blue arrows connecting the current environment state $s_{t}$ and memory $M_{t}$ to the retrieval node $c_{t}$. The policy evaluates both the similarity between the current state and stored experiences, and the local density of memory cases. Conditioned on the retrieved case, the LLM then reflects upon the retrieved experience and generates the environment action $a_{t}\in\mathcal{A}$ (blue nodes) via its generative kernel $p_{\mathrm{LLM}}(a\mid s,c)$, shown by the green arrows. Through this hierarchical process, the retrieval stage provides contextual grounding, while the LLM generates the final action, jointly realising stateful reflection, as illustrated in Figure 1.
A key advantage of this dual-action structure is that it defines a composite policy (Eq. 2) that unifies classical reinforcement learning [SuttonBarto2018] with LLM-based decision-making [christianos2023pangu], thereby combining the strengths of both paradigms. To illustrate, consider two limiting cases. In the first case, a naive LLM produces an action $a\in c$ directly from a retrieved case $c$, effectively treating the case as a prompt. Here, the policy reduces to a standard reinforcement learning or episodic control scheme [blundell2016model, pritzel2017neural, ramani2019short], where the retrieval function $\mu$ selects the case $c$ with the highest $Q$-value, and the corresponding action is executed. In the opposite extreme, when no relevant experience exists in memory, the agent must rely entirely on the internal knowledge of the LLM to generate actions. Between these two cases lies the general setting of the SRDP, where the agent can integrate both sources of information, the reasoning capability of the LLM and the experiential knowledge stored in memory, to produce more adaptive and effective decisions.
The memory evolution mechanism forms a crucial component of the SRDP dynamics. Purple arrows demonstrate how memory grows through $\mathsf{Write}(M,s,a,r)$ operations, with experience tuples $(s_{t},a_{t},r_{t})$ being incorporated into the memory at the next time step. This creates a directed evolution $M_{0}\to M_{1}\to\cdots\to M_{t}$ where the memory size $|M_{t}|$ increases over time, expanding the agent’s retrieval options.
However, the introduction of memory fundamentally alters the dynamics of the system, making the process no longer strictly Markovian if the original state is considered. In a standard MDP, the next state depends only on the current state and action, i.e., the system satisfies $P(s_{t+1}\mid s_{t},a_{t},s_{t-1},a_{t-1},\ldots)=P(s_{t+1}\mid s_{t},a_{t}).$ However, once we introduce an evolving memory $M_{t}$ that accumulates past experience tuples $(s_{i},a_{i},r_{i})$, the transition to the next augmented state depends on the entire history through $M_{t}$. Formally, $P(s_{t+1},M_{t+1}\mid s_{t},a_{t},M_{t})\neq P(s_{t+1},M_{t+1}\mid s_{t},a_{t})$ because $M_{t}$ itself is a summary of all past interactions. We address this issue next by defining a simplified MDP that incorporates memory directly into the state representation while embedding the fixed LLM within the environment dynamics. This formulation enables us to focus on optimising the retrieval policy, with both the LLM’s behaviour and memory evolution captured within the new state–transition structure.
3.2.2 Transforming to Reflected MDP
To recover a Markovian formulation, we redefine the state to include memory, forming an augmented state $x_{t}=(s_{t},M_{t})$. The resulting process over $x_{t}$} is Markovian in this expanded space, even though the underlying environment process over $s_{t}$} is not. Each augmented state $x_{t}=(s_{t},M_{t})$ combines the environment state $s_{t}$ with the episodic memory $M_{t}$, as illustrated by the dashed rectangles in Figure 3.
In addition, since our focus is on the retrieval process rather than on modifying the LLM itself, we treat the LLM as a fixed component of the environment. Its action generation is absorbed into the environment’s transition and reward functions, effectively redefining the interaction dynamics. Under this view, the agent’s controllable decision is the retrieval action $c_{t}\sim\mu(\cdot\mid s_{t},M_{t}),$ while the environment, augmented by the fixed LLM kernel ($p_{\mathrm{LLM}}(a\mid s_{t},c_{t})$), determines both the resulting state transition and reward. This formulation leads naturally to a new MDP in which the retrieval policy ($\mu$) becomes the agent’s primary policy, and the LLM-mediated dynamics are incorporated into the environment model. Overall, we have our new MDP that is simplified by introducing an augmented state space and LLM becoming part of environment, namely,
Definition 3 (Reflected MDP).
Given an underlying SRDP
$\mathcal{M}_{\mathrm{SRDP}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{M},p_{\mathrm{LLM}}\rangle,$
with state space $\mathcal{S}$, action space $\mathcal{A}$, transition kernel $\mathcal{P}$, reward function $\mathcal{R}$, discount factor $\gamma$, episodic memory $\mathcal{M}$, and an LLM kernel $p_{\mathrm{LLM}}(a\mid s,c)$, the Reflected MDP is a tuple
$\mathcal{M}_{\mathrm{LLM}}=\langle\mathcal{X},\mathcal{C},\mathcal{P}_{\mathrm{LLM}},\mathcal{R}_{\mathrm{LLM}},\gamma\rangle,$
where the components are defined as follows:
  • Augmented state space: $\mathcal{X}=\mathcal{S}\times\mathcal{M}$, with each state $x=(s,M)$ consisting of the current environment state $s$ and the current episodic memory $M$.
  • Action space: $\mathcal{C}=M$, corresponding to retrieval actions (selecting cases from memory).
  • Transition kernel: For $x=(s,M)$ and $c\in\mathcal{C}$,
    $\mathcal{P}_{\mathrm{LLM}}(x^{\prime}\mid x,c)=\sum_{a\in\mathcal{A}}p_{\mathrm{LLM}}(a\mid s,c)\,\mathbf{1}\{x^{\prime}=(s^{\prime},\mathsf{Write}(M,s,a,\mathcal{R}(s,a)))\}\,\mathcal{P}(s^{\prime}\mid s,a),$ (3)
    where $\mathsf{Write}(M,s,a,r)$ appends the new experience $(s,a,r)$ to memory.
  • Reward function:
    $\mathcal{R}_{\mathrm{LLM}}(x,c)=\sum_{a\in\mathcal{A}}p_{\mathrm{LLM}}(a\mid s,c)\,\mathcal{R}(s,a).$ (4)
The Reflected MDP provides a fully Markovian formulation over the augmented state space $\mathcal{X}=\mathcal{S}\times\mathcal{M}$, where transitions and rewards correspond to the expected outcomes under the LLM kernel $p_{\mathrm{LLM}}$. Each transition involves four steps: (1) retrieving a case $c_{t}\sim\mu(\cdot\mid s_{t},M_{t})$ from memory, (2) generating an environment action $a_{t}\sim p_{\mathrm{LLM}}(\cdot\mid s_{t},c_{t})$, (3) sampling the next state $s_{t+1}\sim\mathcal{P}(\cdot\mid s_{t},a_{t})$, and (4) updating the memory $M_{t+1}=\mathsf{Write}(M_{t},s_{t},a_{t},\mathcal{R}(s_{t},a_{t}))$.
3.3 Policy Iteration
We now define the retrieval mechanism that underpins the learning method. The central idea is to use local state density to guide which past cases should be retrieved from memory. To capture the similarity structure in the memory bank, we employ Parzen window estimation [Parzen1962, wang2008unified].
Let $K:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ be a smooth kernel (e.g., Gaussian) with bandwidth $h>0$, and let $\psi:\mathcal{S}\to\mathbb{R}^{d}$ be a state embedding. For memory $M=\{c_{i}=(s_{i},a_{i},r_{i})\}_{i=1}^{N}$ and query state $s$, the Parzen similarity weight for case $c\in M$ is:
$w_{\text{parzen}}(s,c)=\frac{K\left(\psi(s)-\psi(s(c))/h\right)}{\sum_{c^{\prime}\in M}K\left(\psi(s)-\psi(s(c^{\prime}))/h\right)},$ (5)
where $s(c)$ denotes the state component of a case $c$. The kernel defines a probability distribution over memory cases that places higher weight on states similar to the current query state $s$. Thus, the Parzen prior $\mu_{0}(c\mid x)$ over retrieval actions at augmented state $x=(s,M)$ is:
$\mu_{0}(c\mid x)\equiv w_{\text{parzen}}(s,c).$ (6)
In order to enable an LLM agent to act even when no relevant past cases exist in memory, we augment the case set $\mathcal{C}$ with a special void case:
$\mathcal{M}_{v}\;=\;\mathcal{M}\cup\{c_{v}\}.$
The void case $c_{v}$ represents the ability of the agent to generate an action directly from its internal world knowledge, independently of retrieved memory. We formulate this by modifying the prior policy $\mu_{0}$ that incorporates a state-dependent mixture between memory-based retrieval and the void case. Let $K(x,c)$ be a similarity kernel for ordinary cases $c\in\mathcal{C}$, and let $K_{v}>0$ be a constant kernel score assigned to the void case $c_{v}$. Then, the prior distribution $\mu_{0}(\cdot\mid x)$ becomes
$\mu_{0}(c\mid x)\equiv\frac{\begin{cases}K(x,c),&c\neq c_{v},\\[3.0pt] K_{v},&c=c_{v},\end{cases}}{\sum_{c^{\prime}\neq c_{v}}K(x,c^{\prime})+K_{v}}.$ (7)
It is useful to express this distribution as a mixture. Define the normalised memory-based distribution
$\mu_{\mathrm{mem}}(c\mid x)=\frac{K(x,c)}{\sum_{c^{\prime}\neq c_{v}}K(x,c^{\prime})},\qquad c\neq c_{v},$
and let
$\lambda(x)=\frac{\sum_{c^{\prime}\neq c_{v}}K(x,c^{\prime})}{\sum_{c^{\prime}\neq c_{v}}K(x,c^{\prime})+K_{v}}.$
Then (7) is equivalent to the mixture representation
$\mu_{0}(c\mid x)=\lambda(x)\,\mu_{\mathrm{mem}}(c\mid x)+\bigl(1-\lambda(x)\bigr)\,\delta_{c_{v}}(c),$ (8)
where $\delta_{c_{v}}$ denotes the unit point mass at $c_{v}$. This mixture form shows that the agent automatically interpolates between reflection(retrieval)-driven behaviour when similarity to stored cases is high ($\lambda(x)\approx 1$) and knowledge-driven discovery through $c_{v}$ when similarity is low ($\lambda(x)\approx 0$). This mechanism mirrors the “discovery” component in infinite-armed bandits [berry1997bandit], complementing exploitation and exploration with the capacity to act beyond the current memory. This mechanism ensures that the agent can always act, even in novel states, while still exploiting relevant past experiences when available.
We aim for the agent’s final policy $\mu(\cdot\mid x)$ to remain close to a prior policy $\mu_{0}$, while still being shaped by reward feedback from the environment. To formalise this trade-off, we define a KL-regularised objective in terms of the state–case value function $Q^{\mu}(x,c)$. The KL-regularised Bellman evaluation operator is
$(\mathcal{T}^{\mu}_{\mathrm{KL}}Q)(x,c)=\mathbb{E}\!\left[r+\gamma\left(\sum_{c^{\prime}}\mu(c^{\prime}\mid x^{\prime})\,Q(x^{\prime},c^{\prime})-\alpha\,\mathrm{KL}\!\bigl(\mu(\cdot\mid x^{\prime})\,\|\,\mu_{0}(\cdot\mid x^{\prime})\bigr)\right)\,\middle|\,x,c\right].$ (9)
For any fixed $\mu$, this operator is a $\gamma$-contraction in the sup-norm and admits a unique fixed point $Q^{\mu}$, representing the KL-regularised value of $\mu$.
KL-Regularised Policy Improvement Given a $Q$-function over retrieval actions, we define policy improvement as the solution to a KL-regularised optimisation problem. Namely, at each augmented state $x$, the improved retrieval policy is:
$\mu^{+}(c\mid x)\in\arg\max_{\nu\in\Delta(M)}\left\{\sum_{c\in M}\nu(c|x)\,Q(x,c)-\alpha\,\mathrm{KL}(\nu(c|x)\|\mu_{0}(\cdot\mid x))\right\}$ (10)
where $\Delta(M)$ is the simplex over finite set, memory $M$, and $\alpha>0$ is the temperature parameter.
Lemma 4 (also in [neu2017unified]).
The optimisation problem (10) has a unique solution:
$\mu^{+}(c\mid x)=\frac{\mu_{0}(c\mid x)\,\exp(Q(x,c)/\alpha)}{\sum_{c^{\prime}\in M}\mu_{0}(c^{\prime}\mid x)\,\exp(Q(x,c^{\prime})/\alpha)}$ (11)
The proof can be found in Appendix A. Because $\mu_{0}(c_{v}\mid x)>0$ for all $x$, the improved policy $\mu^{\prime}$ always assigns nonzero probability to the void case, preserving the agent’s capacity for discovery. To summary, for fixed retrieval policy $\mu$, we define soft state-action values with the Parzen prior incorporated into the evaluation operator, and the soft $Q$-function for retrieval policy $\mu$ satisfies:
$Q^{\mu}(x,c)$ $=\mathbb{E}\left[r+\gamma V^{\mu}(x^{\prime})\mid x,c\right]$ (12)
$V^{\mu}(x)$ $=\sum_{c}\mu(c\mid x)\,Q^{\mu}(x,c)-\alpha\,\mathrm{KL}(\mu(\cdot\mid x)\|\mu_{0}(\cdot\mid x)),$ (13)
where the expectation integrates over $a\sim p_{\mathrm{LLM}}(\cdot\mid s,c)$, $s^{\prime}\sim\mathcal{P}(\cdot\mid s,a)$, and $M^{\prime}=\mathsf{Write}(M,s,a,Q)$.
KL-Regularised Policy Iteration Now, on the basis of policy evaluation in Eq. (9) and policy improvement in Eq. (11), we are ready to give the policy iteration algorithm, as shown in Algorithm 1.
Algorithm 1 Stateful Reflective Learning with Read–Write Loop
1:Input: Temperature $\alpha>0$, kernel bandwidth $h>0$, initial $Q_{0}$, initial reference prior $\mu_{0,0}$ (with void case $c_{v}$), initial memory $M_{0}$
2:for $k=0,1,2,\ldots$ do
3:  (Prior Update) Using memory $M_{k}$, construct the Parzen kernel-based reference prior $\mu_{0,k+1}(\cdot\mid x)$ according to (6), including the void case $c_{v}$ via the mixture form in (8).
4:  (Memory Read, a.k.a, Policy Improvement) Update the policy by the KL-regularised greedy step
$\mu_{k+1}(\cdot\mid x)\propto\mu_{0,k+1}(\cdot\mid x)\,\exp\bigl(Q^{\mu_{k}}(x,\cdot)/\alpha\bigr),$
i.e. using the closed-form solution in (11).
5:  (Interact with Environment) Execute the policy $(\mu_{k+1},p_{\mathrm{LLM}})$ and receive feedback $r$.
6:  (Memory Rewrite, a.k.a., Policy Evaluation) Collect transitions $(s,a,r)$, and update memory
$M_{k+1}\leftarrow\mathsf{Write}(M_{k},s,a,r).$
7:   Given $\mu_{k}$ and $\mu_{0,k}$, approximately solve $Q^{\mu_{k}}$ by iterating
$Q\leftarrow\mathcal{T}^{\mu_{k}}_{\mathrm{KL}}Q$
using the KL-regularised evaluation operator in (9). Two classes of memory representations: (i) a tabular memory that explicitly stores $Q$-values for each state–action tuple, and (ii) a parameterised memory in which the $Q$-function is represented by a neural network.
8:end for
Under standard assumptions, this procedure yields a monotone improvement in the KL-regularised value and converges to the fixed point of the KL-regularised optimality operator,
$(\mathcal{T}^{*}_{\mathrm{KL}}Q)(x,c)=\mathbb{E}\!\left[r+\gamma\,\alpha\log\!\sum_{c^{\prime}}\mu_{0}(c^{\prime}\mid x^{\prime})\,\exp\!\bigl(Q(x^{\prime},c^{\prime})/\alpha\bigr)\,\middle|\,x,c\right].$
Specifically, we define
Assumption 5 (Bounded rewards and discount).
$|r|\leq R_{\max}<\infty$, $\gamma\in[0,1)$.
Assumption 6 (Stationary memory dynamics during evaluation).
Within each policy-evaluation phase, the memory kernel is stationary (equivalently, treat $M$ as part of the state).
Theorem 7 (Convergence of Parzen-KL Soft Policy Iteration).
Under assumptions 5 and 6 (bounded rewards and $\gamma<1$ and stationary memory), the process (the reflected MDP) is an ordinary discounted MDP. Algorithm 1 converges to a fixed point $(Q^{\star},\mu^{\star})$ that is optimal for the KL-regularised objective with Parzen prior.
We refer to Appendix for the proof.
Discussion It is worth mentioning that the void case $c_{v}$ plays a crucial role: by providing a state-independent baseline probability in $\mu_{0}$, it ensures that policy iteration retains the ability to discover new behaviours beyond exploitation or exploration of existing memory. Unlike classical density estimation, our memory cases are not independent and identically distributed but are gathered sequentially through policy execution. Moreover, the finite memory constraint means we perform kernel smoothing over a limited and evolving set of points rather than relying on asymptotic density estimation. As a result, bandwidth selection plays a critical role in balancing local similarity against statistical robustness as the memory continues to grow. Several strategies can be used to choose the bandwidth $h$. Silverman’s rule suggests setting $h\propto n^{-1/(d+4)}$, where $n=|M|$ is the memory size and $d$ the embedding dimension. Another approach is cross-validation, which tunes $h$ by optimising performance on held-out retrieval tasks. Finally, an adaptive strategy decreases $h$ as memory grows, ensuring that neighbourhoods remain local and informative.
3.4 Two-Time-Scale Convergence with Online Memory Re-writing
We next allow memory to evolve on a slower time-scale. Let $\eta_{t}$ be steps for evaluation/improvement and $\rho_{t}$ for memory updates, with $\rho_{t}/\eta_{t}\to 0$, $\sum_{t}\eta_{t}=\infty$, $\sum_{t}\eta_{t}^{2}<\infty$. We make the following assumptions.
Assumption 8 (Two-Time-Scale Conditions).
  1. (i)
    Step sizes: $\eta_{t}>0$, $\rho_{t}>0$ with:
    $\sum_{t=0}^{\infty}\eta_{t}=\infty,\quad\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty,\quad\sum_{t=0}^{\infty}\rho_{t}=\infty,\quad\sum_{t=0}^{\infty}\rho_{t}^{2}<\infty,\quad\lim_{t\to\infty}\frac{\rho_{t}}{\eta_{t}}=0$
  2. (ii)
    Martingale difference noise: $\xi^{(Q)}_{t}$ and $\xi^{(\mu)}_{t}$ are martingale difference sequences with respect to the filtration $\mathcal{F}_{t}$, information available up to time $t$ (random variable with zero conditional mean) :
    $\mathbb{E}[\xi^{(Q)}_{t}|\mathcal{F}_{t}]=0,\quad\mathbb{E}[\xi^{(\mu)}_{t}|\mathcal{F}_{t}]=0,\quad\mathbb{E}[(\xi^{(Q)}_{t})^{2}|\mathcal{F}_{t}]\leq K,\quad\mathbb{E}[(\xi^{(\mu)}_{t})^{2}|\mathcal{F}_{t}]\leq K$
  3. (iii)
    Bounded iterates: The sequences $\{Q_{t}\}$, $\{\mu_{t}\}$, $\{M_{t}\}$ remain bounded almost surely.
  4. (iv)
    Memory attractor: The memory process $\{M_{t}\}$ has a compact attractor set $\mathcal{M}_{\infty}\subset\mathcal{M}$. This means memory size is bounded and the content distribution stabilises over time. In practice, this is achieved by: 1) Using fixed-size memory, 2) Having policies that converge to stationary distributions, 3) Using environments that produce stationary experience streams, or 4) Conservative memory update strategies.
Theorem 9 (Two-Time-Scale Convergence for Parzen-KL Policy Iteration).
Under Assumptions 5 and 8, let $(Q_{t},\mu_{t},M_{t})$ be generated by the following sample-based recursions.
  1. 1.
    Q-update (TD learning, sample form):
    $Q_{t+1}(x_{t},c_{t})$ $=Q_{t}(x_{t},c_{t})+\eta_{t}\Big[\delta_{t}(Q_{t},\mu_{0},M_{t+1})\Big],$ (14)
    $\delta_{t}(Q,\mu_{0},M)$ $:=r_{t}+\gamma\alpha\log\!\Big(\sum_{c^{\prime}\in M}\mu_{0}(c^{\prime}|x_{t+1})e^{Q(x_{t+1},c^{\prime})/\alpha}\Big)-Q(x_{t},c_{t}).$
  2. 2.
    Policy update (Parzen-KL improvement, sample form):
    $\mu_{t+1}(c|x_{t})$ $=\mu_{t}(c|x_{t})+\eta_{t}\Big[\Phi_{t}(c|x_{t};Q_{t},\mu_{0},M_{t})-\mu_{t}(c|x_{t})\Big],$ (15)
    $\Phi_{t}(c|x;Q,\mu_{0},M)$ $:=\frac{\mu_{0}(c|x)\,e^{Q(x,c)/\alpha}}{Z_{t}(x)},\qquad Z_{t}(x):=\sum_{c^{\prime}\in M}\mu_{0}(c^{\prime}|x)\,e^{Q(x,c^{\prime})/\alpha}.$
  3. 3.
    Memory update (experience replay):
    $M_{t+1}=\begin{cases}M_{t}\cup\{(s_{t},a_{t},r_{t})\}&\text{with probability }\rho_{t},\\ M_{t}&\text{otherwise.}\end{cases}$ (16)
Define the filtration $\mathcal{F}_{t}:=\sigma\big((x_{k},c_{k},r_{k}),M_{k},Q_{k},\mu_{k}:k\leq t\big)$, which collects all information available to the agent up to time (t), including past observations, the current memory state, and the value function and policy parameters. Define the mean drifts (expected operators) by
$H_{Q}(Q_{t},\mu_{t},M_{t})(x_{t},c_{t})$ $:=\mathbb{E}\!\left[\,\delta_{t}(Q_{t},\mu_{0},M_{t+1})\mid\mathcal{F}_{t}\right],$ (17)
$H_{\mu}(Q_{t},\mu_{t},M_{t})(c|x_{t})$ $:=\mathbb{E}\!\left[\,\Phi_{t}(c|x_{t};Q_{t},\mu_{0},M_{t})-\mu_{t}(c|x_{t})\mid\mathcal{F}_{t}\right].$ (18)
Then define the martingale difference noises as the residuals
$\xi^{(Q)}_{t+1}$ $:=\delta_{t}(Q_{t},\mu_{0},M_{t+1})-H_{Q}(Q_{t},\mu_{t},M_{t})(x_{t},c_{t}),$ (19)
$\xi^{(\mu)}_{t+1}(c|x_{t})$ $:=\Big(\Phi_{t}(c|x_{t};Q_{t},\mu_{0},M_{t})-\mu_{t}(c|x_{t})\Big)-H_{\mu}(Q_{t},\mu_{t},M_{t})(c|x_{t}).$ (20)
With these definitions, the recursions can be written in stochastic approximation form:
$Q_{t+1}(x_{t},c_{t})$ $=Q_{t}(x_{t},c_{t})+\eta_{t}\Big[H_{Q}(Q_{t},\mu_{t},M_{t})(x_{t},c_{t})+\xi^{(Q)}_{t+1}\Big$], (21)
$\mu_{t+1}(c|x_{t})$ $=\mu_{t}(c|x_{t})+\eta_{t}\Big[H_{\mu}(Q_{t},\mu_{t},M_{t})(c|x_{t})+\xi^{(\mu)}_{t+1}(c|x_{t})\Big].$ (22)
Moreover, $(\xi^{(Q)}_{t+1})$ and $(\xi^{(\mu)}_{t+1})$ are martingale difference sequences w.r.t. $(\mathcal{F}_{t})$:
$\mathbb{E}\!\left[\xi^{(Q)}_{t+1}\mid\mathcal{F}_{t}\right]=0,\qquad\mathbb{E}\!\left[\xi^{(\mu)}_{t+1}(c|x_{t})\mid\mathcal{F}_{t}\right]=0,$
and satisfy the conditional second-moment bounds in Assumption 8 ii.
Finally, the conclusions (a)–(c) hold:
  1. (a)
    The fast variables $(Q_{t},\mu_{t})$ track the stationary-memory soft policy-iteration fixed points:
    $\limsup_{t\to\infty}\inf_{(Q^{*}_{M},\mu^{*}_{M})\in\mathcal{F}(M_{t})}\|(Q_{t},\mu_{t})-(Q^{*}_{M},\mu^{*}_{M})\|\leq\epsilon\quad\text{a.s.}$
  2. (b)
    The joint process converges to the limiting set:
    $(Q_{t},\mu_{t},M_{t})\to\{(Q^{*}_{M},\mu^{*}_{M},M):M\in\mathcal{M}_{\infty},\ (Q^{*}_{M},\mu^{*}_{M})\in\mathcal{F}(M)\}.$
  3. (c)
    If $M_{t}\to M_{\infty}$ a.s., then $(Q_{t},\mu_{t})\to(Q^{*}_{M_{\infty}},\mu^{*}_{M_{\infty}})$.
The above result can be established using the ODE method for stochastic approximation [borkar2009stochastic]. The key insight is that from the perspective of the fast variables $(Q_{t},\mu_{t})$, the slow variable $M_{t}$ appears quasi-static due to the timescale separation $\rho_{t}/\eta_{t}\to 0$. This detailed proof establishes several key properties of our Read–Write Learning iteration. First, it demonstrates robustness to memory updates, showing that the algorithm remains stable even as the memory evolves, provided that memory updates occur on a slower timescale than value and policy updates. Second, it guarantees tracking behaviour, ensuring that the value function and policy rapidly adjust to reflect changes in memory, remaining close to the optimal configuration for the current memory state. Third, the analysis confirms convergence to local optima, proving that for each fixed memory configuration (M), the algorithm converges to the corresponding optimal retrieval policy ($\mu^{*}_{M}$). Finally, the timescale separation result provides practical implementation guidance, indicating that memory should be updated less frequently than the value function and policy. Collectively, these results provide strong theoretical guarantees for continual Read–Write Learning loop under our SRDP, demonstrating that it can adapt effectively to evolving memory while preserving convergence and stability.
3.5 The Role of Reflection and Memory towards Optimal Policy
A key question now is whether the Read–Write Learning agent can approach optimal performance in the underlying MDP as memory grows. We now establish reflection consistency: the property that increasing memory coverage drives the composite policy toward the true optimal policy.
[…]
Figure 4: The illustration of the assumption of LLM local sufficiency.
Algorithm 2 Memory-Based Policy Iteration via Parzen–KL Q-Learning
1:Input: Learning rate $\eta_{t}$, temperature $\alpha$, bandwidth $h$
2:Initialise: Episodic memory $M_{0}$ and Q-function $Q_{0}$
3:
4:(1) Read: Retrieve relevant experiences
5:Given current augmented state $x_{t}=(s_{t},M_{t})$, sample a case $c_{t}\sim\mu(\cdot\mid x_{t})$ from memory.
6:(2) Evaluate: Compute soft target using Parzen–KL smoothing
7:Observe reward $r_{t}$ and next augmented state $x_{t+1}=(s_{t+1},M_{t+1})$
8:Estimate Parzen prior $\mu_{0}(\cdot\mid x_{t+1})$ using (6)
9:Compute the soft target:
$y_{t}=r_{t}+\gamma\alpha\log\sum_{c^{\prime}\in M_{t+1}}\mu_{0}(c^{\prime}\mid x_{t+1})\exp\!\big(Q_{t}(x_{t+1},c^{\prime})/\alpha\big)$
10:(3) Write: Update memory and value function
11:Update Q-value:
$Q_{t+1}(x_{t},c_{t})=(1-\eta_{t})Q_{t}(x_{t},c_{t})+\eta_{t}y_{t}$
12:Store the new experience $(s_{t},c_{t},r_{t},s_{t+1})$ into memory:
$M_{t+1}=\mathsf{Write}(M_{t},s_{t},c_{t},r_{t})$
13:
14:Output: Updated memory $M_{t+1}$ and Q-function $Q_{t+1}$
For each memory $M$, define the coverage radius $r_{M}\triangleq\sup_{s\in\mathcal{S}}\min_{c\in M}d\big(s,s(c)\big)$. Let the (deterministic) retrieval rule $c_{M}(s)$ pick any case achieving the minimum. The LLM-induced composite policy on the original state space is
$\pi_{M}(a\mid s)\;\triangleq\;p_{\mathrm{LLM}}\big(a\mid s,c_{M}(s)\big).$
Assume bounded rewards $|\mathcal{R}(s,a)|\leq R_{\max}$ and $\gamma\in[0,1)$.
Assumption 10 (LLM local consistency).
There exists a modulus $\varepsilon_{\mathrm{LLM}}(r)\rightarrow 0$ as $r\rightarrow 0$ such that for any $s\in\mathcal{S}$ and any case $c$ with $d\big(s,s(c)\big)\leq r$,
$\mathrm{TV}\!\left(p_{\mathrm{LLM}}(\cdot\mid s,c),\,\pi^{*}(\cdot\mid s)\right)\;\leq\;\varepsilon_{\mathrm{LLM}}(r),$
where $\mathrm{TV}(p,q)$ denotes the total variation distance:
$\mathrm{TV}(p,q)=\tfrac{1}{2}\sum_{a\in\mathcal{A}}\left|p(a)-q(a)\right|\quad\text{(discrete case)},$
or
$\mathrm{TV}(p,q)=\tfrac{1}{2}\int_{\mathcal{A}}\left|p(a)-q(a)\right|\,da\quad\text{(continuous case)}.$
Lemma 11 (Policy Approximation at State Level).
Under Assumption 10:
$\mathrm{TV}(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s))\leq\varepsilon_{\mathrm{LLM}}(r_{M}(s))+\delta_{M}(s)$
Proof.
Let $\mathcal{E}_{s}$ denote the event that the retriever selects a good case within distance $r_{M}(s)$ of state $s$. By construction, this event occurs with probability at least $1-\delta_{M}(s)$. Conditioned on $\mathcal{E}_{s}$, Assumption 10 implies
$\mathrm{TV}\!\left(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s)\right)\leq\varepsilon_{\mathrm{LLM}}\!\left(r_{M}(s)\right).$
With the complementary event $\mathcal{E}_{s}^{c}$, the total variation distance is trivially bounded by
$\mathrm{TV}\!\left(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s)\right)\leq 1.$
Taking expectation over the retrieval event yields
$\mathrm{TV}\!\left(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s)\right)$ $\leq(1-\delta_{M}(s))\,\varepsilon_{\mathrm{LLM}}\!\left(r_{M}(s)\right)+\delta_{M}(s)$
$\leq\varepsilon_{\mathrm{LLM}}\!\left(r_{M}(s)\right)+\delta_{M}(s),$
which completes the proof. ∎
Lemma 12 (Immediate reward approximation).
For any $s\in\mathcal{S}$, with $c^{*}=c_{M}(s)$,
$\bigg|\;\mathcal{R}_{\mathrm{LLM}}\big((s,M),c^{*}\big)\;-\;\sum_{a}\pi^{*}(a\mid s)\,\mathcal{R}(s,a)\;\bigg|\;\leq\;2R_{\max}\,\mathrm{TV}\!\Big(p_{\mathrm{LLM}}(\cdot\mid s,c^{*}),\,\pi^{*}(\cdot\mid s)\Big).$
Proof.
Write $\mathcal{R}_{\mathrm{LLM}}\!=\!\sum_{a}p_{\mathrm{LLM}}(a\mid s,c^{*})\,\mathcal{R}(s,a)$ and $R^{*}(s)\!=\!\sum_{a}\pi^{*}(a\mid s)\,\mathcal{R}(s,a)$. By Hölder’s inequality and the definition of total variation, $\big|\sum_{a}(p-\pi^{*})\mathcal{R}\big|\leq\|\mathcal{R}\|_{\infty}\|p-\pi^{*}\|_{1}\leq 2R_{\max}\,\mathrm{TV}(p,\pi^{*})$. ∎
Theorem 13 (Value difference bound via policy TV).
Let $\pi^{*}$ be the optimal for the original MDP, and let $\pi_{M}$ be the composite policy defined in SRDP. Then
$\big\|V^{\pi^{*}}-V^{\pi_{M}}\big\|_{\infty}\;\leq\;\frac{2R_{\max}}{(1-\gamma)^{2}}\;\sup_{s\in\mathcal{S}}\mathrm{TV}\!\big(\pi_{M}(\cdot\mid s),\,\pi^{*}(\cdot\mid s)\big)\;=\;\frac{2R_{\max}}{(1-\gamma)^{2}}\,\Delta_{M}.$
Proof.
Consider the Bellman operators $(T^{\pi}V)(s)=\sum_{a}\pi(a\mid s)\big(\mathcal{R}(s,a)+\gamma\,\mathbb{E}[V(s^{\prime})\mid s,a]\big)$. Since $V^{\pi^{*}}$ satisfies $V^{\pi^{*}}=T^{\pi^{*}}V^{\pi^{*}}$,
$\|V^{\pi^{*}}-V^{\pi_{M}}\|_{\infty}\;\leq\;\frac{1}{1-\gamma}\,\big\|(T^{\pi^{*}}-T^{\pi_{M}})V^{\pi^{*}}\big\|_{\infty}.$
For any $s$,
$\big|(T^{\pi^{*}}-T^{\pi_{M}})V^{\pi^{*}}(s)\big|$
$=\Big|\sum_{a}\big(\pi^{*}-\pi_{M}\big)(a\mid s)\,\mathcal{R}(s,a)\;+\;\gamma\,\sum_{a}\big(\pi^{*}-\pi_{M}\big)(a\mid s)\,\mathbb{E}[V^{\pi^{*}}(s^{\prime})\mid s,a]\Big|$
$\leq 2R_{\max}\,\mathrm{TV}(\pi^{*},\pi_{M})\;+\;\gamma\,\|V^{\pi^{*}}\|_{\infty}\cdot 2\,\mathrm{TV}(\pi^{*},\pi_{M}),$
using Lemma 12’s argument and that $|\mathbb{E}V^{\pi^{*}}|\leq\|V^{\pi^{*}}\|_{\infty}$. Since $\|V^{\pi^{*}}\|_{\infty}\leq R_{\max}/(1-\gamma)$,
$\big\|(T^{\pi^{*}}-T^{\pi_{M}})V^{\pi^{*}}\big\|_{\infty}\;\leq\;2R_{\max}\,\mathrm{TV}(\pi^{*},\pi_{M})\Big(1+\tfrac{\gamma}{1-\gamma}\Big)\;=\;\frac{2R_{\max}}{1-\gamma}\,\mathrm{TV}(\pi^{*},\pi_{M}).$
Divide by $(1-\gamma)$ to get the claimed bound. ∎
Corollary 14 (Asymptotic optimality with growing memory).
Under Assumption 10 and Lemma 11, define
$\delta_{M}\;\triangleq\;\sup_{s\in\mathcal{S}}\delta_{M}(s).$
Then
$\Delta_{M}\;\triangleq\;\sup_{s\in\mathcal{S}}\mathrm{TV}\!\big(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s)\big)\;\leq\;\varepsilon_{\mathrm{LLM}}(r_{M})+\delta_{M}.$
If $r_{M}\to 0$ and $\delta_{M}\to 0$, then $\Delta_{M}\to 0$ and
$\sup_{s\in\mathcal{S}}\big|V^{\pi^{\star}}(s)-V^{\pi_{M}}(s)\big|\;\leq\;\frac{2R_{\max}}{(1-\gamma)^{2}}\,\Delta_{M}\;\longrightarrow\;0.$
Hence, as memory grows, the Read–Write operation drives the agent toward optimal performance in the underlying MDP.
Proof.
Lemma 11 implies $\mathrm{TV}(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s))\leq\varepsilon_{\mathrm{LLM}}(r_{M})+\delta_{M}(s)$ for all $s\in\mathcal{S}$. Taking the supremum over $s$ yields $\Delta_{M}\leq\varepsilon_{\mathrm{LLM}}(r_{M})+\delta_{M}\to 0$. Applying Theorem 13 completes the proof. ∎
4 Read–Write Learning Equivalency
The following theorem shows that iterative memory operations, consisting of the read and write processes, are equivalent to performing policy iteration, which includes policy evaluation and policy improvement. This demonstrates that memory interaction itself realises a learning mechanism, and by the previous theorem, the process converges to the optimal policy in the underlying MDP.
Theorem 15 (Equivalence of Memory and Policy Operations).
The following processes are equivalent:
  1. 1.
    Retrieval Policy Iteration in SRDP:
    • Memory Read: $c\sim\mu(\cdot|x)$
    • Memory Write: $M^{\prime}=\mathsf{Write}(M,s,a,r)$
    • LLM Action: $a\sim p_{\mathrm{LLM}}(\cdot|s,c)$
  2. 2.
    Standard Policy Iteration in Reflected MDP $\widetilde{\mathcal{M}}$:
    • Policy Evaluation: Compute $Q^{\mu}(x,c)$
    • Policy Improvement: Update $\mu$ to be greedy/softmax w.r.t. $Q^{\mu}$
Specifically:
  1. (a)
    The value functions are equal: $V^{\pi_{M}}(s)=\widetilde{V}^{\mu}(x)$
  2. (b)
    Optimal policies correspond: $\pi_{M}^{\star}$ optimal in M-MDP $\iff$ $\mu^{\star}$ optimal in $\widetilde{\mathcal{M}}$
  3. (c)
    Memory write implements policy evaluation data collection
  4. (d)
    Memory read implements policy improvement
Proof.
(a) Value Function Equality:
$V^{\pi_{M}}(s)$ $=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})\middle|s_{0}=s,a_{t}\sim p_{\mathrm{LLM}}(\cdot|s_{t},c_{t}),c_{t}\sim\mu(\cdot|x_{t})\right]$
$=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\widetilde{\mathcal{R}}(x_{t},c_{t})\middle|x_{0}=(s,M_{0}),c_{t}\sim\mu(\cdot|x_{t})\right]$
$=\widetilde{V}^{\mu}(x)$
(b) Policy Optimality: Follows directly from (a) since the objectives are identical.
(c) Memory Write as Evaluation: Storing $(s,a,r,s^{\prime})$ provides the data $(x,c,\widetilde{\mathcal{R}}(x,c),x^{\prime})$ needed for temporal difference learning.
(d) Memory Read as Improvement: Selecting $c$ to maximise expected return is exactly policy improvement in $\widetilde{\mathcal{M}}$. ∎
Therefore, in practical implementations, memory operations can be used as a mechanism for continuously learning and improving the policy over time, as illustrated in Algorithm 2.
5 Related Work
Our work on Memory-Augmented Markov Decision Processes (LLM-M-MDPs) with Parzen-KL retrieval builds upon, yet also departs from, several established lines of research in reinforcement learning and large language models (LLMs). This section situates our contribution within the broader landscape of memory-based RL and clarifies the key distinctions from existing approaches.
The integration of memory mechanisms into reinforcement learning has long been explored to address the challenges of partial observability and long-term credit assignment. Early approaches often relied on recurrent neural networks (RNNs), in particular Long Short-Term Memory (LSTM) networks [hochreiter1997long], to maintain hidden states that capture historical information. While effective in certain tasks, these methods frequently struggle with long-term dependencies and lack explicit mechanisms for memory management [morad2023memory].
More recent methods have taken inspiration from cognitive neuroscience, especially the hippocampal processes of memory formation and retrieval in humans and animals. For instance, the Hippocampal-Augmented Memory Integration (HAMI) framework [hami2025] incorporates symbolic indexing and hierarchical refinement based on hippocampal functions. Although HAMI shares our biological motivation, our Parzen-KL retrieval mechanism diverges in its probabilistic formulation and information-theoretic regularisation, which underpin both stability and scalability.
Episodic control methods form another strand of memory-based RL. Model-Free Episodic Control (MFEC) [blundell2016model] and Neural Episodic Control (NEC) [pritzel2017neural] explicitly store past experiences and retrieve them via similarity measures. These methods, however, generally lack principled retrieval strategies and encounter difficulties in scaling to large memory sizes [ramani2019short]. Our framework addresses these issues by introducing density-based similarity priors and KL-regularised updates, which enable efficient and stable use of memory over time.
Case-based reasoning (CBR) [AamodtPlaza1994, Ashley1990] has also inspired memory-driven approaches to problem solving, emphasising the retrieval and adaptation of relevant past cases. In modern contexts, LLM agents increasingly employ explicit episodic memory systems to extend context and improve performance across diverse tasks [Zhong2024, Chhikara2025, fountas2024human]. Yet, the integration of memory retrieval with decision-making in LLMs still lacks a firm theoretical foundation, particularly regarding the conditions under which such systems converge to optimal behaviour. Our work addresses this gap directly.
Experience replay, a widely used mechanism in deep RL since its adoption in DQN [mnih2015human], has evolved from uniform sampling to prioritised replay [schaul2015prioritized], where transitions are sampled according to temporal-difference errors. While these methods improve sample efficiency, they do not provide the structured, similarity-based retrieval required for generalisation across diverse contexts. Our approach diverges here by employing density-aware retrieval via Parzen window estimation, offering more nuanced and context-sensitive access to stored experiences.
Recent work on Stable Hadamard Memory [hadamard2024] highlights the importance of stability in memory systems, using Hadamard products to calibrate and update stored information. While complementary in spirit, our method focuses instead on retrieval optimisation through information-theoretic regularisation. These two perspectives—stability of storage and optimality of retrieval—could be integrated in future research to build more comprehensive memory architectures.
Our work also connects to broader developments in information-theoretic reinforcement learning. Maximum entropy RL [haarnoja2018soft] introduces entropy regularisation to encourage exploration and robustness, while KL constraints in algorithms such as TRPO [schulman2015trust] and MPO [abdolmaleki2018maximum] ensure stable policy updates. The distinctive aspect of our contribution is the application of KL regularisation not only to the policy but also to the retrieval process itself. This dual-level regularisation balances the exploitation of high-value memories with the exploration of diverse experiences, whilst preserving consistency with the underlying similarity structure.
Finally, from a theoretical perspective, our framework extends established convergence results in reinforcement learning. The convergence proof for our soft policy iteration builds upon entropy-regularised MDPs [neu2017unified] but adapts them to the context of memory-based retrieval. Our two-time-scale analysis further draws on stochastic approximation theory [borkar2009stochastic], extending it to LLM-M-MDPs where value and memory are updated simultaneously. A central outcome is the formal equivalence we establish between retrieval, read, and write operations and the fundamental processes of policy improvement and policy evaluation. This connection makes retrieval-augmented LLMs provably learnable for the first time, opening the way for reinforcement learning theory and tools to be systematically applied in this emerging setting.
6 Conclusions
Our Parzen-KL framework makes several theoretical contributions. First, it introduces a principled notion of similarity, using Parzen windows to provide a statistically grounded approach to similarity-based retrieval. Second, it offers convergence guarantees, showing that both the soft and hard variants of the method converge under appropriate conditions. Third, it establishes reflection consistency, formally linking memory growth to progress toward optimality. Finally, it outlines practical algorithms, presenting clear implementation paths for both exact and approximate variants.
Beyond the theoretical advances, the analysis highlights several practical implications for the design of memory-augmented LLM agents. Effective performance requires density-aware retrieval strategies to balance exploration and exploitation. The void case $c_{0}$ controls the discovery from the LLM itself whose parameter is required to be tuned. Stability can be enhanced by implementing soft, entropy-regularised improvements, while systematic memory growth helps expand state coverage. In addition, similarity metrics and bandwidths should adapt as the memory evolves to maintain efficiency and relevance.
Despite these strengths, the framework has certain limitations. It might face computational challenges as memory size increases. Moreover, the impact of embedding quality on performance is not fully addressed.
Looking ahead, there are several promising directions for future work. These include extending the analysis to continuous state spaces, developing more advanced strategies for memory curation, and conducting large-scale empirical validation on complex domains. Together, these efforts could further strengthen the role of reflection and memory in advancing memory-augmented LLM agents.
7 Acknowledgements
We are grateful to Xue Yuan for valuable discussions on reinforcement learning and kernel methods, Siyuan for insights into bandit algorithms and the discovery–exploration–exploitation trade-off, and Huichi and Yihang for assistance with implementation. We also thank collaborators who reviewed earlier versions of this manuscript and provided helpful feedback.
Appendix A Proof of Lemma 4
Proof.
The Lagrangian for (10) with the constraint $\sum_{c}\nu(c)=1$ is
$L(\nu,\lambda)=\sum_{c}\nu(c)Q(x,c)\;-\;\alpha\sum_{c}\nu(c)\log\frac{\nu(c)}{\mu_{0}(c\mid x)}\;-\;\lambda\Big(\sum_{c}\nu(c)-1\Big).$
Taking the partial derivative w.r.t. $\nu(c)$ and setting it to zero, we use
$\frac{\partial}{\partial\nu}\Big[\nu\log(\nu/\mu_{0})\Big]=\log(\nu/\mu_{0})+1,$
to obtain
$\frac{\partial L}{\partial\nu(c)}=Q(x,c)-\alpha\Big(\log\frac{\nu(c)}{\mu_{0}(c\mid x)}+1\Big)-\lambda=0.$
Equivalently,
$\log\frac{\nu(c)}{\mu_{0}(c\mid x)}=\frac{Q(x,c)-\lambda}{\alpha}-1.$
Therefore,
$\nu(c)=\mu_{0}(c\mid x)\exp\!\Big(\tfrac{1}{\alpha}Q(x,c)\Big)\exp\!\Big(-1-\tfrac{\lambda}{\alpha}\Big).$
Imposing $\sum_{c}\nu(c)=1$ gives the normalising constant
$\exp\!\Big(-1-\tfrac{\lambda}{\alpha}\Big)=\left(\sum_{c}\mu_{0}(c\mid x)\exp\!\Big(\tfrac{1}{\alpha}Q(x,c)\Big)\right)^{-1}.$
Substituting back yields (11). ∎
Appendix B Proof of Theorem 7
Unlike standard soft reinforcement learning [Haarnoja2018], our setting involves state-dependent base measures that vary with memory content, requiring a careful analysis of the induced operators. In particular, extending KL-regularised evaluation requires establishing that the associated log-sum-exp operations preserve contraction for both the value function and the prior distributed resulted from memory updates. Lemma 16 extends the log-sum-exp Lipschitz bound [BoydVandenberghe2004] to this setting with Parzen base measures and memory updates.
B.1 Bi-Lipschitz Properties
Lemma 16 (Bi-Lipschitz property of log-sum-exp with base measure).
Let $q,q^{\prime}\in\mathbb{R}^{n}$, $b,b^{\prime}\in\Delta_{n}$, and $\alpha>0$. Define
$f(b,q)\;\triangleq\;\alpha\log\sum_{i=1}^{n}b_{i}\exp(q_{i}/\alpha).$
Then:
  1. 1.
    (Lipschitz in $q$) For any fixed $b\in\Delta_{n}$,
    $|f(b,q)-f(b,q^{\prime})|\leq\|q-q^{\prime}\|_{\infty}.$
  2. 2.
    (Lipschitz in $b$) If $q_{i}\in[Q_{\min},Q_{\max}]$ for all $i$ and $\Delta_{Q}=Q_{\max}-Q_{\min}$, then for any fixed $q$,
    $|f(b,q)-f(b^{\prime},q)|\leq\alpha e^{\Delta_{Q}/\alpha}\,\|b-b^{\prime}\|_{1}=2\alpha e^{\Delta_{Q}/\alpha}\,\mathrm{TV}(b,b^{\prime}).$
  3. 3.
    (Joint bound) Under the same boundedness condition,
    $|f(b,q)-f(b^{\prime},q^{\prime})|\leq\|q-q^{\prime}\|_{\infty}+\alpha e^{\Delta_{Q}/\alpha}\,\|b-b^{\prime}\|_{1}.$
Proof.
(1) Lipschitz in $q$ (base-measure log-sum-exp). Fix any $b\in\Delta_{n}$ and define the single-argument function
$f(q)\;\triangleq\;f(b,q)=\alpha\log\sum_{i=1}^{n}b_{i}\exp\!\left(\frac{q_{i}}{\alpha}\right).$
The function $f$ is differentiable with gradient
$\nabla f(q)=\left[\frac{b_{i}\exp(q_{i}/\alpha)}{\sum_{j=1}^{n}b_{j}\exp(q_{j}/\alpha)}\right]_{i=1}^{n},$
which is a probability vector (all components are nonnegative and sum to one), hence
$\|\nabla f(q)\|_{1}=1\qquad\text{for all }q\in\mathbb{R}^{n}.$
By the mean value theorem for vector-valued functions (e.g., [BoydVandenberghe2004]), there exists $\xi$ on the line segment between $q$ and $q^{\prime}$ such that
$f(q)-f(q^{\prime})=\nabla f(\xi)^{\top}(q-q^{\prime}).$
Therefore, by Hölder’s inequality,
$|f(b,q)-f(b,q^{\prime})|=|f(q)-f(q^{\prime})|\leq\|\nabla f(\xi)\|_{1}\,\|q-q^{\prime}\|_{\infty}=\|q-q^{\prime}\|_{\infty}.$
(2) Lipschitz in $b$. Fix any $q$ and write $f(b,q)=\alpha\log(b^{\top}w)$ with $w_{i}=e^{q_{i}/\alpha}$. Then
$\nabla_{b}f(b,q)=\alpha\,\frac{w}{b^{\top}w}.$
If $q_{i}\in[Q_{\min},Q_{\max}]$ for all $i$, then
$\min_{i}w_{i}=e^{Q_{\min}/\alpha},\qquad\max_{i}w_{i}=e^{Q_{\max}/\alpha},\qquad b^{\top}w\geq\min_{i}w_{i}=e^{Q_{\min}/\alpha}.$
Hence
$\|\nabla_{b}f(b,q)\|_{\infty}\leq\alpha\,\frac{\max_{i}w_{i}}{\min_{i}w_{i}}=\alpha e^{(Q_{\max}-Q_{\min})/\alpha}=\alpha e^{\Delta_{Q}/\alpha}.$
Applying the mean value theorem in $b$ and $\ell_{1}\ell_{\infty}$ duality gives
$|f(b,q)-f(b^{\prime},q)|\leq\|\nabla_{b}f(\xi,q)\|_{\infty}\,\|b-b^{\prime}\|_{1}\leq\alpha e^{\Delta_{Q}/\alpha}\,\|b-b^{\prime}\|_{1}=2\alpha e^{\Delta_{Q}/\alpha}\,\mathrm{TV}(b,b^{\prime}).$
(3) Joint bound. Add and subtract $f(b,q^{\prime})$ and apply (1)–(2):
$|f(b,q)-f(b^{\prime},q^{\prime})|\leq|f(b,q)-f(b,q^{\prime})|+|f(b,q^{\prime})-f(b^{\prime},q^{\prime})|\leq\|q-q^{\prime}\|_{\infty}+\alpha e^{\Delta_{Q}/\alpha}\,\|b-b^{\prime}\|_{1}.$
Next, we use the above to establish the contraction for the KL-regularised evaluation.
B.2 Contraction of the KL-Regularised Evaluation Operator
We prove Theorem 17 using a single unified Lipschitz lemma for log-sum-exp with a base measure (Lemma 16), which simultaneously provides (i) Lipschitz continuity in the value argument $Q$ and (ii) stability with respect to perturbations of the base measure induced by memory.
Theorem 17 (Contraction of KL-Regularised Evaluation).
Fix a memory state $M$ and a base measure $\mu_{0}^{M}(\cdot\mid x)$ for each context $x$. Define the KL-regularised evaluation operator $\mathcal{T}^{\mu,M}_{\mathrm{KL}}$ acting on $Q:\mathcal{X}\times\mathcal{C}\to\mathbb{R}$ by
$(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q)(x,c)\;\triangleq\;r(x,c)+\gamma\,\alpha\log\!\Big(\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x^{\prime})\exp(Q(x^{\prime},c^{\prime})/\alpha)\Big),$ (23)
where $x^{\prime}$ denotes the next context/state (possibly random) following $(x,c)$, and $\gamma\in(0,1)$ is the discount factor. Then $\mathcal{T}^{\mu,M}_{\mathrm{KL}}$ is a $\gamma$-contraction in the $\|\cdot\|_{\infty}$ norm:
$\|\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q-\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q^{\prime}\|_{\infty}\;\leq\;\gamma\|Q-Q^{\prime}\|_{\infty}.$
Consequently, $\mathcal{T}^{\mu,M}_{\mathrm{KL}}$ admits a unique fixed point $Q_{M}^{\star}$.
Proof.
Fix $M$ and let $Q,Q^{\prime}$ be arbitrary bounded functions. For any $(x,c)$, subtract (23) evaluated at $Q$ and $Q^{\prime}$:
$\big|(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q)(x,c)-(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q^{\prime})(x,c)\big|$
$\qquad=\gamma\left|\alpha\log\!\Big(\sum_{c^{\prime}}\mu_{0}^{M}(c^{\prime}\mid x^{\prime})e^{Q(x^{\prime},c^{\prime})/\alpha}\Big)-\alpha\log\!\Big(\sum_{c^{\prime}}\mu_{0}^{M}(c^{\prime}\mid x^{\prime})e^{Q^{\prime}(x^{\prime},c^{\prime})/\alpha}\Big)\right|.$
Define, for the (fixed) probability vector $b=\mu_{0}^{M}(\cdot\mid x^{\prime})$,
$f(b,q)\triangleq\alpha\log\sum_{i}b_{i}e^{q_{i}/\alpha},\qquad q\triangleq Q(x^{\prime},\cdot),\quad q^{\prime}\triangleq Q^{\prime}(x^{\prime},\cdot).$
By Lemma 16(1) (Lipschitz in $q$ for fixed $b$), we have
$\left|f(b,q)-f(b,q^{\prime})\right|\leq\|q-q^{\prime}\|_{\infty}=\|Q(x^{\prime},\cdot)-Q^{\prime}(x^{\prime},\cdot)\|_{\infty}\leq\|Q-Q^{\prime}\|_{\infty}.$
Therefore, for all $(x,c)$,
$\big|(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q)(x,c)-(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q^{\prime})(x,c)\big|\leq\gamma\|Q-Q^{\prime}\|_{\infty}.$
Taking the supremum over $(x,c)$ yields
$\|\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q-\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q^{\prime}\|_{\infty}\leq\gamma\|Q-Q^{\prime}\|_{\infty},$
so $\mathcal{T}^{\mu,M}_{\mathrm{KL}}$ is a $\gamma$-contraction. Existence and uniqueness of the fixed point follow from Banach’s fixed-point theorem. ∎
Remark 18 (Stability under memory perturbations).
The same Lemma 16 also provides stability of the evaluation operator to changes in memory-induced base measures. In particular, under boundedness of $Q$, Lemma 16(2) implies that for any $M,M^{\prime}$,
$\sup_{x^{\prime}}\left|\alpha\log\!\sum_{c^{\prime}}\mu_{0}^{M}(c^{\prime}\mid x^{\prime})e^{Q(x^{\prime},c^{\prime})/\alpha}-\alpha\log\!\sum_{c^{\prime}}\mu_{0}^{M^{\prime}}(c^{\prime}\mid x^{\prime})e^{Q(x^{\prime},c^{\prime})/\alpha}\right|\leq 2\alpha e^{\Delta_{Q}/\alpha}\,\|M^{\prime}-M\|.$
This perturbation bound is used later to establish the Lipschitz continuity of the moving fixed point $Q_{M}^{\star}$ and to derive the tracking recursion for two-timescale read–write learning.
We next establish monotone improvement results for Read–Write Learning Iterations (Theorem 19).
B.3 Monotone Improvement of Read-Write Iterations
Our analysis relies on the Gibbs variational principle, also known as the log-sum-exp conjugate duality [BoydVandenberghe2004, Section 3.3], which connects entropy-regularised optimisation to exponential-family distributions. In particular, the Gibbs principle states that
$\alpha\log\sum_{i}e^{q_{i}/\alpha}=\max_{\nu\in\Delta_{n}}\left\{\sum_{i}\nu_{i}q_{i}-\alpha H(\nu)\right\},$
where $H(\nu)=-\sum_{i}\nu_{i}\log\nu_{i}$ denotes the Shannon entropy. This identity characterises the log-sum-exp operator as the optimal value of an entropy-regularised linear objective.
Our setting extends this principle by incorporating a base measure through the Kullback–Leibler divergence $\mathrm{KL}(\nu\|\mu_{0})$, replacing entropy regularisation with relative entropy. This yields a Parzen-style exponential-family formulation and corresponds to the Bregman divergence induced by negative entropy [Ziebart2010].
Theorem 19 (Monotone Improvement Under Parzen-KL).
Let $Q^{\mu}$ be the fixed point of $\mathcal{T}^{\mu}_{\mathrm{KL}}$ and define $\mu^{+}$ via (11). Then:
$V^{\mu^{+}}(x)\geq V^{\mu}(x)\quad\text{for all }x$
proof:.
By the generalised Gibbs variational principle with base measure $\mu_{0}$:
$\alpha\log\sum_{c}\mu_{0}(c\mid x)e^{Q^{\mu}(x,c)/\alpha}$ (24)
$=\max_{\nu\in\Delta(M)}\left\{\sum_{c}\nu(c)Q^{\mu}(x,c)-\alpha\,\mathrm{KL}(\nu\|\mu_{0}(\cdot\mid x))\right\}$ (25)
The maximiser is exactly $\mu^{+}$ from (11). Therefore:
$V^{\mu^{+}}(x)$ $=\sum_{c}\mu^{+}(c\mid x)Q^{\mu}(x,c)-\alpha\,\mathrm{KL}(\mu^{+}(\cdot\mid x)\|\mu_{0}(\cdot\mid x))$ (26)
$=\alpha\log\sum_{c}\mu_{0}(c\mid x)e^{Q^{\mu}(x,c)/\alpha}$ (27)
$\geq\sum_{c}\mu(c\mid x)Q^{\mu}(x,c)-\alpha\,\mathrm{KL}(\mu(\cdot\mid x)\|\mu_{0}(\cdot\mid x))$ (28)
$=V^{\mu}(x)$ (29)
The inequality follows since $\mu^{+}$ maximises the objective defining $V^{\mu}$.∎
Finally, we derive the proof of Theorem 7 for the convergence of the Read-Write Learning.
B.4 Convergence of Read-Write Iterations
Proof.
Let $(Q_{k},\mu_{k})$ be the sequence of iterates. By Theorem 19, the soft values satisfy $V^{\mu_{k+1}}(x)\geq V^{\mu_{k}}(x)$ for all $x$.
Since rewards are bounded by $R_{\max}$ and $\gamma<1$, soft values are uniformly bounded:
$V^{\mu}(x)\leq\frac{R_{\max}}{1-\gamma}+\alpha\log|M|.$
The last term uses $\mathrm{KL}(\nu\|\mu_{0})\geq-\log|M|$ for any $\nu\in\Delta(M)$. Since $\{V^{\mu_{k}}\}$ is bounded and monotone increasing, it converges pointwise to some limit $\bar{V}$. By Theorem 17, each evaluation stage converges to the unique $Q^{\mu_{k}}$. The sequence $\{Q^{\mu_{k}}\}$ lies in a compact subset of bounded functions. Any limit point $(\bar{Q},\bar{\mu})$ must satisfy:
  • Bellman consistency: $\bar{Q}=\mathcal{T}^{\bar{\mu}}_{\mathrm{KL}}\bar{Q}$
  • KL optimality: $\bar{\mu}$ maximizes the KL-regularized objective given $\bar{Q}$
By continuity of the improvement mapping (11), these conditions uniquely determine the fixed point, so the entire sequence converges. ∎
Appendix C Lipschitz Properties of Memory Updates
In practice, memory updates typically follow one of these patterns:
  1. 1.
    Experience Replay Buffer:
    $M_{t+1}=\begin{cases}M_{t}\cup\{(s_{t},a_{t},r_{t})\}&\text{with probability }\rho_{t}\\ M_{t}&\text{otherwise}\end{cases}$ (30)
  2. 2.
    Sliding Window Memory:
    $M_{t+1}=(M_{t}\setminus\{\text{oldest case}\})\cup\{(s_{t},a_{t},r_{t})\}\quad\text{with probability }\rho_{t}$ (31)
  3. 3.
    Importance-Weighted Memory:
    $M_{t+1}=M_{t}\cup\{(s_{t},a_{t},r_{t},w_{t})\}\quad\text{where }w_{t}=f(Q_{t},\mu_{t})$ (32)
Throughout, we make explicit the role of memory as an operator-valued state rather than a raw set. In other words, although memory $M_{t}$ is implemented as a buffer of cases, the algorithm accesses memory only through the induced base measure
$\mu_{0}^{M}(\cdot\mid x),$
used by the Parzen–KL prior and the KL-regularised evaluation operator.
Definition 20 (Memory distance induced by base measures).
Let $\mathcal{C}$ be a (countable) universe of case identifiers. Each memory buffer $M\subseteq\mathcal{C}$ induces a memory-dependent base measure
$\mu_{0}^{M}(\cdot\mid x)\in\Delta_{\mathcal{C}},\qquad\text{with }\;\mu_{0}^{M}(c\mid x)=0\;\;\text{for all }c\notin M.$
We define the memory distance between two memory buffers $M$ and $M^{\prime}$ at the operator level as
$\|M^{\prime}-M\|\;\triangleq\;\sup_{x}\mathrm{TV}\!\left(\mu_{0}^{M^{\prime}}(\cdot\mid x),\mu_{0}^{M}(\cdot\mid x)\right)\;=\;\frac{1}{2}\sup_{x}\big\|\mu_{0}^{M^{\prime}}(\cdot\mid x)-\mu_{0}^{M}(\cdot\mid x)\big\|_{1}.$ (33)
All subsequent norms involving memory are understood in this sense.
Assumption 21 (Bounded $Q$-range).
For all $x\in\mathcal{X}$ and $c\in\mathcal{C}$, $Q(x,c)\in[Q_{\min},Q_{\max}]$ and $\Delta_{Q}\triangleq Q_{\max}-Q_{\min}$.
Lemma 22 (Lipschitz continuity of the Parzen–KL targets in memory).
Assume 2021. Fix any $Q$ and define, for each $x$,
$G_{M}(x;Q)\;\triangleq\;\alpha\log\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x)\exp\!\big(Q(x,c^{\prime})/\alpha\big),$
and
$\Phi_{M}(c\mid x;Q)\;\triangleq\;\frac{\mu_{0}^{M}(c\mid x)\exp(Q(x,c)/\alpha)}{\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x)\exp(Q(x,c^{\prime})/\alpha)}.$
Then for any $M,M^{\prime}$ and all $x$,
$\big|G_{M}(x;Q)-G_{M^{\prime}}(x;Q)\big|$ $\leq\alpha e^{\Delta_{Q}/\alpha}\,\|\mu_{0}^{M}(\cdot\mid x)-\mu_{0}^{M^{\prime}}(\cdot\mid x)\|_{1},$ (34)
$\|\Phi_{M}(\cdot\mid x;Q)-\Phi_{M^{\prime}}(\cdot\mid x;Q)\|_{1}$ $\leq 2e^{\Delta_{Q}/\alpha}\,\|\mu_{0}^{M}(\cdot\mid x)-\mu_{0}^{M^{\prime}}(\cdot\mid x)\|_{1}.$ (35)
Consequently,
$\sup_{x}|G_{M}(x;Q)-G_{M^{\prime}}(x;Q)|\leq 2\alpha e^{\Delta_{Q}/\alpha}\,\|M^{\prime}-M\|,$
$\sup_{x}\|\Phi_{M}(\cdot\mid x;Q)-\Phi_{M^{\prime}}(\cdot\mid x;Q)\|_{1}\leq 4e^{\Delta_{Q}/\alpha}\,\|M^{\prime}-M\|.$
Proof.
Fix $x$ and abbreviate $b=\mu_{0}^{M}(\cdot\mid x)$, $b^{\prime}=\mu_{0}^{M^{\prime}}(\cdot\mid x)$, and $w(c)=\exp(Q(x,c)/\alpha)$. By Assumption 21, $w(c)\in[e^{Q_{\min}/\alpha},e^{Q_{\max}/\alpha}]$.
(i) Lipschitzness of $G_{M}$.
Write $G_{M}(x;Q)=\alpha\log(b^{\top}w)$. This map is differentiable in $b$ with $\nabla_{b}(\alpha\log(b^{\top}w))=\alpha\,w/(b^{\top}w)$. Since $b^{\top}w\geq\min_{c}w(c)=e^{Q_{\min}/\alpha}$,
$\|\nabla_{b}(\alpha\log(b^{\top}w))\|_{\infty}\leq\alpha\,\frac{\max_{c}w(c)}{\min_{c}w(c)}=\alpha e^{(Q_{\max}-Q_{\min})/\alpha}=\alpha e^{\Delta_{Q}/\alpha}.$
By the mean value theorem and Hölder’s inequality,
$|G_{M}(x;Q)-G_{M^{\prime}}(x;Q)|=|\nabla_{b}G_{\xi}^{\top}(b-b^{\prime})|\leq\|\nabla_{b}G_{\xi}\|_{\infty}\|b-b^{\prime}\|_{1}\leq\alpha e^{\Delta_{Q}/\alpha}\|b-b^{\prime}\|_{1},$
which proves (34).
(ii) Lipschitzness of $\Phi_{M}$.
Let $w_{i}=e^{q_{i}/\alpha}$ and $Z(b)=b^{\top}w$. Then $\psi(b)_{i}=b_{i}w_{i}/Z(b)$. Write
$\psi(b)-\psi(b^{\prime})=\frac{b\odot w}{Z(b)}-\frac{b^{\prime}\odot w}{Z(b^{\prime})}=\underbrace{\frac{(b-b^{\prime})\odot w}{Z(b)}}_{(A)}+\underbrace{(b^{\prime}\odot w)\Big(\frac{1}{Z(b)}-\frac{1}{Z(b^{\prime})}\Big)}_{(B)},$
where $\odot$ is elementwise product. For (A), using $Z(b)\geq\min_{i}w_{i}=e^{Q_{\min}/\alpha}$,
$\|(A)\|_{1}\leq\frac{\|(b-b^{\prime})\odot w\|_{1}}{Z(b)}\leq\frac{\max_{i}w_{i}}{\min_{i}w_{i}}\,\|b-b^{\prime}\|_{1}=e^{\Delta_{Q}/\alpha}\|b-b^{\prime}\|_{1}.$
For (B), note $|Z(b)-Z(b^{\prime})|\leq\|b-b^{\prime}\|_{1}\max_{i}w_{i}$ and $Z(b)Z(b^{\prime})\geq(\min_{i}w_{i})^{2}$. Also $\|b^{\prime}\odot w\|_{1}=Z(b^{\prime})\leq\max_{i}w_{i}$. Thus
$\|(B)\|_{1}\leq\|b^{\prime}\odot w\|_{1}\frac{|Z(b)-Z(b^{\prime})|}{Z(b)Z(b^{\prime})}\leq(\max w)\frac{\|b-b^{\prime}\|_{1}(\max w)}{(\min w)^{2}}=e^{\Delta_{Q}/\alpha}\|b-b^{\prime}\|_{1}.$
Summing the two bounds yields (35). Finally, applying Definition 20 gives $\|b-b^{\prime}\|_{1}\leq 2\|M^{\prime}-M\|$ uniformly over $x$, proving the stated consequences. ∎
Lemma 23 (Lipschitz continuity of $F_{M}$ in memory for updates (14)–(15)).
Consider the sample-form updates:
$\delta_{t}(Q,\mu_{0},M)$ $:=r_{t}+\gamma\,\alpha\log\!\Big(\sum_{c^{\prime}\in M}\mu_{0}^{M}(c^{\prime}\mid x_{t+1})e^{Q(x_{t+1},c^{\prime})/\alpha}\Big)-Q(x_{t},c_{t}),$
$\Phi_{t}(c\mid x;Q,\mu_{0},M)$ $:=\frac{\mu_{0}^{M}(c\mid x)e^{Q(x,c)/\alpha}}{\sum_{c^{\prime}\in M}\mu_{0}^{M}(c^{\prime}\mid x)e^{Q(x,c^{\prime})/\alpha}}.$
Let $F_{M}(Z)$ denote the mean-field drift of $(Q,\mu)$ induced by these targets (i.e., conditional expectations given the current state). Under 2021, there exists $L_{F}<\infty$ such that for any fixed $Z=(Q,\mu)$,
$\|F_{M^{\prime}}(Z)-F_{M}(Z)\|\leq L_{F}\,\|M^{\prime}-M\|.$
In particular, one may take $L_{F}$ proportional to $e^{\Delta_{Q}/\alpha}$.
Proof.
Memory enters the TD target only through the log-sum-exp term $G_{M}(x_{t+1};Q)$, and enters the policy target only through $\Phi_{M}(\cdot\mid x_{t};Q)$. By Lemma 22, both of these are Lipschitz in $M$ under the distance (20), uniformly in $x$. Taking conditional expectations (to pass from sample form to mean-field drift) preserves Lipschitz constants. Therefore the combined drift mapping $F_{M}(Z)$ is Lipschitz in memory, with constant $L_{F}$ depending on $\gamma,\alpha$ and the bounded range $\Delta_{Q}$, and scaling as $O(e^{\Delta_{Q}/\alpha})$. ∎
Appendix D Proof of Theorem 9
We are not ready to prove Theorem 9 using the ODE method for two-timescale stochastic approximation [borkar2009stochastic, BorkarMeyn2000].
D.1 Fast Timescale: Fixed-Memory Dynamics
Fix a memory state $M$. Let $Z=(Q,\mu)$ denote the fast variables, and define the KL-regularised operator
$Z\;\mapsto\;F_{M}(Z),$
which combines KL-regularised policy evaluation and Parzen–KL policy improvement. The stochastic updates (14) and (15) are stochastic approximations of the limiting ODE
$\frac{dZ}{d\tau}=F_{M}(Z)-Z,$ (36)
where $\tau$ denotes fast time.
Assumption 24 (Uniform Contraction).
There exists $\kappa\in(0,1)$ such that for all $M$ and all $Z,Z^{\prime}$,
$\|F_{M}(Z)-F_{M}(Z^{\prime})\|\leq\kappa\|Z-Z^{\prime}\|.$
By Theorems 17 and 19, Assumption 24 holds and the ODE (36) admits a unique globally asymptotically stable equilibrium
$Z_{M}^{\star}=(Q_{M}^{\star},\mu_{M}^{\star}).$
D.2 Perturbed Contraction and Fixed-Point Lipschitz Continuity
We next quantify how the equilibrium depends on memory.
Lemma 25 (Lipschitz Continuity of the Fixed Point).
Assume:
  1. 1.
    $F_{M}$ is a $\kappa$-contraction uniformly in $M$;
  2. 2.
    $F_{M}$ is Lipschitz in memory:
    $\|F_{M^{\prime}}(Z)-F_{M}(Z)\|\leq L_{F}\|M^{\prime}-M\|\quad\forall Z.$
Then the fixed point satisfies
$\|Z_{M^{\prime}}^{\star}-Z_{M}^{\star}\|\;\leq\;\frac{L_{F}}{1-\kappa}\,\|M^{\prime}-M\|.$
Proof.
Using $Z_{M}^{\star}=F_{M}(Z_{M}^{\star})$,
$\|Z_{M^{\prime}}^{\star}-Z_{M}^{\star}\|$ $=\|F_{M^{\prime}}(Z_{M^{\prime}}^{\star})-F_{M}(Z_{M}^{\star})\|$
$\leq\|F_{M^{\prime}}(Z_{M^{\prime}}^{\star})-F_{M^{\prime}}(Z_{M}^{\star})\|+\|F_{M^{\prime}}(Z_{M}^{\star})-F_{M}(Z_{M}^{\star})\|$
$\leq\kappa\|Z_{M^{\prime}}^{\star}-Z_{M}^{\star}\|+L_{F}\|M^{\prime}-M\|.$
Rearranging yields the claim. ∎
D.3 Slow Timescale: Memory Dynamics
Memory updates take the stochastic approximation form
$M_{t+1}=M_{t}+\rho_{t}\bigl[h(M_{t},Z_{t})+\xi^{(M)}_{t}\bigr$], (37)
where $Z_{t}=(Q_{t},\mu_{t})$, $h$ is the expected update, and $\xi^{(M)}_{t}$ is a martingale difference noise term. The associated slow ODE is
$\frac{dM}{d\varsigma}=h(M),$ (38)
with $\varsigma$ denoting slow time. By Assumption 8(iv), this ODE admits a compact attractor set $\mathcal{M}_{\infty}$.
We also have $\|M_{t+1}-M_{t}\|=O(\rho_{t})$, a.k.a., $\sup_{t}\|M_{t+1}-M_{t}\|/\rho_{t}<\infty$. Under the distance (20), each of the following practical update schemes induces a change of order $\rho_{t}$:
  1. 1.
    Experience replay: a new case is added with probability $\rho_{t}$, changing the induced base measure by $O(\rho_{t})$.
  2. 2.
    Sliding window: one case is replaced with probability $\rho_{t}$, yielding an $O(\rho_{t})$ change.
  3. 3.
    Importance-weighted memory: weights are updated by a step $\rho_{t}$ and the induced base measure changes Lipschitzly.
Hence $\|M_{t+1}-M_{t}\|=O(\rho_{t})$ almost surely.
D.4 Tracking Lemma for Two-Timescale SA
We now state the tracking result used in the proof.
Lemma 26 (Tracking Lemma).
Suppose:
  1. 1.
    For each fixed $M$, the fast ODE (36) has a globally asymptotically stable equilibrium $Z_{M}^{\star}$;
  2. 2.
    $Z_{M}^{\star}$ is Lipschitz in $M$ as in Lemma 25;
  3. 3.
    The step sizes satisfy $\rho_{t}/\eta_{t}\to 0$.
Then
$\|Z_{t}-Z_{M_{t}}^{\star}\|\;\to\;0\quad\text{almost surely}.$
Lemma 26 follows from standard two-timescale stochastic approximation results; see [borkar2009stochastic, BorkarMeyn2000].
D.5 Derivation of the Tracking Recursion
Define the tracking error
$\epsilon_{t}=\|Z_{t}-Z_{M_{t}}^{\star}\|.$
The fast update admits the form
$Z_{t+1}=Z_{t}+\eta_{t}\bigl(F_{M_{t+1}}(Z_{t})-Z_{t}\bigr)+\eta_{t}\xi_{t},$
with martingale noise $\xi_{t}$. Using contraction of $F_{M}$, Lipschitz continuity of $Z_{M}^{\star}$, and the triangle inequality, we obtain
$\epsilon_{t+1}\;\leq\;(1-\eta_{t}\lambda)\epsilon_{t}+\eta_{t}L\|M_{t+1}-M_{t}\|+\eta_{t}\|\xi_{t}\|,$ (39)
where $\lambda=1-\kappa>0$ and $L$ is the constant from Lemma 25. Since $\|M_{t+1}-M_{t}\|=O(\rho_{t})$ and $\rho_{t}=o(\eta_{t})$, the drift term is negligible relative to the contraction.
D.6 Convergence to the Limit Set
By Lemma 26, $\epsilon_{t}\to 0$ almost surely. Since $M_{t}\to\mathcal{M}_{\infty}$, all limit points of $(Z_{t},M_{t})$ lie in
$\mathcal{L}=\{(Z_{M}^{\star},M):M\in\mathcal{M}_{\infty}\}.$
If $M_{t}\to M_{\infty}$, then $(Q_{t},\mu_{t})\to(Q_{M_{\infty}}^{\star},\mu_{M_{\infty}}^{\star})$ almost surely.
This completes the proof.
Appendix E Comparison between SRDP and Reflected MDP
The Stateful Reflection Decision Process (SRDP; Definition 2) specifies the generative mechanism by which an LLM-based agent interacts with its environment through episodic memory, reflection, and action execution. It explicitly models the internal read–write structure of reflection and the stochastic action selection mediated by the LLM.
The Reflected MDP (Definition 3) is an induced control process obtained by collapsing the internal reflection–action pipeline of SRDP into a single Markovian decision layer over augmented states. It captures the effective dynamics perceived by the agent after the reflection mechanism has been fixed, and serves as the appropriate abstraction for theoretical analysis and reinforcement learning.
Table 1 summarises the key distinctions between the two formulations.
Table 1: Comparison between SRDP and Reflected MDP.
Aspect SRDP Reflected MDP
Role in the framework Generative decision process describing how reflection, memory, and actions are produced. Induced Markov decision process capturing the effective control dynamics.
Definition level Explicitly models internal reflection (read/write) and LLM-mediated action sampling. Abstracts away internal mechanisms into a single-stage decision process.
State space $\mathcal{S}$: environment states, with memory handled as an external component. $\mathcal{X}=\mathcal{S}\times\mathcal{M}$: augmented states including memory.
Action space $\mathcal{A}$: primitive environment actions produced by the LLM. $\mathcal{C}$: reflection (retrieval) actions that index memory cases.
Decision mechanism Two-stage: retrieval policy $\mu(s,M)$ followed by LLM kernel $p_{\mathrm{LLM}}(a\mid s,c)$. Single-stage: choosing $c$ directly determines expected transitions and rewards.
Transition kernel Environment dynamics $\mathcal{P}(s^{\prime}\mid s,a)$ applied after LLM action generation. $\widetilde{\mathcal{P}}(x^{\prime}|x,c)=\mathbb{E}_{a\sim p_{\mathrm{LLM}}(\cdot|s,c)}\big[\mathcal{P}(s^{\prime}\mid s,a)\,\delta_{M^{\prime}}(\mathsf{Write}())\big].$
Reward function $\mathcal{R}(s,a)$. $\widetilde{\mathcal{R}}(x,c)=\mathbb{E}_{a\sim p_{\mathrm{LLM}}(\cdot\mid s,c)}[\mathcal{R}(s,a)]$.
Memory dynamics Memory is explicitly read and written as part of the interaction loop. Memory evolution is embedded into state transitions over $\mathcal{X}$.
Markov property Not Markovian in $\mathcal{S}$ due to latent reflection and memory updates. Markovian in the augmented state space $\mathcal{X}$.
Interpretation Reflective process: how the agent reasons and acts. Reflected process: how the environment appears after reflection is folded in.
Primary use Conceptual and mechanistic description of LLM-based reflection. Theoretical analysis, planning, and reinforcement learning.