### v3

# Memento 2: Learning by Stateful Reflective Memory

Abstract
We present a theoretical study of continual and experiential learning in Large Language Model (LLM) agents that integrate episodic memory with reinforcement learning. We identify reflection, the ability of an LLM to leverage past experience to inform future actions, as the key mechanism enabling continual adaptation without fine-tuning model parameters. Empirical results on Memento [64] and Case-Based Reasoning LLMs [24, 23] show that episodic, experience-driven reflection supports generalised adaptation in a wide range of open-ended, long-horizon tasks, implying that efficient learning can occur during deployment and blurring the traditional separation between training and testing in machine learning. Motivated by this observation, we introduce the Stateful Reflective Decision Process (SRDP), a formal abstraction of reflective memory dynamics, where an agent maintains episodic memory and performs two key operations: writing, which stores interaction outcomes and corresponds to policy evaluation, and reading, which retrieves relevant cases to make informed reflective decisions and corresponds to policy improvement. This frames reflective memory as a control-theoretic object amenable to classical reinforcement learning analysis. We instantiate this Read–Write Reflective Learning by integrating retrieval into soft policy iteration, with guaranteed convergence. We formally show that as memory expands and covers the state space more densely, the composite policy converges to the optimal solution. This theoretical framework unifies heuristic approaches such as case-based reasoning and retrieval-augmented generation with principled reinforcement learning and provides a rigorous mathematical basis for building reflective, memory-embedded LLM agents capable of continual general-purpose learning.
Figure 1: Learning by Stateful Reflection with Memory realises continual learning without fine-tuning LLMs by iteratively reading from (policy improvement) and writing to (policy evaluation) an evolving episodic memory that drives the agent’s online adaptation. Empirically, this Read-Write Learning paradigm has shown strong effectiveness across diverse domains, including software testing [24], automated data science [23], and deep research agents [64], demonstrating that memory-driven reflection can endow LLM agents with genuine self-improving capabilities.
1 Introduction
Modern machine learning is about learning from experience [51, 47]; both supervised learning and reinforcement learning can be viewed as processes of optimising a parameterised model through experience, either from human-labelled examples or data-driven feedback. The objective is to learn a mapping from inputs to outputs, or in the case of reinforcement learning, from states to actions, such that the resulting function generalises effectively to new, unseen situations [46, 52]. Formally, the learner assumes that the target function is constrained by a family of parameterised models, for instance, neural networks, where the parameters determine the precise form of the mapping.
In supervised learning, these parameters are typically adjusted by minimising a loss function defined over labelled examples, whereas in reinforcement learning (RL), the supervision signal is indirect: instead of labelled outputs, the agent interacts with an environment, receives rewards, and adjusts its parameters to maximise the expected cumulative return [49]. In both cases, the parameters are updated by backpropagation, a gradient-based optimisation algorithm that propagates errors through the network to refine the model [43].
Despite their substantial empirical successes, current machine learning paradigms suffer from fundamental inefficiencies in sample complexity. Learning via backpropagation and gradient-based optimisation typically requires an enormous number of training examples, whether labelled data in supervised learning or interaction trajectories in reinforcement learning, before competent behaviour emerges [20, 22]. From a theoretical perspective, this reflects the difficulty of identifying generalisable structure in high-dimensional, non-linear hypothesis classes when learning is driven almost entirely by statistical correlations [46]. In the absence of explicit access to semantic abstractions, compositional structure, or task-level meaning, learning algorithms must recover regularities indirectly through averaging over large numbers of samples, leading to intrinsically high sample complexity.
The training of large language models continuously exemplifies this frustration. Although massive pre-training endows them with broad linguistic and reasoning abilities, adapting them to specific tasks or domains still has to resort to additional fine-tuning, such as supervised fine-tuning (SFT) or reinforcement learning with human feedback (RLHF) [37], each demanding further large quantities of curated data or human annotations [32].
In contrast, humans leverage semantic representations, episodic memory, and reflective reuse of experience to adapt with far fewer examples [19]. This contrast suggests that learning mechanisms beyond purely statistical optimisation, based on stateful memory and semantic reflection, may offer a principled path to lower complexity learning without large-scale gradient updates.
Interestingly, despite remaining data hungry and inefficient to train, current LLMs exhibit early signs of human like cognitive abilities that support iterative generalisation to unseen tasks from only a small number of examples. These emerging abilities can be broadly categorised into three forms of what we term generalised reflection, as illustrated in Figure 2:
  1. 1.
    In-context learning. As early as GPT-3, researchers observed that LLMs could learn from a few human-provided examples directly within a prompt, a process known as in-context learning, where a few labelled examples guide behaviour on new tasks [14] (Figure 2 (a)). For instance, given a few examples of English-to-French translations in the prompt, GPT-3 can correctly translate a new English sentence into French without any parameter updates.
  2. 2.
    Feedback-driven reflection. Subsequent studies showed that given a specific unknown task, LLMs could refine their outputs through interaction with the environment, adjusting responses based on external feedback [60] (Figure 2 (b)). For instance, in code generation, the LLM proposes a function, runs tests to observe errors or runtime feedback, and iteratively refines the code to reach a correct solution.
  3. 3.
    Internal reasoning. More recently, LLM models have displayed the ability to perform internal reasoning via chain-of-thought processes before generating final outputs [57] (Figure 2 (c)). For instance, when solving a math word problem, the model explicitly writes intermediate reasoning steps (e.g., decomposing a problem into smaller parts) before computing the final numerical answer.
Although these mechanisms vary in complexity from human-guided examples (Figure 2 (a)) to environment-driven feedback (Figure 2 (b)) to purely internal reasoning (Figure 2 (c)), they share a common principle: each relies on intermediate reflection such as information gathering or internal computation to enable adaptation and generalisation beyond the training data. Such generalised reflection has been observed not only in pre-trained large language models (LLMs) [14, 22], but also in multimodal LLMs [17] and non-language transformers [28, 61].
(a) In-context learning (b) Feedback-driven reflection (c) Chain-of-thought
Figure 2: Three modes of generalised reflection in LLM agents, where reflection broadly refers to information gathering or internal computation conducted before an answer is produced: (a) in-context learning where few-shot exemplars in the prompt steer behaviour on new tasks; (b) post-output reflection via environmental feedback; (c) internal chain-of-thought and planning before emitting an answer.
In our earlier empirical work [23, 24, 64], we investigated whether reflection through memory could enable learning without parameter updates. We first introduced this paradigm with case-based reasoning LLM agents [23], where adaptation arises from updating memory rather than fine-tuning model weights. Building on this foundation, Memento [64] formalised the idea within a Memory-Augmented Markov Decision Process (M-MDP) framework. Across domains such as software testing [24], automated data science [23, 21], and deep research [64], this memory-driven reflection paradigm has consistently shown that LLMs can achieve continual self-improvement through accumulated experience rather than model retraining.
In this paper, we further provide a theoretical analysis and formalisation and leverage the capability of generalised reflection to develop a new paradigm for sample-efficient learning from experience without relying on data-intensive model training. Instead of updating model parameters, the proposed learning process operates through intermediate reflection, driven by relevant past experiences retrieved from an episodic memory [19], as shown in Fig 1. This framework can unify diverse sources of experience, whether originating from internal reasoning (e.g., chain-of-thought), interactions with the environment (e.g., ReAct [60]), or human-provided labelled data, under a single, memory-based learning mechanism.
We establish a theory of redefining how learning and decision-making are modelled through a proposed Stateful Reflective Decision Process (SRDP). Unlike the traditional Markov Decision Process (MDP), our framework introduces several conceptual shifts. First, the agent’s state now combines the current situation with its episodic memory, allowing it to incorporate past experiences directly while still maintaining a coherent Markov structure in this enlarged space (namely reflected MDP). Second, each action unfolds in two stages: the agent first recalls a relevant past case from memory, and then the language model produces a concrete response or behaviour based on that case. Third, the environment’s transitions and rewards are defined through the probabilistic reasoning of the language model itself, meaning that the model’s internal uncertainty is embedded in how the world evolves. Finally, memory evolves over time as the agent writes new experiences into it. This continual update makes the system dynamic and self-adaptive, without violating the overall Markovian formulation.
Within this formulation, learning naturally emerges from the process of retrieval: policy improvement is achieved through more effective retrieval, while policy evaluation is realised through writing experiences back to memory. This mechanism closely mirrors the principles of episodic control and reinforcement learning in humans [19, 18]. Consequently, such learning can be efficiently implemented using retrieval-augmented generation architectures equipped with large-scale memory systems [10, 31].
A key desired property is its self-learnability. With a frozen LLM [^1] While LLM parameters may evolve over time through offline updates given adequate data, such updates typically occur at a much slower timescale. Our approach does not rule out parameter learning, but instead addresses the intermediate regime in which adaptation must occur without immediate model updates. as its reasoning and reflection core, the agent improves through trial and error when data is limited and later reuses and combines past knowledge as experience grows. It balances exploitation (using known cases), exploration (targeted search in uncertain areas), and discovery (seeking new ones) through adaptive memory operations, achieving continual learning without parameter updates.
The paper is organised as follows. We first introduce the stateful reflective decision process and its Read-Write Reflective Learning. We then present theoretical results on convergence and analyse the model’s optimality and its properties as a continuously-learnable agent. Related work is reviewed next, and the paper concludes with directions for future research.
2 Stateful Reflection as Generalised Learning Methods
Recent advances in reflective prompting, such as Tree of Thoughts [59] and Graph of Thoughts [8], have demonstrated the potential of structured reflection to enhance LLM reasoning. From a different perspective, we explore a new learning paradigm that enables models to learn from experience through reflection grounded in episodic memory, rather than through parameter fine-tuning [23, 24, 64]. To formalise this paradigm, we introduce the Stateful Reflective Decision Process (SRDP), which models continual learning as a sequence of reflective interactions with memory.
We next present the core concepts, with the full mathematical formulation deferred to later sections. Covering the three modes of generalised reflection introduced earlier (Fig. 2), we define reflection in this work as follows.
Definition 1 (Reflection).
Reflection (generalised) is an iterative mechanism by which a large language model (LLM) agent progressively improves its effective policy through coordinated interaction with episodic memory, internal reasoning processes, and environmental feedback. It is not a single operation but a two-step procedure consisting of a read step followed by a write step, which together form a closed learning loop.
Figure 1 illustrates the overall workflow of an SRDP. We begin by formally defining episodic memory as a structured repository of past interaction episodes, consistent with the classical notion of episodic memory in cognitive psychology [50].
Definition 2 (Episodic Memory).
An episodic memory $M_{t}$ is a finite (and growing) collection of memory items
$M_{t}=\{m_{i}\}_{i=1}^{N_{t}},\qquad m_{i}:=(s_{i},a_{i},r_{i},s^{\prime}_{i}),$
where each item $m_{i}$ records a past interaction consisting of the environment state $s_{i}\in S$, the executed action $a_{i}\in A$, the received reward $r_{i}\in\mathbb{R}$, and the subsequent state $s^{\prime}_{i}\in S$. The space of all finite episodic memories is denoted by $\mathfrak{M}$. Episodic memory evolves over time through a write operation
${M}_{t+1}=\mathrm{Write}({M}_{t},s_{t},a_{t},r_{t},s_{t+1}).$
Given a task or query represented as a state $s$ (1), the agent enters an internal reflection stage (2), during which the LLM performs a read operation on M{M} to retrieve relevant episodic memory items or their summaries. Conditioned on this retrieved information, the LLM selects an external action or generates an answer $a$ (3), which is executed in the environment (4) and produces evaluative feedback, including rewards, correctness signals, or task level assessments. Finally, the resulting feedback is incorporated into episodic memory via a write operation (5), which stores new interaction episodes.
Formally, let $\pi_{t}$ denote the agent’s effective policy at iteration $k$, implicitly represented by its prompt, episodic memory contents, and internal reasoning configuration. Reflection defines an operator $\mathcal{T}$ such that
$\pi_{t+1}=\mathcal{T}(\pi_{t})=\text{\emph{Read}}\bigl(\text{\emph{Write}}(\pi_{t})\bigr).$
The Read step corresponds to policy improvement: the agent conditions on retrieved episodic memories, contextual exemplars, or intermediate reasoning traces to construct an improved action distribution for the current state,
$\pi_{t}\leftarrow\text{\emph{Read}}(\pi_{t-1}\mid s_{t},M_{t}),$
where $s_{t}$ denotes the current environment state and $M_{t}$ the episodic memory state. In LLM agents, such conditioning reshapes the effective policy directly, without parameter updates.
The Write step corresponds to policy evaluation: the agent generates an action, plan, or answer under $\pi_{t+1}$, producing trajectories
$\tau_{t}:=m_{i}=(s_{i},a_{i},r_{i},s^{\prime}_{i})\sim\pi_{t},$
whose outcomes yield evaluative feedback,
$\hat{V}(\pi_{t+1})\leftarrow\text{Eval}(\tau_{t}),$
which is subsequently stored in episodic memory for future reads,
$M_{t+1}\leftarrow\text{\emph{Write}}(M_{t},\tau_{t}\sim\pi_{t}).$
As we shall show in the following sections, this alternating reflective read–write reflective mechanism is formally analogous to policy iteration in reinforcement learning. Policy improvement is realised through memory-conditioned inference, while policy evaluation is realised through interaction-driven feedback and experience accumulation. Under mild assumptions that reading yields non-decreasing policy quality with respect to accumulated feedback and that writing provides informative evaluation signals, the iterative reflection process converges to a fixed point corresponding to an optimal or self-consistent policy.
SRDP generalises and unifies several existing learning paradigms. Compared with case based reasoning (CBR), which retrieves and adapts discrete past cases for problem solving [23, 1], SRDP embeds retrieval within a sequential decision making framework. This explicitly models temporal dependencies, memory evolution, and long horizon effects under a Markovian assumption, rather than treating cases as isolated problem instances.
In contrast to retrieval augmented generation (RAG), which conditions generation on static external documents or databases [10], SRDP treats memory as a dynamic and stateful component. Episodic memory is continually updated through interaction with the environment and grounded in evaluative feedback, allowing retrieval to reflect the agent’s own experience rather than fixed external knowledge. From the perspective of memory based reinforcement learning, SRDP replaces parameter level updates with reflective memory updates mediated by an LLM. Policy improvement is achieved through structured recall and reflection over episodic memory rather than gradient based optimisation of model parameters [42]. This shift enables continual adaptation while keeping the underlying LLM fixed, positioning reflection as the primary learning mechanism.
This formulation is also strongly inspired by neuroscientific findings on the role of the hippocampus in episodic memory and reflective cognition [39, 33]. In humans and animals, episodic recall supports flexible reasoning and planning by replaying past experiences, an ability mirrored by the reflective cycles of SRDP. Thus, the Stateful Reflective Decision Process provides both a computational and cognitive bridge: it formalises how memory and reflection together enable continual, self-improving intelligence without explicit retraining.
Experimentally, it has been demonstrated that this approach enables learning without fine-tuning across various agentic tasks, including data-science agents [24], software-testing agents [23], and deep-research agents [64].
3 A Model for Reflective Process over Episodic Experience
Mathematically, the core idea is to extend the Markov Decision Process [41] with an intermediate reflection stage, where the agent retrieves relevant experiences from external memory before generating the final output, enabling continual and memory-based learning.
3.1 Preliminary: Markov Decision Process
To make self-contained, we begin by reviewing the Markov Decision Process (MDP) [41], the standard framework for modelling decision-making in agents. An LLM agent, viewed as a large language model capable of reasoning and acting through interaction with its environment [16, 58, 56, 53], can be naturally described within this framework. Formally, an MDP is defined by the tuple
$\mathcal{D}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle,$
where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, and $\mathcal{P}(s^{\prime}\mid s,a)$ denotes the transition kernel, specifying the probability of reaching state $s^{\prime}$ after taking action $a$ in state $s$; $\mathcal{R}(s,a)\in\mathbb{R}$ represents the expected immediate reward; and $\gamma\in[0,1)$ is the discount factor. Here, the state $s$ corresponds to the current prompt or task context. Specifically, in the context of LLM reasoning, states correspond to intermediate reasoning steps [56, 55], and when interacting with the external environment, they may take the form of textual descriptions of the environment state [16]. The action $a$ is the model’s generated output or decision, and the transition describes how the environment or subsequent prompt evolves after the model’s response [56]. A (stochastic) policy $\pi(a\mid s)$ defines a probability distribution over actions conditioned on the current state and, in turn, induces the value function that measures the expected return when following the policy from a given state
$V^{\pi}_{\mathcal{D}}(s)=\mathbb{E}_{\pi,\mathcal{P}}\!\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})\,\bigg|\,s_{0}=s\right],$
with the state–action value function
$Q^{\pi}_{\mathcal{D}}(s,a)=\mathcal{R}(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot\mid s,a)}\!\left[V^{\pi}_{\mathcal{D}}(s^{\prime})\right].$
The optimal value functions are
$V^{\star}_{\mathcal{D}}(s)=\max_{\pi}V^{\pi}_{\mathcal{D}}(s),\qquad Q^{\star}_{\mathcal{D}}(s,a)=\mathcal{R}(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot\mid s,a)}\!\left[\max_{a^{\prime}}Q^{\star}_{\mathcal{D}}(s^{\prime},a^{\prime})\right].$
For large language model (LLM) agents [16, 53, 62], the policy is typically given by the LLM itself, which maps input states or prompts directly to action distributions through its generative reasoning process:
$\pi_{\mathrm{LLM}}(a\mid s)\triangleq p_{\mathrm{LLM}}(a\mid s;\,\theta),$
where $\theta$ denotes the (frozen in our case) parameters of the pre-trained model. Next, we extend the LLM’s stochastic policy by incorporating contributions from episodic memory, allowing past experiences to influence current decisions.
3.2 Stateful Reflective Decision Process
Building on the standard Markov Decision Process (MDP) framework introduced in the previous section, we now present the proposed Stateful Reflective Decision Process (SRDP). The classical MDP $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$ captures agent–environment interaction in its simplest form but lacks mechanisms for reflection and memory. To enable reflective, memory-driven behaviour, we extend this formulation by augmenting the agent with an episodic memory space $M$, a memory retrieval policy $\mu$, and a stochastic LLM kernel $p_{\mathrm{LLM}}$ that generates context-dependent actions conditioned on both the current environment state and retrieved memory, yielding the Stateful Reflective Decision Process (SRDP).
3.2.1 Definition
Definition 3 (Stateful Reflective Decision Process (SRDP)).
SRDP is a tuple
$\mathcal{D}_{\mathrm{SRDP}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathfrak{M},p_{\mathrm{LLM}}\rangle$
that extends the standard MDP formalism by incorporating episodic memory and a language-model-based decision kernel. The components are defined as follows:
  • $\mathcal{S}$: state space;
  • $\mathcal{A}$: environment action space;
  • $\mathcal{P}(\cdot\mid s,a)$: transition kernel;
  • $\mathcal{R}(s,a)\in\mathbb{R}$: reward function;
  • $\gamma\in[0,1)$: discount factor;
  • $\mathfrak{M}$: space of finite episodic memories, where each memory $M\in\mathfrak{M}$ is a multiset of cases $c=(s,a,r,s^{\prime})$ [^2] In multi-state settings, $r$ can be replaced by an estimated value $Q$, where Bellman or TD updates are propagated across memory.;
  • $p_{\mathrm{LLM}}(a\mid s,c)$: the LLM kernel that generates an environment action conditioned on the current state $s$ and a retrieved case $c$.
At time $t$, the agent maintains episodic memory $M_{t}=\{m_{i}\}_{i=1}^{N_{t}}$, where $N_{t}$ is the number of cases in the current memory, and observes the current environment state $s_{t}$. The decision process unfolds in two stages:
  1. 1.
    Retrieval stage: A retrieval action $c_{t}$ is sampled from the current episodic memory according to the retrieval policy
    $c_{t}\sim\mu(\cdot\mid s_{t},M_{t}),$
    where $\mu$ may be similarity-based or learned. The action set is memory-dependent (indexer): for each $M\in\mathfrak{M}$, $\mathcal{C}(M):=M$, and each $c\in\mathcal{C}(M)$ indexes a memory case used for retrieval.
  2. 2.
    Action Stage: The LLM generates an action conditioned on the retrieved case:
    $a_{t}\sim p_{\mathrm{LLM}}(\cdot\mid s_{t},c_{t}).$
This defines a composite policy:
$\pi^{\mu}(a_{t}\mid s_{t},M_{t})=\sum_{c_{t}\in M_{t}}\mu(c_{t}\mid s_{t},M_{t})\,p_{\mathrm{LLM}}(a_{t}\mid s_{t},c_{t}).$ (1)
After executing the action $a_{t}$, the environment changes to $s_{t+1}\sim\mathcal{P}(\cdot\mid s_{t},a_{t})$, yields a reward $r_{t}:=\mathcal{R}(s_{t},a_{t})$, and the memory updates to $M_{t+1}=\mathsf{Write}(M_{t},s_{t},a_{t},r_{t},s^{\prime}_{t})$.
$s_{0}$$M_{0}$$c_{0}$$a_{0}$$r_{0}$$s_{1}$$M_{1}$$c_{1}$$a_{1}$$r_{1}$$s_{t}$$M_{t}$$c_{t}$$a_{t}$$r_{t}$$x_{0}$$x_{1}$$x_{t}$Environment DynamicsMemory Read Policy $\mu(c|x)$LLM Conditioning $p_{\mathrm{LLM}}(a|s,c)$Memory Write $\mathsf{Write}(M,s,a,r)$Reward Generation
Figure 3: A graphical model of the Stateful Reflective Decision Process (SRDP) showing the augmented state space $x_{t}=(s_{t},M_{t})$, retrieval actions $c_{t}$, LLM conditioning, and memory evolution. Dashed boxes indicate augmented states combining environment state and memory.
Figure 3 depicts the graphical model of the Stateful Reflective Decision Process (SRDP), characterised by a dual-action structure. The agent first performs a retrieval action $c_{t}\in M_{t}$ (green nodes), selecting a relevant past case from memory to guide its current decision. This retrieval is controlled by the policy $\mu(c\mid s,M)$, represented by the blue arrows connecting the current environment state $s_{t}$ and memory $M_{t}$ to the retrieval node $c_{t}$. The policy evaluates both the similarity between the current state and stored experiences, and the local density of memory cases. Conditioned on the retrieved case, the LLM then reflects upon the retrieved experience and generates the environment action $a_{t}\in\mathcal{A}$ (blue nodes) via its generative kernel $p_{\mathrm{LLM}}(a\mid s,c)$, shown by the green arrows. Through this hierarchical process, the retrieval stage provides contextual grounding, while the LLM generates the final action, jointly realising stateful reflection, as illustrated in Figure 1.
A key advantage of this dual-action structure is that it defines a composite policy (Eq. 1) that unifies classical reinforcement learning [49] with LLM-based decision-making [16], thereby combining the strengths of both paradigms. To illustrate, consider two limiting cases. In the first case, a naive LLM produces an action $a\in c$ directly from a retrieved case $c$, effectively treating the case as a prompt. Here, the policy reduces to a standard reinforcement learning or episodic control scheme [9, 40, 42], where the retrieval function $\mu$ selects the case $c$ with the highest $Q$-value, and the corresponding action is executed. In the opposite extreme, when no relevant experience exists in memory, the agent must rely entirely on the internal knowledge of the LLM to generate actions. Between these two cases lies the general setting of SRDP, where the agent can integrate both sources of information, the reasoning capability of the LLM and the experiential knowledge stored in memory, to produce more adaptive and effective decisions.
The memory evolution mechanism forms a crucial component of the SRDP dynamics. Purple arrows demonstrate how memory grows through $\mathsf{Write}(M,s,a,r)$ operations, with experience tuples $(s_{t},a_{t},r_{t})$ being incorporated into the memory at the next time step. This creates a directed evolution $M_{0}\to M_{1}\to\cdots\to M_{t}$ where the memory size $|M_{t}|$ increases over time, expanding the agent’s retrieval options.
However, the introduction of memory fundamentally alters the dynamics of the system, making the process no longer strictly Markovian if the original state is considered. In a standard MDP, the next state depends only on the current state and action, i.e., the system satisfies $P(s_{t+1}\mid s_{t},a_{t},s_{t-1},a_{t-1},\ldots)=P(s_{t+1}\mid s_{t},a_{t}).$ However, once we introduce an evolving memory $M_{t}$ that accumulates past experience tuples $(s_{i},a_{i},r_{i})$, the transition to the next augmented state depends on the entire history through $M_{t}$. Formally, $P(s_{t+1},M_{t+1}\mid s_{t},a_{t},M_{t})\neq P(s_{t+1},M_{t+1}\mid s_{t},a_{t})$ because $M_{t}$ itself is a summary of all past interactions. We address this issue next by defining a simplified MDP that incorporates memory directly into the state representation while embedding the fixed LLM within the environment dynamics. This formulation enables us to focus on optimising the retrieval policy, with both the LLM’s behaviour and memory evolution captured within the new state–transition structure.
3.2.2 Transforming to Reflected MDP
To recover a Markovian formulation, we redefine the state to explicitly include memory, forming an augmented state, i.e., at time $t$, the agent occupies an augmented state
$x_{t}:=(s_{t},M_{t})\in\mathcal{X}:=S\times\mathfrak{M}.$
The resulting process over $\{x_{t}\}$ is Markovian in this expanded space, even though the environment process over $\{s_{t}\}$ alone is not. Each augmented state combines the current environment state $s_{t}$ with the agent’s episodic memory $M_{t}$, as illustrated by the dashed rectangles in Figure 3. Augmenting the state with memory restores the Markov property and enables a well-defined control formulation, i.e., for any reflection action $c_{t}$,
$\Pr(x_{t+1}\mid x_{0:t},c_{0:t})=\Pr(x_{t+1}\mid x_{t},c_{t}),$
where the environment transition depends only on $(s_{t},a_{t})$, while the memory update is a deterministic or stochastic function of $(M_{t},s_{t},a_{t},r_{t},s^{\prime}_{t})$. Since the reflection mechanism and the LLM action distribution are fixed functions of $(s_{t},M_{t})$, all sources of dependence on the past are captured by $x_{t}$. Hence the augmented process satisfies the Markov property.
Since our focus is on learning through memory retrieval rather than modifying the LLM itself, we treat the LLM as a fixed or slow-evolving component of the environment. Its action generation is absorbed into the environment dynamics via the effective transition and reward functions $\mathcal{P}^{\mathrm{LLM}}$ and $\mathcal{R}^{\mathrm{LLM}}$. Under this view, the agent’s only controllable decision is the retrieval action
$c_{t}\sim\mu(\cdot\mid s_{t},M_{t}),$
while state transitions and rewards are determined jointly by the environment and the fixed LLM kernel $p_{\mathrm{LLM}}(a\mid s_{t},c_{t})$. The resulting Reflected MDP therefore treats the retrieval policy $\mu$ as the agent’s primary policy, with LLM mediated reasoning incorporated into the environment model. Specifically, we have:
Definition 4 (Reflected MDP).
The Reflected MDP is a tuple
$\mathcal{D}_{\mathrm{ReMDP}}=\langle\mathcal{X},\mathcal{C},\mathcal{P}_{\mathrm{LLM}},\mathcal{R}_{\mathrm{LLM}},\gamma\rangle,$
that transforms the underlying SRDP, $\mathcal{D}_{\mathrm{SRDP}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathfrak{M},p_{\mathrm{LLM}}\rangle$, with the components are defined as follows:
  • Augmented state space: $\mathcal{X}=\mathcal{S}\times\mathfrak{M}$, with each state $x=(s,M)$ consisting of the current environment state $s$ and the current episodic memory $M$.
  • Action space: $\mathcal{C}(M):=M$, consisting of retrieval actions that index and select cases from the current episodic memory.
  • Transition kernel: For $x=(s,M)$ and $c\in\mathcal{C}$,
    $\mathcal{P}_{\mathrm{LLM}}(x^{\prime}\mid x,c)=\sum_{a\in\mathcal{A}}p_{\mathrm{LLM}}(a\mid s,c)\,\mathbf{1}\{x^{\prime}=(s^{\prime},\mathsf{Write}(M,s,a,\mathcal{R}(s,a)))\}\,\mathcal{P}(s^{\prime}\mid s,a),$ (2)
    where $\mathsf{Write}(M,s,a,r)$ appends the new experience $(s,a,r)$ to memory.
  • Reward function:
    $\mathcal{R}_{\mathrm{LLM}}(x,c)=\sum_{a\in\mathcal{A}}p_{\mathrm{LLM}}(a\mid s,c)\,\mathcal{R}(s,a).$ (3)
The Reflected MDP provides a fully Markovian formulation over the augmented state space $\mathcal{X}=\mathcal{S}\times\mathfrak{M}$, in which transitions and rewards correspond to expected outcomes under the fixed LLM kernel $p_{\mathrm{LLM}}$. At each time step, the transition unfolds as follows: the agent first retrieves a memory case $c_{t}\sim\mu(\cdot\mid s_{t},M_{t})$; conditioned on this retrieval, the LLM generates an environment action $a_{t}\sim p_{\mathrm{LLM}}(\cdot\mid s_{t},c_{t})$; the environment then transitions to $s_{t+1}\sim\mathcal{P}(\cdot\mid s_{t},a_{t})$ and produces a reward $\mathcal{R}(s_{t},a_{t})$; finally, episodic memory is updated according to $M_{t+1}=\mathsf{Write}(M_{t},s_{t},a_{t},r_{t},s^{\prime}_{t})$.
By absorbing the retrieval and LLM generation mechanisms into effective transition and reward functions over the augmented state space $\mathcal{X}$, this interaction induces a simplified control process. A policy in the Reflected MDP is defined as a retrieval policy
$\mu:\mathcal{X}\to\Delta(\mathcal{C}),$
which maps augmented states x=(s,M)x=(s,{M}) to probability distributions over retrieval actions. Environment actions are generated implicitly by the fixed LLM kernel $p_{\mathrm{LLM}}(a\mid s,c)$ and absorbed into the transition dynamics. This abstraction captures the full reflective behaviour of the agent in a compact Markovian form that is amenable to classical reinforcement learning analysis and learning algorithms.
It is worth mentioning that this augmented state admits a natural cognitive interpretation. The environment state $s_{t}$ encodes the agent’s external sensory or task-level signal, while the episodic memory $M_{t}$ represents an internal cognitive state that stores and organises past experiences, outcomes, and contextual information. The pair $(s_{t},M_{t})$ therefore constitutes the agent’s effective decision state, integrating external observation with internal experience. This perspective parallels belief-state constructions in partially observable Markov decision processes, where internal memory summarises relevant history to support optimal control [29], and aligns with cognitive theories that view episodic memory as a core substrate for human decision making and learning [50, 19].
The underlying MDP captures the dynamics of the environment alone, independent of the agent’s internal mechanisms. In contrast, the SRDP explicitly models the coupled interaction among the agent, its episodic memory, and the LLM during decision making. The Reflected MDP then provides a Markovian reformulation of this process by lifting the SRDP to an augmented state space, in which the combined agent–memory–LLM dynamics become Markovian. The relationship between the underlying environment, the SRDP, and the resulting Reflected MDP is summarised as
$\underbrace{\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle}_{\text{Underlying MDP}}\xrightarrow{\;\mathrm{ext.}\;}\underbrace{\Big\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,M,\mu,p_{\mathrm{LLM}}\Big\rangle}_{\text{SRDP}}\xrightarrow{\;\mathrm{equiv.}\;}\underbrace{\Big\langle\mathcal{X}:=\mathcal{S}\times M,\mathcal{C},\mathcal{P}^{\mathrm{LLM}},\mathcal{R}^{\mathrm{LLM}},\gamma\Big\rangle}_{\text{Reflected MDP}},$ (4)
where a detailed comparison between the SRDP and the Reflected MDP is provided in Appendix E.
4 Read–Write Reflective Learning
Under the reflected MDP abstraction, continual learning reduces to an iterative read–write reflective procedure over episodic memory, which we formalise as a learning algorithm next.
4.1 Policy Iteration
We now define the retrieval mechanism that underpins the learning method. The central idea is to use local state density to guide which past cases should be retrieved from memory. To capture the similarity structure in the memory bank, we employ Parzen window estimation [38, 54].
Let $K:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ be a smooth kernel (e.g., Gaussian) with bandwidth $h>0$, and let $\psi:\mathcal{S}\to\mathbb{R}^{d}$ be a state embedding. For memory $M=\{c_{i}=(s_{i},a_{i},r_{i})\}_{i=1}^{N}$ and query state $s$, the Parzen similarity weight for case $c\in M$ is:
$w_{\text{parzen}}(s,c)=\frac{K\left(\psi(s)-\psi(s(c))/h\right)}{\sum_{c^{\prime}\in M}K\left(\psi(s)-\psi(s(c^{\prime}))/h\right)},$ (5)
where $s(c)$ denotes the state component of a case $c$. The kernel defines a probability distribution over memory cases that places higher weight on states similar to the current query state $s$. Thus, the Parzen prior $\mu_{0}(c\mid x)$ over retrieval actions at augmented state $x=(s,M)$ is:
$\mu_{0}(c\mid x)\equiv w_{\text{parzen}}(s,c).$ (6)
In order to enable an LLM agent to act even when no relevant past cases exist in memory, we augment the case set $\mathcal{C}$ with a special void case: $M\leftarrow M\cup\{c_{\varnothing}\}$. The void case $c_{\varnothing}$ represents the ability of the agent to generate an action directly from its internal world knowledge, independently of retrieved memory. We formulate this by modifying the prior policy $\mu_{0}$ that incorporates a state-dependent mixture between memory-based retrieval and the void case. Let $K(x,c)$ be a similarity kernel for ordinary cases $c\in\mathcal{C}$, and let $K_{\varnothing}>0$ be a constant kernel score assigned to the void case $c_{\varnothing}$. Then, the prior distribution $\mu_{0}(\cdot\mid x)$ becomes
$\mu_{0}(c\mid x)\equiv\frac{\begin{cases}K(x,c),&c\neq c_{\varnothing},\\[3.0pt] K_{\varnothing},&c=c_{\varnothing},\end{cases}}{\sum_{c^{\prime}\neq c_{\varnothing}}K(x,c^{\prime})+K_{\varnothing}},$ (7)
where $K(x,c):=K\!\left(\frac{\psi(s(x))-\psi(s(c))}{h}\right)$.
It is useful to express this distribution as a mixture. Define the normalised memory-based distribution
$\mu_{\mathrm{mem}}(c\mid x)=\frac{K(x,c)}{\sum_{c^{\prime}\neq c_{\varnothing}}K(x,c^{\prime})},\qquad c\neq c_{\varnothing},$
and let
$\lambda(x)=\frac{\sum_{c^{\prime}\neq c_{\varnothing}}K(x,c^{\prime})}{\sum_{c^{\prime}\neq c_{\varnothing}}K(x,c^{\prime})+K_{\varnothing}}.$
Then (7) is equivalent to the mixture representation
$\mu_{0}(c\mid x)=\lambda(x)\,\mu_{\mathrm{mem}}(c\mid x)+\bigl(1-\lambda(x)\bigr)\,\delta_{c_{\varnothing}}(c),$ (8)
where $\delta_{c_{\varnothing}}$ denotes the unit point mass at $c_{\varnothing}$. This mixture form shows that the agent automatically interpolates between reflection(retrieval)-driven behaviour when similarity to stored cases is high ($\lambda(x)\approx 1$) and knowledge-driven discovery through $c_{\varnothing}$ when similarity is low ($\lambda(x)\approx 0$). This mechanism mirrors the “discovery” component in infinite-armed bandits [7], complementing exploitation and exploration with the capacity to act beyond the current memory. This mechanism ensures that the agent can always act, even in novel states, while still exploiting relevant past experiences when available.
We aim for the agent’s final policy $\mu(\cdot\mid x)$ to remain close to a prior policy $\mu_{0}$, while still being shaped by reward feedback from the environment. To formalise this trade-off, we define a KL-regularised objective in terms of the state–case value function $Q^{\mu}(x,c)$. The KL-regularised Bellman evaluation operator is
$(\mathcal{T}^{\mu}_{\mathrm{KL}}Q)(x,c):=\mathbb{E}_{a,s^{\prime},M^{\prime}}\!\left[r+\gamma\left(\sum_{c^{\prime}}\mu(c^{\prime}\mid x^{\prime})\,Q(x^{\prime},c^{\prime})-\alpha\,\mathrm{KL}\!\bigl(\mu(\cdot\mid x^{\prime})\,\|\,\mu_{0}(\cdot\mid x^{\prime})\bigr)\right)\,\middle|\,x,c\right],$ (9)
where $\mathcal{T}^{\mu}_{\mathrm{KL}}:\mathbb{R}^{\mathcal{X}\times\mathcal{C}}\to\mathbb{R}^{\mathcal{X}\times\mathcal{C}}$. It is worth emphasising that all value functions $Q(x,c)$ and $V(x)$ defined on the augmented state space $\mathcal{X}$ of the Reflected MDP are distinct from the value functions $Q^{\mathcal{D}}$ and $V^{\mathcal{D}}$ of the original MDP (see Section 3.1). The latter are defined over the environment state space $S$, whereas the former explicitly incorporate episodic memory and LLM-mediated action generation through the Reflected MDP dynamics. We shall show that, by appropriately leveraging these LLM-induced dynamics, the Reflected MDP can approximate the optimal value functions of the original MDP in a more sample-efficient manner.
For any fixed $\mu$, this operator is a $\gamma$-contraction in the sup-norm and admits a unique fixed point $Q^{\mu}$, representing the KL-regularised value of $\mu$.
KL-Regularised Policy Improvement Given a $Q$-function over retrieval actions, we define policy improvement as the solution to a KL-regularised optimisation problem. Namely, at each augmented state $x$, the improved retrieval policy is:
$\mu^{+}(c\mid x)\in\arg\max_{\nu\in\Delta(M)}\left\{\sum_{c\in M}\nu(c|x)\,Q(x,c)-\alpha\,\mathrm{KL}(\nu(c|x)\|\mu_{0}(\cdot\mid x))\right\},$ (10)
where $\Delta(M)$ is the simplex over finite set, memory $M$, and $\alpha>0$ is the temperature parameter.
Lemma 5 (also in [35]).
The optimisation problem (10) has a unique solution:
$\mu^{+}(c\mid x)=\frac{\mu_{0}(c\mid x)\,\exp(Q(x,c)/\alpha)}{\sum_{c^{\prime}\in M}\mu_{0}(c^{\prime}\mid x)\,\exp(Q(x,c^{\prime})/\alpha)}.$ (11)
The proof can be found in Appendix A. Because $\mu_{0}(c_{\varnothing}\mid x)>0$ for all $x$, the improved policy $\mu^{\prime}$ always assigns nonzero probability to the void case, preserving the agent’s capacity for discovery.
KL-Regularised Policy Iteration Now, on the basis of policy evaluation in Eq. (9) and policy improvement in Eq. (11), we are ready to give the policy iteration algorithm, as shown in Algorithm 1. Under standard assumptions, this procedure yields a monotone improvement in the KL-regularised value and converges to the fixed point of the KL-regularised optimality operator,
$(\mathcal{T}^{*}_{\mathrm{KL}}Q)(x,c)=\mathbb{E}\!\left[r+\gamma\,\alpha\log\!\sum_{c^{\prime}}\mu_{0}(c^{\prime}\mid x^{\prime})\,\exp\!\bigl(Q(x^{\prime},c^{\prime})/\alpha\bigr)\,\middle|\,x,c\right].$
Algorithm 1 Soft Policy Iteration with Fixed Memory under Reflected MDP
1:Input: Temperature $\alpha>0$, kernel bandwidth $h>0$
2:Initialise: Fixed episodic memory $M$, reference prior $\mu_{0}(\cdot\mid x)$ (including void case $c_{\varnothing}$), initial policy $\mu_{0}$, initial $Q_{0}$
3:for $t=0,1,2,\ldots$ do
4:  (Policy Evaluation)
5:  Compute the KL-regularised value function $Q^{\mu_{t}}$ as the unique fixed point of
$Q=\mathcal{T}^{\mu_{t}}_{\mathrm{KL}}Q,$
where (Eq. (9))
$(\mathcal{T}^{\mu_{t}}_{\mathrm{KL}}Q)(x,c)=\mathbb{E}\!\left[r+\gamma\Big(\sum_{c^{\prime}\in M}\mu_{t}(c^{\prime}\mid x^{\prime})Q(x^{\prime},c^{\prime})-\alpha\mathrm{KL}(\mu_{t}(\cdot\mid x^{\prime})\|\mu_{0}(\cdot\mid x^{\prime}))\Big)\,\middle|\,x,c\right].$
6:  (Policy Improvement)
7:  Update the retrieval policy by solving
$\mu_{t+1}(\cdot\mid x)\in\arg\max_{\nu\in\Delta(M)}\left\{\sum_{c}\nu(c\mid x)Q^{\mu_{t}}(x,c)-\alpha\mathrm{KL}(\nu(\cdot\mid x)\|\mu_{0}(\cdot\mid x))\right\},$
which admits the closed-form solution (Eq. (11))
$\mu_{t+1}(c\mid x)=\frac{\mu_{0}(c\mid x)\exp(Q^{\mu_{t}}(x,c)/\alpha)}{\sum_{c^{\prime}}\mu_{0}(c^{\prime}\mid x)\exp(Q^{\mu_{t}}(x,c^{\prime})/\alpha)}.$
8:end for
For the convergence analysis, we start with assumptions followed with the theorem:
Assumption 6 (Bounded rewards and discount).
$|r|\leq R_{\max}<\infty$, $\gamma\in[0,1)$.
Assumption 7 (Stationary memory dynamics during evaluation).
Within each policy-evaluation phase, the memory kernel is stationary (equivalently, treat $M$ as part of the state).
Theorem 8 (Convergence of Parzen-KL Soft Policy Iteration).
Under assumptions 6 and 7 (bounded rewards and $\gamma<1$ and stationary memory), the process (the reflected MDP) is an ordinary discounted MDP. Iterations between (9) and (11) (Algorithm 1) converge to a fixed point $(Q^{\star},\mu^{\star})$ that is optimal for the KL-regularised objective with Parzen prior.
We refer to Appendix for the proof.
Some discussions: It is worth mentioning that the void case $c_{\varnothing}$ plays a crucial role: by providing a state-independent baseline probability in $\mu_{0}$, it ensures that policy iteration retains the ability to discover new behaviours beyond exploitation or exploration of existing memory. Unlike classical density estimation, our memory cases are not independent and identically distributed but are gathered sequentially through policy execution. Moreover, the finite memory constraint means we perform kernel smoothing over a limited and evolving set of points rather than relying on asymptotic density estimation. As a result, bandwidth selection plays a critical role in balancing local similarity against statistical robustness as the memory continues to grow. Several strategies can be used to choose the bandwidth $h$. Silverman’s rule suggests setting $h\propto n^{-1/(d+4)}$, where $n=|M|$ is the memory size and $d$ the embedding dimension. Another approach is cross-validation, which tunes $h$ by optimising performance on held-out retrieval tasks. Finally, an adaptive strategy decreases $h$ as memory grows, ensuring that neighbourhoods remain local and informative.
4.2 Two-Time-Scale Convergence with Online Memory Rewriting
The preceding analysis assumes that episodic memory remains fixed. We now relax this assumption and allow memory to evolve on a slower time scale. Let $\eta_{t}$ denote the step size used for policy evaluation and improvement, and let $\rho_{t}$ denote the step size for memory updates. We assume a two time scale regime in which
$\frac{\rho_{t}}{\eta_{t}}\to 0,\qquad\sum_{t=1}^{\infty}\eta_{t}=\infty,\qquad\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty.$
Under this separation of time scales, policy evaluation and improvement operate on a faster time scale, while episodic memory evolves more slowly. Formally, we make the following assumptions:
Assumption 9 (Two-Time-Scale Conditions).
  1. (i)
    Step sizes: $\eta_{t}>0$, $\rho_{t}>0$ with:
    $\sum_{t=0}^{\infty}\eta_{t}=\infty,\quad\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty,\quad\sum_{t=0}^{\infty}\rho_{t}=\infty,\quad\sum_{t=0}^{\infty}\rho_{t}^{2}<\infty,\quad\lim_{t\to\infty}\frac{\rho_{t}}{\eta_{t}}=0.$
  2. (ii)
    Martingale difference noise: $\xi^{(Q)}_{t}$ and $\xi^{(\mu)}_{t}$ are martingale difference sequences with respect to the filtration $\mathcal{F}_{t}$, information available up to time $t$ (random variable with zero conditional mean) :
    $\mathbb{E}[\xi^{(Q)}_{t}|\mathcal{F}_{t}]=0,\quad\mathbb{E}[\xi^{(\mu)}_{t}|\mathcal{F}_{t}]=0,\quad\mathbb{E}[(\xi^{(Q)}_{t})^{2}|\mathcal{F}_{t}]\leq K,\quad\mathbb{E}[(\xi^{(\mu)}_{t})^{2}|\mathcal{F}_{t}]\leq K.$
  3. (iii)
    Bounded iterates: The sequences $\{Q_{t}\}$, $\{\mu_{t}\}$, $\{M_{t}\}$ remain bounded almost surely.
  4. (iv)
    Memory attractor: The memory process $\{M_{t}\}$ has a compact attractor set $M_{\infty}\subset M$. This means memory size is bounded and the content distribution stabilises over time. In practice, this is achieved by: 1) Using fixed-size memory, 2) Having policies that converge to stationary distributions, 3) Using environments that produce stationary experience streams, or 4) Conservative memory update strategies.
Theorem 10 (Two-Time-Scale Convergence for Parzen-KL Policy Iteration).
Under Assumptions 6 and 9, let $(Q_{t},\mu_{t},M_{t})$ be generated by the following sample-based recursions.
  1. 1.
    Q-update (TD learning, sample form):
    $Q_{t+1}(x_{t},c_{t})$ $=Q_{t}(x_{t},c_{t})+\eta_{t}\Big[\delta_{t}(Q_{t},\mu_{0},M_{t+1})\Big],$ (12)
    $\delta_{t}(Q,\mu_{0},M)$ $:=r_{t}+\gamma\alpha\log\!\Big(\sum_{c^{\prime}\in M}\mu_{0}(c^{\prime}|x_{t+1})e^{Q(x_{t+1},c^{\prime})/\alpha}\Big)-Q(x_{t},c_{t}).$
  2. 2.
    Policy update (Parzen-KL improvement, sample form):
    $\mu_{t+1}(c|x_{t})$ $=\mu_{t}(c|x_{t})+\eta_{t}\Big[\Phi_{t}(c|x_{t};Q_{t},\mu_{0},M_{t})-\mu_{t}(c|x_{t})\Big],$ (13)
    $\Phi_{t}(c|x;Q,\mu_{0},M)$ $:=\frac{\mu_{0}(c|x)\,e^{Q(x,c)/\alpha}}{Z_{t}(x)},\qquad Z_{t}(x):=\sum_{c^{\prime}\in M}\mu_{0}(c^{\prime}|x)\,e^{Q(x,c^{\prime})/\alpha}.$
  3. 3.
    Memory update (experience replay):
    $M_{t+1}=\begin{cases}M_{t}\cup\{(s_{t},a_{t},r_{t})\}&\text{with probability }\rho_{t},\\ M_{t}&\text{otherwise.}\end{cases}$ (14)
Define the filtration $\mathcal{F}_{t}:=\sigma\big((x_{k},c_{k},r_{k}),M_{k},Q_{k},\mu_{k}:k\leq t\big)$, which collects all information available to the agent up to time (t), including past observations, the current memory state, and the value function and policy parameters. Define the mean drifts (expected operators) by
$H_{Q}(Q_{t},\mu_{t},M_{t})(x_{t},c_{t})$ $:=\mathbb{E}\!\left[\,\delta_{t}(Q_{t},\mu_{0},M_{t+1})\mid\mathcal{F}_{t}\right],$ (15)
$H_{\mu}(Q_{t},\mu_{t},M_{t})(c|x_{t})$ $:=\mathbb{E}\!\left[\,\Phi_{t}(c|x_{t};Q_{t},\mu_{0},M_{t})-\mu_{t}(c|x_{t})\mid\mathcal{F}_{t}\right].$ (16)
Then define the martingale difference noises as the residuals
$\xi^{(Q)}_{t+1}$ $:=\delta_{t}(Q_{t},\mu_{0},M_{t+1})-H_{Q}(Q_{t},\mu_{t},M_{t})(x_{t},c_{t}),$ (17)
$\xi^{(\mu)}_{t+1}(c|x_{t})$ $:=\Big(\Phi_{t}(c|x_{t};Q_{t},\mu_{0},M_{t})-\mu_{t}(c|x_{t})\Big)-H_{\mu}(Q_{t},\mu_{t},M_{t})(c|x_{t}).$ (18)
With these definitions, the recursions can be written in stochastic approximation form:
$Q_{t+1}(x_{t},c_{t})$ $=Q_{t}(x_{t},c_{t})+\eta_{t}\Big[H_{Q}(Q_{t},\mu_{t},M_{t})(x_{t},c_{t})+\xi^{(Q)}_{t+1}\Big$], (19)
$\mu_{t+1}(c|x_{t})$ $=\mu_{t}(c|x_{t})+\eta_{t}\Big[H_{\mu}(Q_{t},\mu_{t},M_{t})(c|x_{t})+\xi^{(\mu)}_{t+1}(c|x_{t})\Big].$ (20)
Moreover, $(\xi^{(Q)}_{t+1})$ and $(\xi^{(\mu)}_{t+1})$ are martingale difference sequences w.r.t. $(\mathcal{F}_{t})$:
$\mathbb{E}\!\left[\xi^{(Q)}_{t+1}\mid\mathcal{F}_{t}\right]=0,\qquad\mathbb{E}\!\left[\xi^{(\mu)}_{t+1}(c|x_{t})\mid\mathcal{F}_{t}\right]=0,$
and satisfy the conditional second-moment bounds in Assumption 9 ii.
Finally, the conclusions (a)–(c) hold:
  1. (a)
    The fast variables $(Q_{t},\mu_{t})$ track the stationary-memory soft policy-iteration fixed points:
    $\limsup_{t\to\infty}\inf_{(Q^{*}_{M},\mu^{*}_{M})\in\mathcal{F}(M_{t})}\|(Q_{t},\mu_{t})-(Q^{*}_{M},\mu^{*}_{M})\|\leq\epsilon\quad\text{a.s.}$
  2. (b)
    The joint process converges to the limiting set:
    $(Q_{t},\mu_{t},M_{t})\to\{(Q^{*}_{M},\mu^{*}_{M},M):M\in M_{\infty},\ (Q^{*}_{M},\mu^{*}_{M})\in\mathcal{F}(M)\}.$
  3. (c)
    If $M_{t}\to M_{\infty}$ a.s., then $(Q_{t},\mu_{t})\to(Q^{*}_{M_{\infty}},\mu^{*}_{M_{\infty}})$.
Appendix C presents the required Lipschitz properties of the memory updates, and Appendix D provides the detailed proof of Theorem 10. The novelty lies in extending the classical ODE method to handle structured, state-dependent episodic memory updates and their coupling with fast policy dynamics under a two–time-scale regime. The key insight is that from the perspective of the fast variables $(Q_{t},\mu_{t})$, the slow variable $M_{t}$ appears quasi-static due to the timescale separation $\rho_{t}/\eta_{t}\to 0$. This detailed proof establishes several key properties of our Read–Write Learning iteration. First, it demonstrates robustness to memory updates, showing that the algorithm remains stable even as the memory evolves, provided that memory updates occur on a slower timescale than value and policy updates. Second, it guarantees tracking behaviour, ensuring that the value function and policy rapidly adjust to reflect changes in memory, remaining close to the optimal configuration for the current memory state. Third, the analysis confirms convergence to local optima, proving that for each fixed memory configuration (M), the algorithm converges to the corresponding optimal retrieval policy ($\mu^{*}_{M}$). Finally, the timescale separation result provides practical implementation guidance, indicating that memory should be updated less frequently than the value function and policy.
Algorithm 2 Read–Write Reflective Learning (Fast Policy Update, Slow Memory Write)
1:Input: Temperature $\alpha>0$, kernel bandwidth $h>0$, discount $\gamma\in(0,1)$
2:Stepsizes: $\{\eta_{t}\}$ for policy evaluation/improvement, $\{\rho_{t}\}$ for memory updates, with $\rho_{t}/\eta_{t}\to 0$, $\sum_{t=1}^{\infty}\eta_{t}=\infty$, $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$
3:Initialise: Memory $M_{0}$, Q-function $Q_{0}$, initial Parzen prior $\mu_{0,0}(\cdot\mid x)$ (incl. void case $c_{\varnothing}$)
4:for $t=0,1,2,\ldots$ do
5:  (1) State and Prior Construction
6:  Observe $s_{t}$ and form augmented state $x_{t}=(s_{t},M_{t})$.
7:  Construct the reference prior $\mu_{0,t}(\cdot\mid x_{t})$ from $M_{t}$ using (6) and (8).
8:  (2) Read: Policy Improvement (Fast Time Scale)
9:  Define the KL-greedy retrieval policy (Eq. 11) induced by $Q_{t}$:
$\mu_{t}(c\mid x_{t})=\frac{\mu_{0,t}(c\mid x_{t})\exp(Q_{t}(x_{t},c)/\alpha)}{\sum_{c^{\prime}}\mu_{0,t}(c^{\prime}\mid x_{t})\exp(Q_{t}(x_{t},c^{\prime})/\alpha)}.$
10:  Sample a retrieval action $c_{t}\sim\mu_{t}(\cdot\mid x_{t})$.
11:  (3) LLM Act
12:  Sample an environment action $a_{t}\sim p_{\mathrm{LLM}}(\cdot\mid s_{t},c_{t})$ and execute it.
13:  (4) Environment Feedback
14:  Observe reward/feedback $r_{t}$ and next state $s_{t+1}$
15:  (5) a. Write (Slow Memory Update): update episodic memory,
$M_{t+1}\leftarrow(1-\rho_{t})M_{t}+\rho_{t}\,\mathrm{Write}(M_{t},s_{t},a_{t},r_{t},s_{t+1})\,$,
which denotes a generic slow update consistent with the assumed Lipschitz conditions.
16:  Next augmented state $x_{t+1}:=(s_{t+1},M_{t+1})$; Next prior $\mu_{0,t+1}(\cdot\mid x_{t+1})$ from $M_{t+1}$
17:   b. Write (Fast KL-regularised Evaluation): compute the target
$y_{t}:=r_{t}+\gamma\Bigg(\sum_{c^{\prime}\in\mathcal{C}(M_{t+1})}\mu_{t}(c^{\prime}\mid x_{t+1})\,Q_{t}(x_{t+1},c^{\prime})-\alpha\,\mathrm{KL}\!\big(\mu_{t}(\cdot\mid x_{t+1})\;\|\;\mu_{0,t+1}(\cdot\mid x_{t+1})\big)\Bigg),$
and update
$Q_{t+1}(x_{t},c_{t})\leftarrow(1-\eta_{t})Q_{t}(x_{t},c_{t})+\eta_{t}\,y_{t},$
which is stored in memory in the tabular case, or updated via function approximation.
18:end for
4.3 The Role of Reflection and Memory towards Optimal Policy
We are now ready to present the complete Read–Write Reflective Learning pseudocode in Algorithm 2, consistent with the workflow illustrated in Fig. 1.
To ensure practical feasibility, policy evaluation is implemented in an approximate form using feedback from the environment. Steps (2)–(4) and (5.b) constitute a fast policy iteration loop. The Read step performs a KL-regularised greedy update of the retrieval policy, while Step (5.b) approximately evaluates the resulting policy using feedback from the environment. The policy improvement is monotonic because the Read step solves a KL-regularised optimisation problem with respect to the current value function (Lemma 5). Given a fixed reference prior, the resulting update maximises the regularised expected return and therefore yields a policy that is no worse than the previous one in terms of the KL-regularised objective. Since the subsequent evaluation step provides an unbiased (or consistent) estimate of the policy value under environmental feedback, repeated application of Steps (2)–(4) and (5.b) leads to monotonic improvement of the retrieval policy on the fast time scale (Theorem 19). Step (5.a) implements the slow evolution of episodic memory. Under the two-time-scale assumptions, Theorem 10 establishes that the resulting algorithm converges to the optimal policy of the reflected MDP.
Crucially, episodic memory does not merely augment the state representation but directly parameterises the effective Bellman operator. Each memory update induces a new Reflected MDP and hence a new policy-evaluation problem. Although the $Q$-update takes the form of a standard temporal-difference step, the associated KL-evaluation operator is indexed by the current memory $M$ through the induced action set $\mathcal{C}(M)$, the memory-conditioned prior $\mu_{0}(\cdot\mid x,M)$, and the memory-dependent transition $x^{\prime}=(s^{\prime},\mathrm{Write}(M,\cdot))$.
Consequently, the read–write cycle realises a two-time-scale policy-iteration procedure: for fixed $M$, the effective MDP is stationary and the algorithm converges to the locally optimal policy induced by the current memory coverage, while memory growth drives successive improvements of the effective control problem. Since the Reflected MDP approximates the underlying environment MDP through LLM-mediated interaction (cf. (4)), a central question is whether a Read–Write Reflective Learning agent can asymptotically achieve optimal MDP performance as its episodic memory expands. We formalise this requirement via reflection consistency, which states that increasing memory coverage drives the composite policy—combining retrieval and LLM action selection—toward the true optimal policy.
[…]
Figure 4: The illustration of the assumption of LLM local sufficiency.
For an episodic memory $M\subset\mathfrak{M}$, define the coverage radius
$r_{M}:=\sup_{s\in\mathcal{S}}\min_{c\in M}d\bigl(s,s(c)\bigr),$
where $s(c)$ denotes the state component of memory case $c$. Equivalently, the statewise coverage radius is $r_{M}(s):=\min_{c\in M}d\bigl(s,s(c)\bigr)$.
Let the (deterministic) retrieval rule $c_{M}(s)$ pick any case achieving the minimum. The LLM-induced composite policy on the original state space is
$\pi_{M}(a\mid s)\;:=\;p_{\mathrm{LLM}}\big(a\mid s,c_{M}(s)\big).$
Assume bounded rewards $|\mathcal{R}(s,a)|\leq R_{\max}$ and $\gamma\in[0,1)$.
Assumption 11 (LLM local consistency).
There exists a modulus $\varepsilon_{\mathrm{LLM}}(r)\rightarrow 0$ as $r\rightarrow 0$ such that for any $s\in\mathcal{S}$ and any case $c$ with $d\big(s,s(c)\big)\leq r$,
$\mathrm{TV}\!\left(p_{\mathrm{LLM}}(\cdot\mid s,c),\,\pi^{*}(\cdot\mid s)\right)\;\leq\;\varepsilon_{\mathrm{LLM}}(r),$
where $\mathrm{TV}(p,q)$ denotes the total variation distance:
$\mathrm{TV}(p,q)=\tfrac{1}{2}\sum_{a\in\mathcal{A}}\left|p(a)-q(a)\right|\quad\text{(discrete case)},$
or
$\mathrm{TV}(p,q)=\tfrac{1}{2}\int_{\mathcal{A}}\left|p(a)-q(a)\right|\,da\quad\text{(continuous case)}.$
The function $\varepsilon_{\mathrm{LLM}}(r)$ characterises the local decision-making capability of the LLM and can be interpreted as a measure of its local optimality gap. Specifically, it quantifies how closely the LLM-induced action distribution $p_{\mathrm{LLM}}(\cdot\mid s,c)$ approximates the optimal policy $\pi^{\star}(\cdot\mid s)$ when the retrieved memory case $c$ lies within a neighbourhood of radius $r$ around the true state $s$.
A small value of $\varepsilon_{\mathrm{LLM}}(r)$ indicates that the LLM is able to infer near-optimal actions using only coarse or approximate contextual information. In other words, even when the retrieved memory does not exactly match the current state, a strong LLM can leverage its internal semantic representations, world knowledge, and reasoning ability to interpolate correctly within a local region of the state space. Consequently, a good LLM is characterised by an error function $\varepsilon_{\mathrm{LLM}}(r)$ that remains small even for relatively large values of $r$, reflecting robustness to imprecise memory retrieval and limited state coverage.
Intuitively, $\varepsilon_{\mathrm{LLM}}(r)$ captures the effective “radius of competence” of the LLM: it measures the size of the neighbourhood within which the LLM can reliably recover near-optimal behaviour from local contextual cues. This property directly determines how densely episodic memory must cover the state space in order to achieve near-optimal performance. A smaller $\varepsilon_{\mathrm{LLM}}(r)$ for larger $r$ implies that fewer memory cases are required, leading to improved sample efficiency and faster convergence of the Read–Write Reflective Learning process.
We now examine the retrieval capability. For a given state $s$, define $\delta_{M}(s)$ as the probability that the retriever selects a case whose associated state lies outside radius $r_{M}(s)$, conditioned on the existence of at least one case within this radius, due to auxiliary retrieval errors (e.g., finite-bandwidth smoothing or finite-sample density estimation):
$\delta_{M}(s)\;:=\;\mathbb{P}\!\left(d\!\left(s,s(c)\right)>r_{M}(s)\;\middle|\;\exists\,c^{\prime}\in M:d\!\left(s,s(c^{\prime})\right)\leq r_{M}(s)\right).$
Lemma 12 (Retrieval Capability).
Under Assumption 11:
$\mathrm{TV}(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s))\leq\varepsilon_{\mathrm{LLM}}(r_{M}(s))+\delta_{M}(s)$
Proof.
Let $\mathcal{E}_{s}$ denote the event that the retriever selects a good case within distance $r_{M}(s)$ of state $s$. By construction, this event occurs with probability at least $1-\delta_{M}(s)$. Conditioned on $\mathcal{E}_{s}$, Assumption 11 implies
$\mathrm{TV}\!\left(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s)\right)\leq\varepsilon_{\mathrm{LLM}}\!\left(r_{M}(s)\right).$
With the complementary event $\mathcal{E}_{s}^{c}$, the total variation distance is trivially bounded by
$\mathrm{TV}\!\left(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s)\right)\leq 1.$
Taking expectation over the retrieval event yields
$\mathrm{TV}\!\left(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s)\right)$ $\leq(1-\delta_{M}(s))\,\varepsilon_{\mathrm{LLM}}\!\left(r_{M}(s)\right)+\delta_{M}(s)$
$\leq\varepsilon_{\mathrm{LLM}}\!\left(r_{M}(s)\right)+\delta_{M}(s),$
which completes the proof. ∎
Lemma 13 (Immediate reward approximation).
For any $s\in\mathcal{S}$, with $c^{*}=c_{M}(s)$,
$\bigg|\;\mathcal{R}_{\mathrm{LLM}}\big((s,M),c^{*}\big)\;-\;\sum_{a}\pi^{*}(a\mid s)\,\mathcal{R}(s,a)\;\bigg|\;\leq\;2R_{\max}\,\mathrm{TV}\!\Big(p_{\mathrm{LLM}}(\cdot\mid s,c^{*}),\,\pi^{*}(\cdot\mid s)\Big).$
Proof.
Write $\mathcal{R}_{\mathrm{LLM}}\!=\!\sum_{a}p_{\mathrm{LLM}}(a\mid s,c^{*})\,\mathcal{R}(s,a)$ and $R^{*}(s)\!=\!\sum_{a}\pi^{*}(a\mid s)\,\mathcal{R}(s,a)$. By Hölder’s inequality and the definition of total variation, $\big|\sum_{a}(p-\pi^{*})\mathcal{R}\big|\leq\|\mathcal{R}\|_{\infty}\|p-\pi^{*}\|_{1}\leq 2R_{\max}\,\mathrm{TV}(p,\pi^{*})$. ∎
Theorem 14 (Value difference bound via policy TV).
Let $\pi^{*}$ be the optimal for the original MDP, and let $\pi_{M}$ be the composite policy defined in SRDP. Then
$\big\|V^{\pi^{*}}-V^{\pi_{M}}\big\|_{\infty}\;\leq\;\frac{2R_{\max}}{(1-\gamma)^{2}}\;\sup_{s\in\mathcal{S}}\mathrm{TV}\!\big(\pi_{M}(\cdot\mid s),\,\pi^{*}(\cdot\mid s)\big)\;=\;\frac{2R_{\max}}{(1-\gamma)^{2}}\,\Delta_{M}.$
Proof.
Consider the Bellman operators $(T^{\pi}V)(s)=\sum_{a}\pi(a\mid s)\big(\mathcal{R}(s,a)+\gamma\,\mathbb{E}[V(s^{\prime})\mid s,a]\big)$. Since $V^{\pi^{*}}$ satisfies $V^{\pi^{*}}=T^{\pi^{*}}V^{\pi^{*}}$,
$\|V^{\pi^{*}}-V^{\pi_{M}}\|_{\infty}\;\leq\;\frac{1}{1-\gamma}\,\big\|(T^{\pi^{*}}-T^{\pi_{M}})V^{\pi^{*}}\big\|_{\infty}.$
For any $s$,
$\big|(T^{\pi^{*}}-T^{\pi_{M}})V^{\pi^{*}}(s)\big|$
$=\Big|\sum_{a}\big(\pi^{*}-\pi_{M}\big)(a\mid s)\,\mathcal{R}(s,a)\;+\;\gamma\,\sum_{a}\big(\pi^{*}-\pi_{M}\big)(a\mid s)\,\mathbb{E}[V^{\pi^{*}}(s^{\prime})\mid s,a]\Big|$
$\leq 2R_{\max}\,\mathrm{TV}(\pi^{*},\pi_{M})\;+\;\gamma\,\|V^{\pi^{*}}\|_{\infty}\cdot 2\,\mathrm{TV}(\pi^{*},\pi_{M}),$
using Lemma 13’s argument and that $|\mathbb{E}V^{\pi^{*}}|\leq\|V^{\pi^{*}}\|_{\infty}$. Since $\|V^{\pi^{*}}\|_{\infty}\leq R_{\max}/(1-\gamma)$,
$\big\|(T^{\pi^{*}}-T^{\pi_{M}})V^{\pi^{*}}\big\|_{\infty}\;\leq\;2R_{\max}\,\mathrm{TV}(\pi^{*},\pi_{M})\Big(1+\tfrac{\gamma}{1-\gamma}\Big)\;=\;\frac{2R_{\max}}{1-\gamma}\,\mathrm{TV}(\pi^{*},\pi_{M}).$
Divide by $(1-\gamma)$ to get the claimed bound. ∎
Corollary 15 (Asymptotic optimality with growing memory).
Under Assumption 11 and Lemma 12, define
$\delta_{M}\;:=\;\sup_{s\in\mathcal{S}}\delta_{M}(s).$
Then
$\Delta_{M}\;:=\;\sup_{s\in\mathcal{S}}\mathrm{TV}\!\big(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s)\big)\;\leq\;\varepsilon_{\mathrm{LLM}}(r_{M})+\delta_{M}.$
If $r_{M}\to 0$ and $\delta_{M}\to 0$, then $\Delta_{M}\to 0$ and
$\sup_{s\in\mathcal{S}}\big|V^{\pi^{\star}}(s)-V^{\pi_{M}}(s)\big|\;\leq\;\frac{2R_{\max}}{(1-\gamma)^{2}}\,\Delta_{M}\;\longrightarrow\;0.$
Hence, as memory grows, the Read–Write operation drives the agent toward optimal performance in the underlying MDP.
Proof.
Lemma 12 implies $\mathrm{TV}(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s))\leq\varepsilon_{\mathrm{LLM}}(r_{M})+\delta_{M}(s)$ for all $s\in\mathcal{S}$. Taking the supremum over $s$ yields $\Delta_{M}\leq\varepsilon_{\mathrm{LLM}}(r_{M})+\delta_{M}\to 0$. Applying Theorem 14 completes the proof. ∎
Corollary 15 formalises the role of episodic memory in enabling asymptotically optimal behaviour without parameter updates. The result shows that the discrepancy between the memory-induced policy and the optimal policy of the underlying MDP decomposes into an LLM approximation error and a memory coverage error. As episodic memory grows and achieves sufficient coverage of the state space, in the sense that locally relevant memory cases enable the LLM to approximate optimal actions within neighbourhoods of each state, both error terms vanish. Consequently, the composite policy and its associated value function converge to the optimal solution. This establishes that iterative read–write reflection over expanding memory is sufficient to recover optimal performance in the limit.
5 Related Work
Our work on reflective memory dynamics and Read-Write Reflective Learning builds upon, yet also departs from, several established lines of research in reinforcement learning and large language models (LLMs). This section situates our contribution within the broader landscape of memory-based RL and clarifies the key distinctions from existing approaches.
The integration of memory mechanisms into reinforcement learning has long been explored to address the challenges of partial observability and sample complexity. Early approaches often relied on recurrent neural networks (RNNs), in particular Long Short-Term Memory (LSTM) networks [27], to maintain hidden states that capture historical information. For instance, [26] extends DQN with an LSTM to summarise past observations for Q-learning under partial observability, providing an early neural-memory approach to POMDPs [29]. However, its memory is implicit and entangled in the recurrent hidden state, making long-horizon retention, interpretability, and revision of past information difficult.
More recent methods have taken inspiration from cognitive neuroscience, especially the hippocampal processes of memory formation and retrieval in humans and animals. For instance, the Hippocampal-Augmented Memory Integration (HAMI) framework [39] incorporates symbolic indexing and hierarchical refinement based on hippocampal functions. Although HAMI shares our biological motivation, our Parzen-KL retrieval mechanism diverges in its probabilistic formulation and information-theoretic regularisation, which underpin both stability and scalability.
From another strand of memory-based RL, episodic control methods [42] such as Model Free Episodic Control [9] and Neural Episodic Control [40] use memory primarily as a direct control mechanism, retrieving past transitions to approximate values or reuse actions via similarity matching. In this setting, memory effectively acts as a nonparametric policy, and generalisation is limited to local interpolation over stored experiences [42].
In contrast, our Read–Write Reflective Learning uses episodic memory to enable reflection in large language models. Retrieved memory does not prescribe actions directly; instead, it provides local contextual grounding that allows the LLM to reason and generate actions within a semantically relevant neighbourhood. Memory thus serves as a substrate for reflective generalisation rather than as a policy itself. By combining density based similarity priors with KL regularised updates, our approach yields a principled, scalable retrieval mechanism that supports stable continual learning while exploiting the LLM’s inherent capacity for local semantic generalisation.
Recent work on Stable Hadamard Memory [30] highlights the importance of stability in memory systems, using Hadamard products to calibrate and update stored information. While complementary in spirit, our method focuses instead on retrieval optimisation through information-theoretic regularisation. These two perspectives, stability of storage and optimality of retrieval, could be integrated in future research to build more comprehensive memory architectures.
Case-based reasoning (CBR) [1, 3] has long inspired memory-driven problem solving through the retrieval and adaptation of past cases. Our prior work on CBR-based LLMs [23, 24] provides empirical evidence, particularly in data science and coding tasks, that LLMs are highly effective at revising retrieved cases, enabling strong generalisation to new tasks and overcoming a critical bottleneck in traditional CBR systems. Within our framework, CBR-based LLMs can be viewed as a stateless special case, for which we provide a principled and unified treatment.
Experience replay, a widely used mechanism in deep RL since its adoption in DQN [34], has evolved from uniform sampling to prioritised replay [44], where transitions are sampled according to temporal-difference errors. While these methods improve sample efficiency, they do not provide the structured, similarity-based retrieval required for generalisation across diverse contexts. Our approach diverges here by employing density-aware retrieval via Parzen window estimation, offering more nuanced and context-sensitive access to stored experiences.
Our work also connects to broader developments in information-theoretic reinforcement learning. Maximum entropy RL [25] introduces entropy regularisation to encourage exploration and robustness, while KL constraints in algorithms such as TRPO [45] and MPO [2] ensure stable policy updates. The distinctive aspect of our contribution is the application of KL regularisation not only to the policy but also to the retrieval process itself. This dual-level regularisation balances the exploitation of high-value memories with the exploration of diverse experiences, whilst preserving consistency with the underlying similarity structure.
Finally, from a theoretical perspective, our framework extends established convergence results in reinforcement learning. The convergence proof for our soft policy iteration builds upon entropy-regularised MDPs [35] but adapts them to the context of memory-based retrieval. Our two-time-scale analysis further draws on stochastic approximation theory [12], extending it to SRDPs where value and memory are updated simultaneously. A central outcome is the formal equivalence we establish between read (retrieval) and write operations and the fundamental processes of policy improvement and policy evaluation. This connection makes retrieval-augmented LLMs provably learnable for the first time, opening the way for reinforcement learning theory and tools to be systematically applied in this emerging setting.
6 Conclusions and Future Work
In this paper, we have established a formal theoretical foundation for continual learning in LLM agents through episodic memory and reflection, without requiring parameter fine-tuning. The core contribution is the Stateful Reflective Decision Process (SRDP), which extends the classical MDP framework by modeling reflection as a dual-action process: a retrieval stage that reads from episodic memory (policy improvement) and an action generation stage conditioned on retrieved cases. By augmenting the state space to include memory, we prove this system is Markovian and can be analysed as a Reflected MDP. We then introduce Read-Write Reflective Learning, a practical algorithm that integrates Parzen-window-based retrieval with KL-regularised soft policy iteration. Key theoretical results include: (1) guaranteed convergence of the policy iteration algorithm under bounded rewards, (2) two-timescale convergence when memory evolves slowly relative to policy updates, and (3) asymptotic optimality: as episodic memory grows to cover the state space, the composite policy provably converges to the optimal policy of the underlying MDP. This framework unifies case-based reasoning [23], retrieval-augmented generation, and reinforcement learning under a single mathematical formalism, providing the first rigorous convergence guarantees for memory-driven LLM agents.
Beyond the theoretical advances, the analysis highlights several practical implications for the design of memory-augmented LLM agents. Effective performance requires density-aware retrieval strategies to balance exploration and exploitation. The void case $c_{0}$ controls the discovery from the LLM itself whose parameter is required to be tuned. Stability can be enhanced by implementing soft, entropy-regularised improvements, while systematic memory growth helps expand state coverage. In addition, similarity metrics and bandwidths should adapt as the memory evolves to maintain efficiency and relevance.
Despite these strengths, the framework has certain limitations. It might face computational challenges as memory size increases. Moreover, the impact of embedding quality on performance is not fully addressed.
Looking ahead, several promising directions for future work emerge. In particular, we plan to internalise episodic memory within the LLM architecture itself, yielding a unified stateful neural machine, which we outline next.
6.1 Internalising Episodic Memory into the LLM Architecture
In modern contexts, LLM agents increasingly employ explicit episodic memory systems to extend context and improve performance across diverse tasks [63, 15, 18]. Yet, the integration of memory retrieval with decision-making in LLMs still lacks a firm theoretical foundation, particularly regarding the conditions under which such systems converge to optimal behaviour. Our work addresses this gap directly.
A natural next step is to integrate the episodic memory and read–write reflection mechanism of SRDP directly into the LLM architecture, yielding a single neural system whose internal state corresponds to the augmented state $(s,M)$. Recent “test-time memorisation” lines of work provide useful architectural blueprints. In particular, Titans introduces a neural long-term memory module that updates while the model is running, thereby supporting continual memorisation beyond the context window without offline retraining [6]. MIRAS further reframes modern sequence models as instances of associative memory with explicit choices of memory structure, retention, and online optimisation [4]. Our EM-LLM work in [18] introduces a human-inspired episodic memory architecture that enables large language models to operate with effectively infinite context by selectively storing, retrieving, and reusing past experiences to support long-horizon reasoning and decision-making. These developments suggest a promising direction: replace external retrieval of SRDP over a growing case base by an internal memory state that is updated online, while retaining the SRDP control view in which “read” implements policy improvement and “write” implements policy evaluation.
Our Read-Write loop differs from Titans-style test-time updates in a key way: SRDP performs stateful adaptation through experience-grounded reflection and memory evolution, rather than gradient-based parameter updates of the base model. A concrete research direction is to design a hybrid architecture in which the LLM remains fixed or slow-updated, but an internal memory module (or a small retrieval policy head) is updated online using environment feedback, with a provable two-time-scale separation between fast policy iteration and slow memory evolution. This would preserve the stability benefits of our reflected-MDP analysis while inheriting the efficiency and scalability motivations emphasised in Titans/MIRAS [6, 4] and EM-LLM [18].
More broadly, recent work on test-time training layers and nested optimisation provides a conceptual bridge between architecture and learning dynamics. Test-time training (TTT) instantiates sequence layers whose hidden state is itself a learnable model updated during inference [48]. Nested Learning argues that modern deep models can be viewed as systems of nested optimisation problems with their own internal workflows, offering a principled lens on continual learning and catastrophic forgetting [5]. An important future direction is to recast SRDP within this nested-optimisation view: the outer loop corresponds to fixed LLM representations and long-term objectives, while the inner loop implements read–write reflective adaptation as an online optimisation over an internal memory state. Such a unification could clarify when reflection-based memory updates are sufficient for continual learning, when parameter adaptation is necessary, and how to combine both without destabilising long-horizon control.
Another dimension worth exploring is principled memory consolidation. Recent work on KV-state consolidation in Transformer architectures (Bottlenecked Transformers) demonstrates that selectively rewriting and compressing internal memory can improve predictive efficiency and long-horizon reasoning without expanding context length [36]. From the perspective of SRDP, such mechanisms provide an architectural realisation of the write operation, transforming episodic experience into a compact latent state that supports future reflection. This view is further supported by our work on Agent K, which shows that structured cycles of interaction, reflection, abstraction, and consolidation enable effective experiential learning without parameter fine-tuning [21]. Together, these results point toward an architecture in which episodic memory is consolidated on a slow time scale and queried on a fast time scale, aligning architectural dynamics with the SRDP formalism.
Finally, these connections open new theoretical questions. MIRAS links sequence modelling to online optimisation and retention mechanisms [4]; integrating SRDP into this framework may yield sharper characterisations of memory coverage, approximation error, and the conditions under which reflective memory updates guarantee monotone improvement and asymptotic optimality. Establishing such results for neural internal memories (rather than external case stores) would move SRDP toward a unified neural state machine architecture in which episodic experience, reflection, and control are implemented end-to-end together.
7 Acknowledgements
We are grateful to Xue Yan for valuable discussions on reinforcement learning and kernel methods, Siyuan Guo for insights into bandit algorithms and the discovery–exploration–exploitation trade-off, and Huichi Zhou and Yihang Chen and Yuxiang Chen for assistance with implementation and proof-read. We also thank collaborators who reviewed earlier versions of this manuscript and provided helpful feedback.
Appendix A Proof of Lemma 5
Proof.
The Lagrangian for (10) with the constraint $\sum_{c}\nu(c)=1$ is
$L(\nu,\lambda)=\sum_{c}\nu(c)Q(x,c)\;-\;\alpha\sum_{c}\nu(c)\log\frac{\nu(c)}{\mu_{0}(c\mid x)}\;-\;\lambda\Big(\sum_{c}\nu(c)-1\Big).$
Taking the partial derivative w.r.t. $\nu(c)$ and setting it to zero, we use
$\frac{\partial}{\partial\nu}\Big[\nu\log(\nu/\mu_{0})\Big]=\log(\nu/\mu_{0})+1,$
to obtain
$\frac{\partial L}{\partial\nu(c)}=Q(x,c)-\alpha\Big(\log\frac{\nu(c)}{\mu_{0}(c\mid x)}+1\Big)-\lambda=0.$
Equivalently,
$\log\frac{\nu(c)}{\mu_{0}(c\mid x)}=\frac{Q(x,c)-\lambda}{\alpha}-1.$
Therefore,
$\nu(c)=\mu_{0}(c\mid x)\exp\!\Big(\tfrac{1}{\alpha}Q(x,c)\Big)\exp\!\Big(-1-\tfrac{\lambda}{\alpha}\Big).$
Imposing $\sum_{c}\nu(c)=1$ gives the normalising constant
$\exp\!\Big(-1-\tfrac{\lambda}{\alpha}\Big)=\left(\sum_{c}\mu_{0}(c\mid x)\exp\!\Big(\tfrac{1}{\alpha}Q(x,c)\Big)\right)^{-1}.$
Substituting back yields (11). ∎
Appendix B Proof of Theorem 8
Unlike standard soft reinforcement learning [25], our setting involves state-dependent base measures that vary with memory content, requiring a careful analysis of the induced operators. In particular, extending KL-regularised evaluation requires establishing that the associated log-sum-exp operations preserve contraction for both the value function and the prior distributed resulted from memory updates. Lemma 16 extends the log-sum-exp Lipschitz bound [13] to this setting with Parzen base measures and memory updates.
B.1 Bi-Lipschitz Properties
Lemma 16 (Bi-Lipschitz property of log-sum-exp with base measure).
Let $q,q^{\prime}\in\mathbb{R}^{n}$, $b,b^{\prime}\in\Delta_{n}$, and $\alpha>0$. Define
$f(b,q)\;:=\;\alpha\log\sum_{i=1}^{n}b_{i}\exp(q_{i}/\alpha).$
Then:
  1. 1.
    (Lipschitz in $q$) For any fixed $b\in\Delta_{n}$,
    $|f(b,q)-f(b,q^{\prime})|\leq\|q-q^{\prime}\|_{\infty}.$
  2. 2.
    (Lipschitz in $b$) If $q_{i}\in[Q_{\min},Q_{\max}]$ for all $i$ and $\Delta_{Q}=Q_{\max}-Q_{\min}$, then for any fixed $q$,
    $|f(b,q)-f(b^{\prime},q)|\leq\alpha e^{\Delta_{Q}/\alpha}\,\|b-b^{\prime}\|_{1}=2\alpha e^{\Delta_{Q}/\alpha}\,\mathrm{TV}(b,b^{\prime}).$
  3. 3.
    (Joint bound) Under the same boundedness condition,
    $|f(b,q)-f(b^{\prime},q^{\prime})|\leq\|q-q^{\prime}\|_{\infty}+\alpha e^{\Delta_{Q}/\alpha}\,\|b-b^{\prime}\|_{1}.$
Proof.
(1) Lipschitz in $q$ (base-measure log-sum-exp). Fix any $b\in\Delta_{n}$ and define the single-argument function
$f(q)\;:=\;f(b,q)=\alpha\log\sum_{i=1}^{n}b_{i}\exp\!\left(\frac{q_{i}}{\alpha}\right).$
The function $f$ is differentiable with gradient
$\nabla f(q)=\left[\frac{b_{i}\exp(q_{i}/\alpha)}{\sum_{j=1}^{n}b_{j}\exp(q_{j}/\alpha)}\right]_{i=1}^{n},$
which is a probability vector (all components are nonnegative and sum to one), hence
$\|\nabla f(q)\|_{1}=1\qquad\text{for all }q\in\mathbb{R}^{n}.$
By the mean value theorem for vector-valued functions (e.g., [13]), there exists $\xi$ on the line segment between $q$ and $q^{\prime}$ such that
$f(q)-f(q^{\prime})=\nabla f(\xi)^{\top}(q-q^{\prime}).$
Therefore, by Hölder’s inequality,
$|f(b,q)-f(b,q^{\prime})|=|f(q)-f(q^{\prime})|\leq\|\nabla f(\xi)\|_{1}\,\|q-q^{\prime}\|_{\infty}=\|q-q^{\prime}\|_{\infty}.$
(2) Lipschitz in $b$. Fix any $q$ and write $f(b,q)=\alpha\log(b^{\top}w)$ with $w_{i}=e^{q_{i}/\alpha}$. Then
$\nabla_{b}f(b,q)=\alpha\,\frac{w}{b^{\top}w}.$
If $q_{i}\in[Q_{\min},Q_{\max}]$ for all $i$, then
$\min_{i}w_{i}=e^{Q_{\min}/\alpha},\qquad\max_{i}w_{i}=e^{Q_{\max}/\alpha},\qquad b^{\top}w\geq\min_{i}w_{i}=e^{Q_{\min}/\alpha}.$
Hence
$\|\nabla_{b}f(b,q)\|_{\infty}\leq\alpha\,\frac{\max_{i}w_{i}}{\min_{i}w_{i}}=\alpha e^{(Q_{\max}-Q_{\min})/\alpha}=\alpha e^{\Delta_{Q}/\alpha}.$
Applying the mean value theorem in $b$ and $\ell_{1}\ell_{\infty}$ duality gives
$|f(b,q)-f(b^{\prime},q)|\leq\|\nabla_{b}f(\xi,q)\|_{\infty}\,\|b-b^{\prime}\|_{1}\leq\alpha e^{\Delta_{Q}/\alpha}\,\|b-b^{\prime}\|_{1}=2\alpha e^{\Delta_{Q}/\alpha}\,\mathrm{TV}(b,b^{\prime}).$
(3) Joint bound. Add and subtract $f(b,q^{\prime})$ and apply (1)–(2):
$|f(b,q)-f(b^{\prime},q^{\prime})|\leq|f(b,q)-f(b,q^{\prime})|+|f(b,q^{\prime})-f(b^{\prime},q^{\prime})|\leq\|q-q^{\prime}\|_{\infty}+\alpha e^{\Delta_{Q}/\alpha}\,\|b-b^{\prime}\|_{1}.$
Next, we use the above to establish the contraction for the KL-regularised evaluation.
B.2 Contraction of the KL-Regularised Evaluation Operator
We prove Theorem 17 using a single unified Lipschitz lemma for log-sum-exp with a base measure (Lemma 16), which simultaneously provides (i) Lipschitz continuity in the value argument $Q$ and (ii) stability with respect to perturbations of the base measure induced by memory.
Theorem 17 (Contraction of KL-Regularised Evaluation).
Fix a memory state $M$ and a base measure $\mu_{0}^{M}(\cdot\mid x)$ for each context $x$. Define the KL-regularised evaluation operator $\mathcal{T}^{\mu,M}_{\mathrm{KL}}$ acting on $Q:\mathcal{X}\times\mathcal{C}\to\mathbb{R}$ by
$(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q)(x,c)\;:=\;r(x,c)+\gamma\,\alpha\log\!\Big(\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x^{\prime})\exp(Q(x^{\prime},c^{\prime})/\alpha)\Big),$ (21)
where $x^{\prime}$ denotes the next context/state (possibly random) following $(x,c)$, and $\gamma\in(0,1)$ is the discount factor. Then $\mathcal{T}^{\mu,M}_{\mathrm{KL}}$ is a $\gamma$-contraction in the $\|\cdot\|_{\infty}$ norm:
$\|\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q-\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q^{\prime}\|_{\infty}\;\leq\;\gamma\|Q-Q^{\prime}\|_{\infty}.$
Consequently, $\mathcal{T}^{\mu,M}_{\mathrm{KL}}$ admits a unique fixed point $Q_{M}^{\star}$.
Proof.
Fix $M$ and let $Q,Q^{\prime}$ be arbitrary bounded functions. For any $(x,c)$, subtract (21) evaluated at $Q$ and $Q^{\prime}$:
$\big|(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q)(x,c)-(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q^{\prime})(x,c)\big|$
$\qquad=\gamma\left|\alpha\log\!\Big(\sum_{c^{\prime}}\mu_{0}^{M}(c^{\prime}\mid x^{\prime})e^{Q(x^{\prime},c^{\prime})/\alpha}\Big)-\alpha\log\!\Big(\sum_{c^{\prime}}\mu_{0}^{M}(c^{\prime}\mid x^{\prime})e^{Q^{\prime}(x^{\prime},c^{\prime})/\alpha}\Big)\right|.$
Define, for the (fixed) probability vector $b=\mu_{0}^{M}(\cdot\mid x^{\prime})$,
$f(b,q):=\alpha\log\sum_{i}b_{i}e^{q_{i}/\alpha},\qquad q:=Q(x^{\prime},\cdot),\quad q^{\prime}:=Q^{\prime}(x^{\prime},\cdot).$
By Lemma 16(1) (Lipschitz in $q$ for fixed $b$), we have
$\left|f(b,q)-f(b,q^{\prime})\right|\leq\|q-q^{\prime}\|_{\infty}=\|Q(x^{\prime},\cdot)-Q^{\prime}(x^{\prime},\cdot)\|_{\infty}\leq\|Q-Q^{\prime}\|_{\infty}.$
Therefore, for all $(x,c)$,
$\big|(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q)(x,c)-(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q^{\prime})(x,c)\big|\leq\gamma\|Q-Q^{\prime}\|_{\infty}.$
Taking the supremum over $(x,c)$ yields
$\|\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q-\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q^{\prime}\|_{\infty}\leq\gamma\|Q-Q^{\prime}\|_{\infty},$
so $\mathcal{T}^{\mu,M}_{\mathrm{KL}}$ is a $\gamma$-contraction. Existence and uniqueness of the fixed point follow from Banach’s fixed-point theorem. ∎
Remark 18 (Stability under memory perturbations).
The same Lemma 16 also provides stability of the evaluation operator to changes in memory-induced base measures. In particular, under boundedness of $Q$, Lemma 16(2) implies that for any $M,M^{\prime}$,
$\sup_{x^{\prime}}\left|\alpha\log\!\sum_{c^{\prime}}\mu_{0}^{M}(c^{\prime}\mid x^{\prime})e^{Q(x^{\prime},c^{\prime})/\alpha}-\alpha\log\!\sum_{c^{\prime}}\mu_{0}^{M^{\prime}}(c^{\prime}\mid x^{\prime})e^{Q(x^{\prime},c^{\prime})/\alpha}\right|\leq 2\alpha e^{\Delta_{Q}/\alpha}\,\|M^{\prime}-M\|.$
This perturbation bound is used later to establish the Lipschitz continuity of the moving fixed point $Q_{M}^{\star}$ and to derive the tracking recursion for two-timescale read–write learning.
We next establish monotone improvement results for Read–Write Learning Iterations (Theorem 19).
B.3 Monotone Improvement of Read-Write Iterations
Our analysis relies on the Gibbs variational principle, also known as the log-sum-exp conjugate duality [13, Section 3.3], which connects entropy-regularised optimisation to exponential-family distributions. In particular, the Gibbs principle states that
$\alpha\log\sum_{i}e^{q_{i}/\alpha}=\max_{\nu\in\Delta_{n}}\left\{\sum_{i}\nu_{i}q_{i}-\alpha H(\nu)\right\},$
where $H(\nu)=-\sum_{i}\nu_{i}\log\nu_{i}$ denotes the Shannon entropy. This identity characterises the log-sum-exp operator as the optimal value of an entropy-regularised linear objective.
Our setting extends this principle by incorporating a base measure through the Kullback–Leibler divergence $\mathrm{KL}(\nu\|\mu_{0})$, replacing entropy regularisation with relative entropy. This yields a Parzen-style exponential-family formulation and corresponds to the Bregman divergence induced by negative entropy [65].
Theorem 19 (Monotone Improvement Under Parzen-KL).
Let $Q^{\mu}$ be the fixed point of $\mathcal{T}^{\mu}_{\mathrm{KL}}$ and define $\mu^{+}$ via (11). Then:
$V^{\mu^{+}}(x)\geq V^{\mu}(x)\quad\text{for all }x$
Proof:.
By the generalised Gibbs variational principle with base measure $\mu_{0}$:
$\alpha\log\sum_{c}\mu_{0}(c\mid x)e^{Q^{\mu}(x,c)/\alpha}$ (22)
$=\max_{\nu\in\Delta(M)}\left\{\sum_{c}\nu(c)Q^{\mu}(x,c)-\alpha\,\mathrm{KL}(\nu\|\mu_{0}(\cdot\mid x))\right\}$ (23)
The maximiser is exactly $\mu^{+}$ from (11). Therefore:
$V^{\mu^{+}}(x)$ $=\sum_{c}\mu^{+}(c\mid x)Q^{\mu}(x,c)-\alpha\,\mathrm{KL}(\mu^{+}(\cdot\mid x)\|\mu_{0}(\cdot\mid x))$ (24)
$=\alpha\log\sum_{c}\mu_{0}(c\mid x)e^{Q^{\mu}(x,c)/\alpha}$ (25)
$\geq\sum_{c}\mu(c\mid x)Q^{\mu}(x,c)-\alpha\,\mathrm{KL}(\mu(\cdot\mid x)\|\mu_{0}(\cdot\mid x))$ (26)
$=V^{\mu}(x)$ (27)
The inequality follows since $\mu^{+}$ maximises the objective defining $V^{\mu}$.∎
Finally, we derive the proof of Theorem 8 for the convergence of the Read-Write Learning.
B.4 Convergence of Read-Write Iterations
Proof.
Let $(Q_{t},\mu_{t})$ be the sequence of iterates. By Theorem 19, the soft values satisfy $V^{\mu_{t+1}}(x)\geq V^{\mu_{t}}(x)$ for all $x$.
Since rewards are bounded by $R_{\max}$ and $\gamma<1$, soft values are uniformly bounded:
$V^{\mu}(x)\leq\frac{R_{\max}}{1-\gamma}+\alpha\log|M|.$
The last term uses $\mathrm{KL}(\nu\|\mu_{0})\geq-\log|M|$ for any $\nu\in\Delta(M)$. Since $\{V^{\mu_{t}}\}$ is bounded and monotone increasing, it converges pointwise to some limit $\bar{V}$. By Theorem 17, each evaluation stage converges to the unique $Q^{\mu_{t}}$. The sequence $\{Q^{\mu_{t}}\}$ lies in a compact subset of bounded functions. Any limit point $(\bar{Q},\bar{\mu})$ must satisfy:
  • Bellman consistency: $\bar{Q}=\mathcal{T}^{\bar{\mu}}_{\mathrm{KL}}\bar{Q}$
  • KL optimality: $\bar{\mu}$ maximizes the KL-regularised objective given $\bar{Q}$
By continuity of the improvement mapping (11), these conditions uniquely determine the fixed point, so the entire sequence converges. ∎
Appendix C Lipschitz Properties of Memory Updates
In practice, memory updates typically follow one of these patterns:
  1. 1.
    Experience Replay Buffer:
    $M_{t+1}=\begin{cases}M_{t}\cup\{(s_{t},a_{t},r_{t})\}&\text{with probability }\rho_{t}\\ M_{t}&\text{otherwise}\end{cases}$ (28)
  2. 2.
    Sliding Window Memory:
    $M_{t+1}=(M_{t}\setminus\{\text{oldest case}\})\cup\{(s_{t},a_{t},r_{t})\}\quad\text{with probability }\rho_{t}$ (29)
  3. 3.
    Importance-Weighted Memory:
    $M_{t+1}=M_{t}\cup\{(s_{t},a_{t},r_{t},w_{t})\}\quad\text{where }w_{t}=f(Q_{t},\mu_{t})$ (30)
Throughout, we make explicit the role of memory as an operator-valued state rather than a raw set. In other words, although memory $M_{t}$ is implemented as a buffer of cases, the algorithm accesses memory only through the induced base measure
$\mu_{0}^{M}(\cdot\mid x),$
used by the Parzen–KL prior and the KL-regularised evaluation operator.
Definition 20 (Memory distance induced by base measures).
Let $\mathcal{C}$ be a (countable) universe of case identifiers. Each memory buffer $M\subseteq\mathcal{C}$ induces a memory-dependent base measure
$\mu_{0}^{M}(\cdot\mid x)\in\Delta_{\mathcal{C}},\qquad\text{with }\;\mu_{0}^{M}(c\mid x)=0\;\;\text{for all }c\notin M.$
We define the memory distance between two memory buffers $M$ and $M^{\prime}$ at the operator level as
$\|M^{\prime}-M\|\;:=\;\sup_{x}\mathrm{TV}\!\left(\mu_{0}^{M^{\prime}}(\cdot\mid x),\mu_{0}^{M}(\cdot\mid x)\right)\;=\;\frac{1}{2}\sup_{x}\big\|\mu_{0}^{M^{\prime}}(\cdot\mid x)-\mu_{0}^{M}(\cdot\mid x)\big\|_{1}.$ (31)
All subsequent norms involving memory are understood in this sense.
Assumption 21 (Bounded $Q$-range).
For all $x\in\mathcal{X}$ and $c\in\mathcal{C}$, $Q(x,c)\in[Q_{\min},Q_{\max}]$ and $\Delta_{Q}:=Q_{\max}-Q_{\min}$.
Lemma 22 (Lipschitz continuity of the Parzen–KL targets in memory).
Assume 2021. Fix any $Q$ and define, for each $x$,
$G_{M}(x;Q)\;:=\;\alpha\log\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x)\exp\!\big(Q(x,c^{\prime})/\alpha\big),$
and
$\Phi_{M}(c\mid x;Q)\;:=\;\frac{\mu_{0}^{M}(c\mid x)\exp(Q(x,c)/\alpha)}{\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x)\exp(Q(x,c^{\prime})/\alpha)}.$
Then for any $M,M^{\prime}$ and all $x$,
$\big|G_{M}(x;Q)-G_{M^{\prime}}(x;Q)\big|$ $\leq\alpha e^{\Delta_{Q}/\alpha}\,\|\mu_{0}^{M}(\cdot\mid x)-\mu_{0}^{M^{\prime}}(\cdot\mid x)\|_{1},$ (32)
$\|\Phi_{M}(\cdot\mid x;Q)-\Phi_{M^{\prime}}(\cdot\mid x;Q)\|_{1}$ $\leq 2e^{\Delta_{Q}/\alpha}\,\|\mu_{0}^{M}(\cdot\mid x)-\mu_{0}^{M^{\prime}}(\cdot\mid x)\|_{1}.$ (33)
Consequently,
$\sup_{x}|G_{M}(x;Q)-G_{M^{\prime}}(x;Q)|\leq 2\alpha e^{\Delta_{Q}/\alpha}\,\|M^{\prime}-M\|,$
$\sup_{x}\|\Phi_{M}(\cdot\mid x;Q)-\Phi_{M^{\prime}}(\cdot\mid x;Q)\|_{1}\leq 4e^{\Delta_{Q}/\alpha}\,\|M^{\prime}-M\|.$
Proof.
Fix $x$ and abbreviate $b=\mu_{0}^{M}(\cdot\mid x)$, $b^{\prime}=\mu_{0}^{M^{\prime}}(\cdot\mid x)$, and $w(c)=\exp(Q(x,c)/\alpha)$. By Assumption 21, $w(c)\in[e^{Q_{\min}/\alpha},e^{Q_{\max}/\alpha}]$.
(i) Lipschitzness of $G_{M}$.
Write $G_{M}(x;Q)=\alpha\log(b^{\top}w)$. This map is differentiable in $b$ with $\nabla_{b}(\alpha\log(b^{\top}w))=\alpha\,w/(b^{\top}w)$. Since $b^{\top}w\geq\min_{c}w(c)=e^{Q_{\min}/\alpha}$,
$\|\nabla_{b}(\alpha\log(b^{\top}w))\|_{\infty}\leq\alpha\,\frac{\max_{c}w(c)}{\min_{c}w(c)}=\alpha e^{(Q_{\max}-Q_{\min})/\alpha}=\alpha e^{\Delta_{Q}/\alpha}.$
By the mean value theorem and Hölder’s inequality,
$|G_{M}(x;Q)-G_{M^{\prime}}(x;Q)|=|\nabla_{b}G_{\xi}^{\top}(b-b^{\prime})|\leq\|\nabla_{b}G_{\xi}\|_{\infty}\|b-b^{\prime}\|_{1}\leq\alpha e^{\Delta_{Q}/\alpha}\|b-b^{\prime}\|_{1},$
which proves (32).
(ii) Lipschitzness of $\Phi_{M}$.
Let $w_{i}=e^{q_{i}/\alpha}$ and $Z(b)=b^{\top}w$. Then $\psi(b)_{i}=b_{i}w_{i}/Z(b)$. Write
$\psi(b)-\psi(b^{\prime})=\frac{b\odot w}{Z(b)}-\frac{b^{\prime}\odot w}{Z(b^{\prime})}=\underbrace{\frac{(b-b^{\prime})\odot w}{Z(b)}}_{(A)}+\underbrace{(b^{\prime}\odot w)\Big(\frac{1}{Z(b)}-\frac{1}{Z(b^{\prime})}\Big)}_{(B)},$
where $\odot$ is elementwise product. For (A), using $Z(b)\geq\min_{i}w_{i}=e^{Q_{\min}/\alpha}$,
$\|(A)\|_{1}\leq\frac{\|(b-b^{\prime})\odot w\|_{1}}{Z(b)}\leq\frac{\max_{i}w_{i}}{\min_{i}w_{i}}\,\|b-b^{\prime}\|_{1}=e^{\Delta_{Q}/\alpha}\|b-b^{\prime}\|_{1}.$
For (B), note $|Z(b)-Z(b^{\prime})|\leq\|b-b^{\prime}\|_{1}\max_{i}w_{i}$ and $Z(b)Z(b^{\prime})\geq(\min_{i}w_{i})^{2}$. Also $\|b^{\prime}\odot w\|_{1}=Z(b^{\prime})\leq\max_{i}w_{i}$. Thus
$\|(B)\|_{1}\leq\|b^{\prime}\odot w\|_{1}\frac{|Z(b)-Z(b^{\prime})|}{Z(b)Z(b^{\prime})}\leq(\max w)\frac{\|b-b^{\prime}\|_{1}(\max w)}{(\min w)^{2}}=e^{\Delta_{Q}/\alpha}\|b-b^{\prime}\|_{1}.$
Summing the two bounds yields (33). Finally, applying Definition 20 gives $\|b-b^{\prime}\|_{1}\leq 2\|M^{\prime}-M\|$ uniformly over $x$, proving the stated consequences. ∎
Lemma 23 (Lipschitz continuity of $F_{M}$ in memory for updates (14)–(15)).
Consider the sample-form updates:
$\delta_{t}(Q,\mu_{0},M)$ $:=r_{t}+\gamma\,\alpha\log\!\Big(\sum_{c^{\prime}\in M}\mu_{0}^{M}(c^{\prime}\mid x_{t+1})e^{Q(x_{t+1},c^{\prime})/\alpha}\Big)-Q(x_{t},c_{t}),$
$\Phi_{t}(c\mid x;Q,\mu_{0},M)$ $:=\frac{\mu_{0}^{M}(c\mid x)e^{Q(x,c)/\alpha}}{\sum_{c^{\prime}\in M}\mu_{0}^{M}(c^{\prime}\mid x)e^{Q(x,c^{\prime})/\alpha}}.$
Let $F_{M}(Z)$ denote the mean-field drift of $(Q,\mu)$ induced by these targets (i.e., conditional expectations given the current state). Under 2021, there exists $L_{F}<\infty$ such that for any fixed $Z=(Q,\mu)$,
$\|F_{M^{\prime}}(Z)-F_{M}(Z)\|\leq L_{F}\,\|M^{\prime}-M\|.$
In particular, one may take $L_{F}$ proportional to $e^{\Delta_{Q}/\alpha}$.
Proof.
Memory enters the TD target only through the log-sum-exp term $G_{M}(x_{t+1};Q)$, and enters the policy target only through $\Phi_{M}(\cdot\mid x_{t};Q)$. By Lemma 22, both of these are Lipschitz in $M$ under the distance (20), uniformly in $x$. Taking conditional expectations (to pass from sample form to mean-field drift) preserves Lipschitz constants. Therefore the combined drift mapping $F_{M}(Z)$ is Lipschitz in memory, with constant $L_{F}$ depending on $\gamma,\alpha$ and the bounded range $\Delta_{Q}$, and scaling as $O(e^{\Delta_{Q}/\alpha})$. ∎
Appendix D Proof of Theorem 10
We are not ready to prove Theorem 10 using the ODE method for two-timescale stochastic approximation [12, 11].
D.1 Fast Timescale: Fixed-Memory Dynamics
Fix a memory state $M$. Let $Z=(Q,\mu)$ denote the fast variables, and define the KL-regularised operator
$Z\;\mapsto\;F_{M}(Z),$
which combines KL-regularised policy evaluation and Parzen–KL policy improvement. The stochastic updates (12) and (13) are stochastic approximations of the limiting ODE
$\frac{dZ}{d\tau}=F_{M}(Z)-Z,$ (34)
where $\tau$ denotes fast time.
Assumption 24 (Uniform Contraction).
There exists $\kappa\in(0,1)$ such that for all $M$ and all $Z,Z^{\prime}$,
$\|F_{M}(Z)-F_{M}(Z^{\prime})\|\leq\kappa\|Z-Z^{\prime}\|.$
By Theorems 17 and 19, Assumption 24 holds and the ODE (34) admits a unique globally asymptotically stable equilibrium
$Z_{M}^{\star}=(Q_{M}^{\star},\mu_{M}^{\star}).$
D.2 Perturbed Contraction and Fixed-Point Lipschitz Continuity
We next quantify how the equilibrium depends on memory.
Lemma 25 (Lipschitz Continuity of the Fixed Point).
Assume:
  1. 1.
    $F_{M}$ is a $\kappa$-contraction uniformly in $M$;
  2. 2.
    $F_{M}$ is Lipschitz in memory:
    $\|F_{M^{\prime}}(Z)-F_{M}(Z)\|\leq L_{F}\|M^{\prime}-M\|\quad\forall Z.$
Then the fixed point satisfies
$\|Z_{M^{\prime}}^{\star}-Z_{M}^{\star}\|\;\leq\;\frac{L_{F}}{1-\kappa}\,\|M^{\prime}-M\|.$
Proof.
Using $Z_{M}^{\star}=F_{M}(Z_{M}^{\star})$,
$\|Z_{M^{\prime}}^{\star}-Z_{M}^{\star}\|$ $=\|F_{M^{\prime}}(Z_{M^{\prime}}^{\star})-F_{M}(Z_{M}^{\star})\|$
$\leq\|F_{M^{\prime}}(Z_{M^{\prime}}^{\star})-F_{M^{\prime}}(Z_{M}^{\star})\|+\|F_{M^{\prime}}(Z_{M}^{\star})-F_{M}(Z_{M}^{\star})\|$
$\leq\kappa\|Z_{M^{\prime}}^{\star}-Z_{M}^{\star}\|+L_{F}\|M^{\prime}-M\|.$
Rearranging yields the claim. ∎
D.3 Slow Timescale: Memory Dynamics
Memory updates take the stochastic approximation form
$M_{t+1}=M_{t}+\rho_{t}\bigl[h(M_{t},Z_{t})+\xi^{(M)}_{t}\bigr$], (35)
where $Z_{t}=(Q_{t},\mu_{t})$, $h$ is the expected update, and $\xi^{(M)}_{t}$ is a martingale difference noise term. The associated slow ODE is
$\frac{dM}{d\varsigma}=h(M),$ (36)
with $\varsigma$ denoting slow time. By Assumption 9(iv), this ODE admits a compact attractor set $M_{\infty}$.
We also have $\|M_{t+1}-M_{t}\|=O(\rho_{t})$, a.k.a., $\sup_{t}\|M_{t+1}-M_{t}\|/\rho_{t}<\infty$. Under the distance (20), each of the following practical update schemes induces a change of order $\rho_{t}$:
  1. 1.
    Experience replay: a new case is added with probability $\rho_{t}$, changing the induced base measure by $O(\rho_{t})$.
  2. 2.
    Sliding window: one case is replaced with probability $\rho_{t}$, yielding an $O(\rho_{t})$ change.
  3. 3.
    Importance-weighted memory: weights are updated by a step $\rho_{t}$ and the induced base measure changes Lipschitzly.
Hence $\|M_{t+1}-M_{t}\|=O(\rho_{t})$ almost surely.
D.4 Tracking Lemma for Two-Timescale SA
We now state the tracking result used in the proof.
Lemma 26 (Tracking Lemma).
Suppose:
  1. 1.
    For each fixed $M$, the fast ODE (34) has a globally asymptotically stable equilibrium $Z_{M}^{\star}$;
  2. 2.
    $Z_{M}^{\star}$ is Lipschitz in $M$ as in Lemma 25;
  3. 3.
    The step sizes satisfy $\rho_{t}/\eta_{t}\to 0$.
Then
$\|Z_{t}-Z_{M_{t}}^{\star}\|\;\to\;0\quad\text{almost surely}.$
Lemma 26 follows from standard two-timescale stochastic approximation results; see [12, 11].
D.5 Derivation of the Tracking Recursion
Define the tracking error
$\epsilon_{t}=\|Z_{t}-Z_{M_{t}}^{\star}\|.$
The fast update admits the form
$Z_{t+1}=Z_{t}+\eta_{t}\bigl(F_{M_{t+1}}(Z_{t})-Z_{t}\bigr)+\eta_{t}\xi_{t},$
with martingale noise $\xi_{t}$. Using contraction of $F_{M}$, Lipschitz continuity of $Z_{M}^{\star}$, and the triangle inequality, we obtain
$\epsilon_{t+1}\;\leq\;(1-\eta_{t}\lambda)\epsilon_{t}+\eta_{t}L\|M_{t+1}-M_{t}\|+\eta_{t}\|\xi_{t}\|,$ (37)
where $\lambda=1-\kappa>0$ and $L$ is the constant from Lemma 25. Since $\|M_{t+1}-M_{t}\|=O(\rho_{t})$ and $\rho_{t}=o(\eta_{t})$, the drift term is negligible relative to the contraction.
D.6 Convergence to the Limit Set
By Lemma 26, $\epsilon_{t}\to 0$ almost surely. Since $M_{t}\to M_{\infty}$, all limit points of $(Z_{t},M_{t})$ lie in
$\mathcal{L}=\{(Z_{M}^{\star},M):M\in M_{\infty}\}.$
If $M_{t}\to M_{\infty}$, then $(Q_{t},\mu_{t})\to(Q_{M_{\infty}}^{\star},\mu_{M_{\infty}}^{\star})$ almost surely.
This completes the proof.
Appendix E Notations of MDP, SRDP, and Reflected MDP
Table 1 summarises and contrasts the notation used for the underlying Markov Decision Process (MDP), the Stateful Reflective Decision Process (SRDP) (Definition 3), and its Markovian reformulation as the Reflected MDP (Definition 4). The underlying MDP models only the environment dynamics, abstracting away any internal structure of the agent. The SRDP extends this framework by explicitly modelling episodic memory and a two-stage decision mechanism, in which a retrieval action selects a memory case that conditions the LLM’s action generation. Since the evolving memory breaks the Markov property in the original state space, the Reflected MDP restores Markovianity by augmenting the state with memory and absorbing the fixed LLM behaviour into effective transition and reward kernels. This reformulation allows classical reinforcement learning analysis to be applied while preserving the full agent–memory–LLM interaction dynamics.
Concept Underlying MDP SRDP Reflected MDP
Decision process $\mathcal{D}=\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$ $\mathcal{D}_{\mathrm{SRDP}}=\langle\mathcal{S},\mathcal{A},P,R,\gamma,\mathcal{M},p_{\mathrm{LLM}}\rangle$ $\mathcal{D}_{\mathrm{ReMDP}}=\langle\mathcal{X},\mathcal{C},P_{\mathrm{LLM}},R_{\mathrm{LLM}},\gamma\rangle$
State space $s\in\mathcal{S}$ $s\in\mathcal{S}$ (environment state) $x=(s,M)\in\mathcal{X}:=\mathcal{S}\times\mathcal{M}$
Action space $a\in\mathcal{A}$ Two-stage: retrieval $c\in\mathcal{C}(M)$, then $a\in\mathcal{A}$ Retrieval only: $c\in\mathcal{C}(M)$
Memory None Episodic memory $M\in\mathcal{M}$, evolving via $\mathrm{Write}$ Included in state: $M$ is part of $x=(s,M)$
Retrieval action
set
N/A $\mathcal{C}(M):=M$ (memory indexers) $\mathcal{C}(M):=M$
Policy $\pi(a\mid s)$ Composite policy: $\pi^{\mu}(a\mid s,M)=\sum_{c\in M}\mu(c\mid s,M)\,p_{\mathrm{LLM}}(a\mid s,c)$ Retrieval policy $\mu(c\mid x)$
LLM kernel N/A $p_{\mathrm{LLM}}(a\mid s,c)$ Absorbed into $P_{\mathrm{LLM}},R_{\mathrm{LLM}}$
Transition kernel $P(s^{\prime}\mid s,a)$ $s^{\prime}\sim P(\cdot\mid s,a)$,
$M^{\prime}=\mathrm{Write}(M,s,a,r,s^{\prime})$ $P_{\mathrm{LLM}}(x^{\prime}\mid x,c)=\sum_{a}p_{\mathrm{LLM}}(a\mid s,c)\,P(s^{\prime}\mid s,a)$
Reward $R(s,a)$ $r=R(s,a)$ $R_{\mathrm{LLM}}(x,c)=\sum_{a}p_{\mathrm{LLM}}(a\mid s,c)R(s,a)$
Markov property Markov in $s$ Non-Markov in $s$ due to evolving $M$ Markov in augmented state $x=(s,M)$
Learning focus Action selection Retrieval + memory growth Retrieval policy optimisation
Role Models environment
dynamics
Models agent–memory–LLM interaction Markovian control abstraction of SRDP
Table 1: Comparison of notations for the underlying MDP, Stateful Reflective Decision Process (SRDP), and the Reflected MDP.